### v1

# Recent Advances of Foundation Language Models-based Continual Learning: A Survey

Abstract.
Recently, foundation language models (LMs) have marked significant achievements in the domains of natural language processing (NLP) and computer vision (CV). Unlike traditional neural network models, foundation LMs obtain a great ability for transfer learning by acquiring rich commonsense knowledge through pre-training on extensive unsupervised datasets with a vast number of parameters. However, they still can not emulate human-like continuous learning due to catastrophic forgetting. Consequently, various continual learning (CL)-based methodologies have been developed to refine LMs, enabling them to adapt to new tasks without forgetting previous knowledge. However, a systematic taxonomy of existing approaches and a comparison of their performance are still lacking, which is the gap that our survey aims to fill. We delve into a comprehensive review, summarization, and classification of the existing literature on CL-based approaches applied to foundation language models, such as pre-trained language models (PLMs), large language models (LLMs) and vision-language models (VLMs). We divide these studies into offline CL and online CL, which consist of traditional methods, parameter-efficient-based methods, instruction tuning-based methods and continual pre-training methods. Offline CL encompasses domain-incremental learning, task-incremental learning, and class-incremental learning, while online CL is subdivided into hard task boundary and blurry task boundary settings. Additionally, we outline the typical datasets and metrics employed in CL research and provide a detailed analysis of the challenges and future work for LMs-based continual learning.
Continual Learning, Foundation Language Models, Pre-trained Language Models, Large Language Models, Vision-Language Models, Survey
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Scene understandingccs: Computing methodologies Cognitive roboticsccs: Computing methodologies Cognitive scienceccs: Computing methodologies Intelligent agents
1. Introduction
Recent advancements in foundation language models (LMs) have set new benchmarks in both natural language processing (NLP) (Zhao et al., 2023; Min et al., 2023; Zhou et al., 2023a) and computer vision (CV) (Wang et al., 2023b). Foundation LMs encompass three primary categories: Pre-trained Language Models (PLMs) (Min et al., 2023), Large Language Models (LLMs) (Zhao et al., 2023), and Vision-Language Models (VLMs) (Du et al., 2022). PLMs such as BERT (Kenton and Toutanova, 2019), RoBERTa (Liu et al., 2019a), and BART (Lewis et al., 2020) focus on text-based tasks and are crucial for understanding and generating language by leveraging tasks like masked language modeling during pre-training. LLMs, including models like GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), extend the capabilities of PLMs by increasing the scale of model architecture and training data, thus enhancing their generality and adaptability across a broader range of tasks. VLMs, represented by VisualBERT (Li et al., 2019), CLIP (Radford et al., 2021), LLaVA (Liu et al., 2024) and DALL-E (Ramesh et al., 2021), integrate text and image modalities to enable complicated interactions between visual and textual information. The underlying paradigm of these models involves pre-training on extensive, often unlabeled datasets to capture rich semantic information, which is subsequently fine-tuned for specific tasks or domains. This methodology not only boosts performance across various applications but also significantly enhances the models’ flexibility and task adaptability.
Figure 1. Comparison between traditional CL and Foundation language models (LMs)-Based CL.
However, these foundation models often demonstrate limitations in dynamic environments with a sequence of tasks, primarily due to their fixed parameters once training is completed. These models generally lack the capability to integrate new data or concepts without undergoing a retraining process. A significant challenge associated with training on a sequence of tasks is “catastrophic forgetting” (Kirkpatrick et al., 2017), a phenomenon where a model loses previously acquired knowledge upon learning new information. This is in stark contrast to human learning processes, which are inherently continuous and adaptive. Despite the successes of multi-task learning (MTL) and transfer learning (TL) in certain applications, they have limitations in real-world scenarios. MTL necessitates having all tasks and their data available upfront, which poses a challenge when launching a new service as the model must be retrained with all the data. Furthermore, TL is typically done with only two tasks, i.e., the source and the target, rendering it impractical for real-world online platforms with multiple target tasks. To address these challenges, it is crucial for models to process and learn the continuously expanding and diversifying datasets. This requires mechanisms that allow models to adapt to new linguistic phenomena and trends without compromising the accuracy and sensitivity towards historical data.
Consequently, continual learning (CL) (Wang et al., 2024b; Van de Ven et al., 2022), also referred to as lifelong learning (Parisi et al., 2019) or incremental learning (Zhou et al., 2023b), is a crucial area in artificial intelligence that seeks to develop systems capable of continuously updating themselves and acquiring new knowledge, without forgetting previously learned information, similar to human learning (De Lange et al., 2021). This paradigm is particularly relevant in the context of foundation language models (LMs), which are challenged by specific issues such as catastrophic forgetting (CF) and cross-task knowledge transfer (KT). Catastrophic forgetting represents a significant challenge, where a model tends to lose previously acquired knowledge upon learning new information. To address this, language models must maintain a robust grasp of past language data while adapting to new linguistic trends. Furthermore, cross-task knowledge transfer is essential for enhancing the continual learning process. Effective KT not only accelerates the learning curve for new tasks (forward transfer) but also enhances the model’s performance on prior tasks via the feedback of new knowledge (backward transfer).
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.4em,font=,, where level=2text width=8.4em,font=,, where level=3text width=10.0em,font=,, where level=4text width=10.5em,font=,, [ Foundation LMs-based CL, ver [ Offline Continual
Learning  (§4) [ Domain-Incremental
Learning  (§4.1) [ PLMs-based DIL  (§4.1.1) [ LFPT5 (Qin and Joty, 2022b), B-CL (Ke and Xu, 2021), ELLE (Qin et al., 2022), AdapterCL (Madotto et al., 2020), RMR_ DSE (Li et al., 2022a), DEMIX (Geng et al., 2021) , CLASSIC (Ke et al., 2021b),
CPT (Ke et al., 2022), C-PT (Zhu et al., 2022), CL-KD (Castellucci et al., 2021), PlugLM (Cheng et al., 2023b), Pretr (Cossu et al., 2022), AEWC (Lee, 2017), Continual DAP-training (Ke et al., 2023) , leaf, text width=44.6em ] ] [ LLM-based DIL  (§4.1.2) [ COPR (Zhang et al., 2023b), LAMOL (Sun et al., 2020a), RVAE_LAMOL (Wang et al., 2022a), Adapt-Retrieve-Revise (Zhang et al., 2023c), Lifelong-MoE (Chen et al., 2023),
CKL (Jang et al., 2022b), DACP (Xie et al., 2023), CPPO (Zhang et al., 2024), EcomGPT-CT (Ma et al., 2023) , leaf, text width=44.6em ] ] [ VLMs-based DIL  (§4.1.3) [ S-Prompt (Wang et al., 2022b), Medical AI (Yi et al., 2023), VQACL (Zhang et al., 2023d) , leaf, text width=44.6em ] ] ] [ Task-Incremental
Learning  (§4.2) [ PLMs-based TIL  (§4.2.1) [ PP (Razdaibiedina et al., 2023), CTR (Ke et al., 2021a), MeLL (Wang et al., 2021a), LINC (Liu and Mazumder, 2021), ERDA (Qin and Joty, 2022a), PCLL (Zhao et al., 2022), CLIF (Jin et al., 2021a), ConTinTin (Yin et al., 2022),
HMI (Maekawa et al., 2023), Adaptive Compositional Modules (Zhang et al., 2022b), DYNAINST (Mok et al., 2023), Conure (Yuan et al., 2021), TERACON (Kim et al., 2023),
ERNIE 2.0 (Sun et al., 2020b), RecyclableTuning (Qin et al., 2023) , leaf, text width=44.6em ] ] [ LLMs-based TIL  (§4.2.2) [ Conpet (Song et al., 2023), InstructAlign (Cahyawijaya et al., 2023), Continual-T0 (Scialom et al., 2022a), DynaMind (Du et al., 2023), ELM (Jang et al., 2023), O-LoRA (Wang et al., 2023a), JARe (PENG et al., 2024) , leaf, text width=44.6em ] ] [ VLMs-based TIL  (§4.2.3) [ Medical AI (Yi et al., 2023), CTP (Zhu et al., 2023), ZSCL (Zheng et al., 2023), MoE-Adapters4CL (Yu et al., 2024), TRIPLET (Qian et al., 2023) , leaf, text width=44.6em ] ] ] [ Class-Incremental
Learning  (§4.3) [ PLMs-based CIL  (§4.3.1) [ EPI (Wang et al., 2023c), IDBR (Huang et al., 2021), PAGeR (Varshney et al., 2022), ENTAILMENT (Xia et al., 2021), ExtendNER (Monaikul et al., 2021), PLE (Li et al., 2022b), DE&E (Wójcik et al., 2023),
SRC (Liu et al., 2019b) , leaf, text width=44.6em ] ] [ VLMs-based CIL  (§4.3.2) [ MoE-Adapters4CL (Yu et al., 2024), VLM-PL (Kim et al., 2024), Adaptation-CLIP (Liu et al., 2023a), PROOF (Zhou et al., 2023c), LGCL (Khan et al., 2023), ZSCL (Zheng et al., 2023),
CLAP (Jha et al., 2024), GMM (Cao et al., 2024) , leaf, text width=44.6em ] ] ] ] [ Online Continual
Learning  (§5) [ Hard Task
Boundary  (§5.1) [ PLMs-based HTB  (§5.1.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b), ProgModel (Shen et al., 2019) , leaf, text width=44.6em ] ] ] [ Blurry Task
Boundary  (§5.2) [ PLMs-based BTB  (§5.2.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b) , leaf, text width=44.6em ] ] [ VLMs-based BTB  (§5.2.2) [ CBA (Wang et al., 2023d), MVP (Moon et al., 2023), DKR (Cui et al., 2024) , leaf, text width=44.6em ] ] ] ] ]
Figure 2. Taxonomy of foundation language models for continual learning.
Recent advancements in continual learning methodologies have substantially enhanced the adaptability and knowledge retention capabilities of foundational language models (LMs). These developments are crucial for addressing complex challenges previously observed in CL. Researchers have formulated innovative strategies to mitigate these challenges, thereby enabling LMs to maintain high performance across a variety of tasks while continually integrating new knowledge (Mehta et al., 2021; Cossu et al., 2022; Lee et al., 2023). Notable successes have been documented in diverse downstream tasks, such as aspect-based sentiment analysis, where continual learning enables dynamic adaptation to evolving aspects and sentiments (Ke et al., 2021b). Similarly, in dialogue generation, the novel technologies assist models in refining and expanding their conversational capabilities through ongoing interactions (Scialom et al., 2022b). In text classification, continual learning facilitates the incorporation of new categories and adjustments to shifts in text distributions without the need for complete retraining (Razdaibiedina et al., 2023). Moreover, in the realm of visual question answering, continual learning is essential for updating the models’ capabilities to process and respond to new types of visual content and queries (Zhang et al., 2023d; Qian et al., 2023). The aforementioned works underscore the potential of continual learning to significantly boost the performance of foundation LMs.
In the domain of continual learning, there has been a significant paradigm shift from traditional methodologies to those that integrate foundation LMs (See Figure 1). First, foundation LMs demonstrate enhanced generalization and transfer learning abilities across diverse tasks owing to their broad pre-training on large-scale datasets. The model has specialized transfer capability to quickly adapt to downstream tasks with only a few samples. Consequently, it is crucial to mitigate the degradation of both the zero-shot transfer and history task abilities in LMs while facilitating the acquisition of new skills. Second, due to the substantial number of parameters in foundation LMs, it is crucial to employ parameter-efficient techniques (Han et al., 2024), such as prompt tuning (Liu et al., 2022) and adapters (Mundra et al., 2024), to update parameters without comprehensive retraining. Third, the foundation LMs possess the capability to follow instructions through instructional learning (Dong et al., 2022; Ouyang et al., 2022), enabling more dynamic and context-aware interactions.
This review systematically categorizes these strategies and technologies into two core areas: offline continual learning and online continual learning (Figure 2). We first give detailed definitions and scenarios to format the setting of offline and online CL, where offline CL includes domain-incremental, task-incremental and class-incremental CL, and online CL includes hard task boundary and blurry task boundary. These learning strategies are further subdivided into methods based on Pre-trained Language Models (PLMs), Large Language Models (LLMs), and Vision-Language Models (VLMs). Then, we summarize the related papers about traditional methods, continual pre-training methods, parameter-efficient tuning methods and instruction-based methods. Finally, we static the main datasets from various perspectives and review the key metrics to evaluate the forgetting and transferring of the models.
The main contributions of this survey paper can be summarized as follows.
  • We thoroughly review the existing literature on foundation LMs-based CL approaches, which integrate foundation LMs with CL to learn new knowledge without retraining the models. It is quite different from traditional CL since foundation LMs have great abilities of transfer learning, zero-shot and instruction following with huge parameters.
  • We give the definitions of different settings and categorize these studies into various classes to better understand the development of this domain. In addition to the traditional methods like replay, regularization and parameter-isolation-based algorithms, we also summarize the works about continual pre-training methods, parameter-efficient tuning methods and instruction tuning-based methods.
  • We provide the characters of existing datasets for CL and present the main metrics to evaluate the performance of preventing forgetting and knowledge transfer.
  • We discuss the most challenging problems of foundation LMs-based CL and point out promising future research directions in this field.
The paper is organized as follows. In Section 2, we review the mainly related surveys about continual learning. Then, we introduce the base settings and learning modes of continual learning in Section 3, including the definitions and scenarios of CL. Furthermore, we present the related studies about offline continual learning, which can be divided into domain-incremental learning, task-incremental learning and class-incremental learning in Section 4. In Section 5, we focus on online continual learning, including hard task boundary and blurry task boundary settings. The typical datasets and metrics are provided in Section 6 and 7. Finally, we analyze the challenge and further work in Section 8 and give the conclusion in Section 9.
2. Related Surveys
2.1. Continual Learning
Early examinations of Continual Learning (CL) have provided broad coverage, as observed in surveys such as Parisi et al. (Parisi et al., 2019). Recently, Wang et al. (Wang et al., 2024b) conduct a comprehensive survey that categorizes five key strategies in CL: regularization-based, replay-based, optimization-based, representation-based, and architecture-based approaches. This survey reflects an effort to organize and understand the diverse methodologies employed in the field. Notably, there is a growing focus on class-incremental setting (Zhou et al., 2023b; Masana et al., 2022; Belouadah et al., 2021) and replay-based approaches (Hayes et al., 2021), reflecting the increasing granularity of research interests within the CL domain.
2.2. Continual Learning for Computer Vision
In the realm of computer vision, De et al. (De Lange et al., 2021) address the pressing challenge of continual learning, specifically in the task-incremental setting, where tasks arrive sequentially with clear boundaries. They introduce a novel stability-plasticity trade-off framework tailored for continual learners and undertake a comprehensive experimental analysis, comparing the efficacy of 11 methodologies across three benchmarks. Qu et al. (Qu et al., 2021) present a comprehensive examination of continual learning, highlighting its vital role in the accumulation of knowledge from sequential data streams. This research investigates a range of approaches, encompassing regularization, knowledge distillation, memory-based methods, and more. These approaches are systematically categorized by their characteristics and applications in computer vision. Moreover, Mai et al. (Mai et al., 2022) focus on the realm of online continual learning within image classification, addressing catastrophic forgetting. This study evaluates the efficacy of state-of-the-art methodologies across diverse memory and data configurations. Masana et al. (Masana et al., 2022) present a comprehensive performance evaluation of class-incremental methods applied to image classification. The experiments span a range of scenarios, incorporating large-scale datasets and varied network architectures. Belouadah et al. (Belouadah et al., 2021) pay more attention to class-incremental learning algorithms for visual tasks. This study defines the essential properties of incremental learning algorithms, offers a unified formalization of the class-incremental learning problem, and provides an evaluation framework for detailed analysis.
2.3. Continual Learning for NLP
Biesialska et al. (Biesialska et al., 2020) address the challenge of continual learning within Natural Language Processing (NLP), wherein conventional architectures struggle to accommodate new tasks without compromising previously acquired knowledge. The research presents an extensive review of methods, including rehearsal, regularization, and architectural approaches, all designed to mitigate the aforementioned challenge. In a similar vein, Ke et al. (Ke and Liu, 2022) offer a focused survey on continual learning within the NLP domain, providing a comprehensive examination of various continual learning settings, methodologies, and challenges. This work presents an in-depth analysis of state-of-the-art approaches and extends original CL settings to be more general and up-to-date. Additionally, it emphasizes the significance of knowledge transfer within NLP and the challenges posed by inter-task class separation.
2.4. Continual Learning for Other Domains
Recent surveys, surveys like (Zhang and Kim, 2023; Shaheen et al., 2022; Lesort et al., 2020) investigate the advancements in incremental learning for neural recommendation systems and continual learning (CL) in robotics, respectively. Zhang et al. (Zhang and Kim, 2023) make a notable contribution to narrow the gap between academic research and industrial applications through the introduced Incremental Update Recommendation Systems (IURS). They highlight the imperative for real-time updates with streaming data and focus on the distinctive challenges posed by IURS in contrast to traditional Batch Update Recommendation Systems (BURS) and offer a thorough examination of existing literature and evaluation methodologies in this domain. Shaheen et al. (Shaheen et al., 2022) provide a comprehensive overview of contemporary approaches for CL within real-world contexts. Their analysis focuses on learning algorithms that efficiently handle large sequential datasets within computational and memory constraints. The survey also explores challenges in applying CL to autonomous systems, comparing methods across metrics such as computational efficiency, memory utilization, and network complexity. In the field of robotics, it is essential for agents to adapt and interact with their environment using a continuous stream of observations. Thus, Lesort et al. (Lesort et al., 2020) explore CL within this domain, defining CL as a paradigm where both data distribution and learning objectives evolve dynamically. They emphasize the challenges in evaluating CL algorithms in robotic applications and introduce a novel framework alongside metrics tailored to effectively present and assess CL methodologies.
This paper centers on the crucial advancements in CL as applied to foundational language models, which have obtained significant success in the fields of NLP and multimodal. We categorize existing works into offline and online CL based on PLMs, LLMs, and VLMs.
Figure 3. The setting of different offline continual learning tasks, including task-incremental learning, class-incremental learning and domain-incremental learning. The samples with different classes (domains) are marked with various shapes (colors).
Figure 4. The setting of different online continual learning tasks, including hard task boundary arriving and blurry task boundary arriving. The samples with different classes (domains) are marked with various shapes (colors).
3. Settings and Learning Modes of CL
3.1. Basic Formulation
Continual learning is an advanced method in machine learning. Within this framework, the model is sequentially trained across a diverse array of tasks denoted as $t$ within the set $T=\{1,2,...,N\}$, where each task $t$ is associated with its individual dataset $X_{t}=\{(x^{(t)}_{i},y^{(t)}_{i})\}^{|X_{t}|}_{i=1}$. Here, $x^{(t)}_{i}$ represents an individual training example, and $y^{(t)}_{i}$ denotes the corresponding class label for task $t$, while $|X_{t}|$ indicates the total number of samples in task $t$. However, the data distributions between any two tasks $t$ and $t^{\prime}$ are distinct ($p(X_{t})\neq p(X_{t^{\prime}})$ for all $t\neq t^{\prime}$). This distinction presents a fundamental challenge in managing the diversity of data distributions across multiple tasks. This setup necessitates the model to learn new knowledge while retaining past information.
Continual learning encompasses two principal paradigms: offline and online continual learning. These paradigms define how data arrives and how the model updates its knowledge over time.
  • Offline Continual Learning: This setting involves learning across a series of tasks, with each task fully presented before handling the next task. For each task $t$, the model trains on the entire dataset $D_{t}$ through multiple epochs. The model progresses to task $t+1$ only upon achieving the desired proficiency on task $t$.
  • Online Continual Learning: This setting operates within a dynamic framework wherein the model learns knowledge from a stream of data points or mini-batches presented sequentially. Additionally, the model lacks access to the entire dataset for a given task. This setting closely mirrors real-world scenarios characterized by continuous data flow, compelling the model to adapt in real time.
3.2. Typical Scenarios
3.2.1. Offline Continual Learning
Offline CL (Figure 3) comprises three principal scenarios, each distinguished by distinct characteristics: Domain-Incremental Learning, Task-Incremental Learning, and Class-Incremental Learning.
  • Domain-Incremental Learning (DIL): The model aims to process diverse data distributions. Specifically, in DIL, while the data distributions $p(X_{t})$ in task $t$ and $p(X_{t^{\prime}})$ in task $t^{\prime}$ are different, their task types and class labels remain consistent. The task identities (task IDs) are not required.
  • Task-Incremental Learning (TIL): The model is designed to handle a series of tasks, each with unique objectives. The classes within these tasks may or may not be disjoint. The boundaries of each task are clear, and task IDs are provided during both the training and testing phases.
  • Class-Incremental Learning (CIL): The model is designed to continually learn new class information while retaining knowledge of previously learned classes. For tasks $t$ and $t^{\prime}$, while they might share the same task type (such as classification), their class sets $C_{t}$ and $C_{t^{\prime}}$ are distinct. Moreover, the task IDs are only available during training.
In summary, Domain-Incremental Learning concentrates on adapting the model to the shifts in input data distributions while maintaining consistency in tasks and classes. Task-Incremental Learning necessitates the model’s ability to learn and retain task-specific knowledge over successive tasks. On the other hand, Class-Incremental Learning highlights the gradual integration of new classes into the model’s recognition capabilities without compromising knowledge of previously learned classes.
3.2.2. Online Continual Learning
In online continual learning (See Figure 4), existing research is categorized into two configurations based on the arrival pattern of tasks: ”Hard Task Boundary” and ”Blurry Task Boundary”:
  • Hard Task Boundary: The arrival of tasks follows a strictly structured and sequential process. Data from the preceding task is completely processed before transitioning to the next task, ensuring no overlap of data between tasks.
  • Blurry Task Boundary: The distinction between tasks is less clear, similar to real-world scenarios. Data from different tasks are intermixed, making it difficult to pinpoint when one task ends and another begins.
In both setups, the main challenge lies in achieving the balance of learning new data while preserving previously gained knowledge, often termed as catastrophic forgetting. Numerous approaches, such as experience replay (Qin et al., 2022; Sun et al., 2020a), elastic weight consolidation (EWC) (Kirkpatrick et al., 2017), and progressive neural networks (Ke and Xu, 2021; Ke et al., 2021a), have emerged to address this issue. Each method comes with its unique strengths and weaknesses upon the task arrival configuration.
4. Offline Continual Learning
4.1. Domain-Incremental Learning
4.1.1. PLMs-based DIL
Traditional Methods.
Continual Learning methodologies are frequently employed in the context of Pre-trained Language Models (PLMs), encompassing approaches such as replay-based, regularization-based, and parameter-isolation-based algorithms. Li et al. (Li et al., 2022a) present a regularization-centric framework for Lifelong Learning (LLL) termed RMR-DSE, specifically tailored for sequential operation across multiple domains. Unlike conventional strategies requiring incremental memory allocation, RMR-DSE employs a recall optimization mechanism, driven by regularization, to selectively retain important parameters from preceding tasks. Furthermore, it incorporates a domain drift estimation algorithm to address embedding space shifts. Castellucci et al. (Castellucci et al., 2021) propose a Knowledge Distillation-based Continual Learning (CL-KD) approach with a Teacher-Student framework. When the student model is trained on a new language, the teacher model also transfers its knowledge of supported languages to the student model. Lee et al. (Lee, 2017) introduce a domain-agnostic framework that employs a two-phase training process, initially using synthetic data to learn general conversational patterns and subsequently using human-computer dialogs in customer support. Moreover, the Adaptive Elastic Weight Consolidation algorithm is applied to adjust the loss function, ensuring the balance between acquiring new knowledge and retaining previously learned information.
Gururangan et al. (Gururangan et al., 2022) develop a parameter-isolation-based method, named DEMIX layers, which consists of a collection of experts. The dynamic addition or removal of these experts can enhance the model’s ability to adapt to new domains while maintaining robust performance in previously established ones. PlugLM (Cheng et al., 2023b) is a pre-training model equipped with a differentiable plug-in memory (DPM) designed for domain-adaptive continual training. The core concept behind this approach is to separate the knowledge storage from model parameters using an adaptable key-value memory structure. It enables the explicit retrieval and utilization of knowledge stored within the DPM.
(a) CLASSIC
(b) Lifelong-MoE
(c) S-Prompts
Figure 5. Frameworks in DIL: CLASSIC (PLM-based) (Ke et al., 2021b), Lifelong-MoE (LLM-based) (Chen et al., 2023), S-Prompts (VLM-based) (Wang et al., 2022b).
Continual Pre-training Methods.
Continual domain-adaptive pre-training (DAP-training) (Ke et al., 2023) is based on two main ideas: (1) the general knowledge in the LM and the knowledge gained from prior domains are crucial to mitigating CF and enhancing cross-task knowledge transfer. This is achieved through soft-masking units based on their importance, and (2) the model is designed to develop complementary representations of both the current domain and prior domains, thereby facilitating the integration of knowledge. The key novelty of Continual DAP-training is a soft-masking mechanism that directly controls the update to the LM. Cossu et al. (Cossu et al., 2022) formalize and explore the dynamics of continual pre-training (Pretr) scenarios across language and vision domains. In this framework, models undergo continuous pre-training on a sequential stream of data before subsequent fine-tuning for various downstream tasks.
Parameter-Efficient Tuning Methods.
Due to the huge parameters of LMs, parameter-efficient tuning methods like adaptors (Houlsby et al., 2019; Poth et al., 2023) and p-tuning (Liu et al., 2022) are used for domain-incremental CL (Madotto et al., 2020; Ke and Xu, 2021; Ke et al., 2021b; Zhu et al., 2022).
The adapter architecture incorporates a skip-connection to minimize the number of parameters. A notable exemplar of this approach is AdapterCL (Madotto et al., 2020), which employs residual adapters tailored specifically for task-oriented dialogue systems. This framework, comprising 37 domains, is structured to facilitate continual learning across four important aspects: intent recognition, state tracking, natural language generation, and end-to-end processing. In a related vein, Ke et al. (Ke and Xu, 2021) introduce B-CL model to tackle critical challenges in CL for Aspect-Based Sentiment Classification. B-CL integrates continual learning adapters within capsule network architectures. Aiming at mitigating catastrophic forgetting during fine-tuning, the CLASSIC model (Ke et al., 2021b) presents an innovative solution by deploying adapters to tap into BERT’s capabilities (Figure 5(a)). A novel contrastive continual learning strategy is used to facilitate the transfer of knowledge across tasks and distill insights from previous tasks to subsequent ones. It also effectively eliminates the necessity for task identifiers during testing. Furthermore, Continual PostTraining (CPT) (Ke et al., 2022) introduces two continual learning plug-in modules, termed CL-plugins, embedded within each transformer layer of RoBERTa.
Prompt tuning (Liu et al., 2022), or P-tuning, introduces trainable continuous prompts into the sequence of input word embeddings, while the language model remains frozen. To tackle the challenge of CL under limited labeled data, Qin et al. (Qin and Joty, 2022b) propose a Lifelong Few-shot Language Learning framework (LFPT5). In this framework, prompt tuning, replay and regularization strategies are leveraged. When presented with a new task, the model generates pseudo-labeled samples representative of prior domains. The training process then incorporates these pseudo-labeled samples alongside new task-specific data. Additionally, the KL divergence loss is employed to maintain label consistency between the previous and the current model. Furthermore, Zhu et al. (Zhu et al., 2022) introduced Continual Prompt Tuning (C-PT) as a methodology to address the challenges of continual learning within dialogue systems. C-PT facilitates knowledge transfer between tasks through continual prompt initialization, query fusion, memory replay, and a memory-guided technique.
Instruction Tuning-based Methods.
Instruction tuning-based methods involve transforming a given task into natural language instructions. Qin et al. (Qin et al., 2022) propose ELLE, a novel approach aimed at effectively incorporating continuously expanding streaming data into pre-trained language models (PLMs). It consists of two fundamental components: (1) function-preserved model expansion, which enhances knowledge acquisition efficiency by changing the width and depth of an existing PLM, and (2) pre-trained domain prompts, which significantly enhance the adaptation for downstream tasks by effectively segregating the diverse knowledge acquired during pre-training phases.
4.1.2. LLMs-based DIL
Traditional Methods.
In many practical scenarios, retraining Language Models (LMs) is challenging due to resource constraints and data privacy concerns. Zhang et al. (Zhang et al., 2024) introduce Continual Proximal Policy Optimization (CPPO) to address this issue. CPPO integrates sample-wise weighting into the Proximal Policy Optimization (PPO) algorithm, effectively balancing policy learning and knowledge retention. Zhang et al. (Zhang et al., 2023b) propose Continual Optimal Policy Regularization (COPR), which calculates the optimal policy distribution without the partition function and uses the previous optimal policy to regularize the current policy. Sun et al. (Sun et al., 2020a) introduce LAMOL, which generates pseudo-samples from previous tasks while training on a new task. It effectively mitigates knowledge loss without requiring additional memory or computational resources. Building on this framework, Wang et al. (Wang et al., 2022a) developed RVAE_LAMOL, which integrates a residual variational autoencoder (RVAE) to encode input data into a unified semantic space, thereby enhancing task representation. This model also incorporates an identity task to enhance the model’s discriminative ability for task identification. To enhance training efficacy, the Alternate Lag Training (ALT) is devised to segment the training process into multiple phases.
To reduce hallucinations in specialized domains such as the Chinese legal domain, Zhang et al. (Zhang et al., 2023c) propose a novel domain adaptation framework, named Adapt-Retrieve-Revise (ARR). It consists of three steps: adapting a 7-billion-parameter language model for initial responses, retrieving corroborative evidence from an external knowledge base, and integrating these to refine the final response with GPT-4. Gogoulou et al. (Gogoulou et al., 2023) study the pros and cons of updating a language model when new data comes from new languages – the case of continual learning under language shift. They feed various languages into the model to examine the impact of pre-training sequence and linguistic characteristics on both forward and backward transfer effects across three distinct model sizes. A new continual learning (CL) problem, named Continual Knowledge Learning (CKL), is introduced by Jang et al. (Jang et al., 2022b). To assess CKL approaches, the authors establish a benchmark and metric measuring knowledge retention, updating, and acquisition.
Continual Pre-training Methods
LLMs have demonstrated remarkable proficiency in tackling open-domain tasks. However, their application in specific domains faces notable challenges, encompassing the lack of domain-specific knowledge, limited capacity to utilize such knowledge, and inadequate adaptation to domain-specific data formats. To address these issues, researchers have explored a novel approach known as continual pre-training, aiming to adapt LLMs to specific domains (Xie et al., 2023; Ma et al., 2023; Cheng et al., 2023a). Among these studies, Cheng et al. (Cheng et al., 2023a) draw inspiration from human learning patterns to develop a novel method that transforms raw corpora into reading comprehension texts. Furthermore, they discover that while training directly on raw data enhances the model’s domain knowledge, it significantly hurts the question-answering capability of the model.
Domain-adaptive Continual Pre-training (DACP) uses a large domain corpus, leading to high costs. To reduce these, Xie et al. (Xie et al., 2023) propose two strategies: Efficient Task-Similar Domain-Adaptive Continual Pre-training (ETS-DACP) and Efficient Task-Agnostic Domain-Adaptive Continual Pre-training (ETA-DACP). ETS-DACP is tailored to improve performance on specific tasks, building task-specific foundational LLMs. Conversely, ETA-DACP selects the most informative samples across the domain. Given the high cost of training LLMs from scratch and limited annotated data in certain domains, Ma et al. (Ma et al., 2023) propose a novel model called EcomGPT-CT. It employs a fusion strategy to exploit semi-structured E-commerce data. Moreover, multiple tasks are designed to assess LLMs’ few-shot in-context learning ability and zero-shot performance after fine-tuning.
Parameter-Efficient Tuning Methods.
Chen et al. (Chen et al., 2023) present an innovative Lifelong Learning framework, termed Lifelong-MoE, which leverages a Mixture-of-Experts (MoE) architecture (Figure 5(b)). This architecture enhances the model’s capacity by incorporating new experts, where previously trained experts and gating mechanisms are frozen.
4.1.3. VLMs-based DIL
Vision-language models (VLMs) have demonstrated superiority in domain-incremental learning contexts. Yi et al. (Yi et al., 2023) integrate VLMS with continual learning methodologies to develop a general-purpose medical AI. Moreover, their study highlights the significance of data-efficient adaptation algorithms in minimizing the necessity for extensive labeling when transitioning to new domains or tasks. Furthermore, the prompt text is utilized to master the pre-trained knowledge embedded within VLMs. Aiming to independently learn prompts across disparate domains by using pre-trained VLMs, S-Prompt (Wang et al., 2022b) is devised (Figure 5(c)). This method encompasses techniques for acquiring image prompts and introduces an innovative methodology for language-image prompt acquisition. Prompt learning is conducted separately, utilizing a unified cross-entropy loss function during training. During inference, a K-NN (k-nearest neighbors) technique is employed to discern the domain.
In the domain of visual question answering, Zhang et al. (Zhang et al., 2023d) introduce VQACL, a novel framework designed to effectively integrate data from both visual and linguistic modalities. This integration is achieved through a dual-level task sequence that enhances the model’s performance on complex multimodal tasks. Central to VQACL is a compositionality test that evaluates the model’s ability to generalize new skill and concept combinations. The framework also incorporates a novel representation learning strategy that differentiates between sample-specific (SS) and sample-invariant (SI) features. SS features capture distinctive attributes of individual inputs, enhancing output uniqueness, while SI features, derived from category prototypes, ensure essential characteristics are retained.
(a) HMI
(b) DynaMind
(c) TRIPLET
Figure 6. Frameworks in TIL: HMI (PLM-based) (Maekawa et al., 2023), DynaMind (LLM-based) (Du et al., 2023), TRIPLET (VLM-based) (Qian et al., 2023).
4.2. Task-Incremental Learning
4.2.1. PLMs-based TIL
Traditional Methods.
Drawing inspiration from neurobiological mechanisms, Maekawa et al. (Maekawa et al., 2023) present an inventive approach known as Hippocampal Memory Indexing (HMI) to augment the generative replay technique. HMI leverages hippocampal memory indexing to integrate compressed representations of prior training instances, facilitating selective guidance for the generation of training samples. This methodological refinement contributes to heightened specificity, balance, and overall quality of the replayed samples. In tackling the Continual Few-Shot Relation Learning (CFRL) challenge, Qin et al. (Qin and Joty, 2022a) propose ERDA as a solution, drawing upon replay- and regularization-based techniques. The ERDA framework integrates embedding space regularization and data augmentation strategies to effectively confront the task of acquiring new relational patterns from scarce labeled instances while mitigating the risk of catastrophic forgetting. Wang et al. (Wang et al., 2019b) introduced a memory-based approach to continual learning, termed Episodic Memory Replay (EMR). This method leverages working memory by selectively replaying stored samples during each iteration of learning new tasks, thereby facilitating the integration of new knowledge while preserving previously acquired information.
Conure (Yuan et al., 2021) is a framework that effectively manages multiple tasks by leveraging the redundancy of parameters in deep user representation models. Initially, it prunes less critical parameters to make room for new, task-specific ones. Subsequently, it incorporates these new parameters while retaining key parameters from prior tasks, facilitating positive transfer learning. To prevent the loss of previously acquired knowledge, it maintains these essential parameters in a fixed state. Notably, TERACON (Kim et al., 2023) utilizes task-specific soft masks to isolate parameters, which not only targets parameter updates during training but also clarifies the relationships between tasks. This method includes a novel knowledge retention module that utilizes pseudo-labeling to mitigate the well-known problem of catastrophic forgetting.
Ke et al. (Ke et al., 2021a) introduced a model known as CTR, which employs innovative techniques such as CL-plugin and task masking to tackle the issue of catastrophic forgetting and to facilitate knowledge transfer across tasks. These strategies are particularly effective when utilized in conjunction with pre-trained language models, such as BERT, enhancing their adaptability and efficacy. In the specific context of User Intent Classification (UIC) within large-scale industrial applications, Wang et al. (Wang et al., 2021a) introduce a novel methodology, MeLL, which utilizes a BERT-based text encoder to generate robust, dynamically updated text representations for continual learning. MeLL combines global and local memory networks to preserve prototype representations across tasks, acting as a meta-learner that rapidly adapts to new challenges. It employs a Least Recently Used (LRU) policy for efficient global memory management and minimizes parameter growth.
Recent advancements in conversational AI have concentrated on mitigating the limitations of traditional chatbots, which are dependent on static knowledge bases and extensive manual data annotation. The Lifelong INteractive learning in Conversation (LINC) methodology represents a significant stride in this direction, embodying a dynamic learning framework that mimics human cognitive processes during interactions (Liu and Mazumder, 2021; Mazumder and Liu, 2024; Liu, 2020). Structured around an Interaction Module, a Task Learner, and a Knowledge Store, LINC enables chatbots to dynamically integrate and utilize knowledge within conversational contexts (Figure LABEL:fig:LINC). This approach not only facilitates real-time information extraction and learning from user interactions but also addresses challenges such as managing erroneous inputs, refining conversational skills, and maintaining user engagement, thereby enhancing the chatbots’ linguistic and interactive capabilities.
Continual Pre-training Methods.
Continual pre-training represents a paradigm wherein pre-trained language models (PLMs) are progressively enhanced through the assimilation of new knowledge from expanding datasets. Sun et al. (Sun et al., 2020b) introduced a framework called ERNIE 2.0, which incrementally constructs pre-training tasks, enabling the acquisition of complex lexical, syntactic, and semantic nuances embedded within the training corpora. This model deviates from traditional fixed-task training, instead employing continual multi-task learning to refine its capabilities continually. Further advancing the field of continual pre-training, RecyclableTuning (Qin et al., 2023) introduces a novel concept of recyclable tuning that features two distinct strategies: initialization-based and distillation-based methods. The former uses fine-tuned weights of existing PLMs as the basis for further enhancements, capitalizing on the established parametric relationships. Conversely, the distillation-based approach harnesses outdated weights to maintain knowledge continuity and efficiency in successor models.
Parameter-Efficient Tuning Methods.
Expounding upon the crucial need for more efficacious knowledge integration, Zhang et al. (Zhang et al., 2022b) propose a pioneering strategy that entails the integration of Adaptive Compositional Modules alongside a replay mechanism. These modules are designed to dynamically adjust to new tasks and are supplemented by pseudo-experience replay, significantly enhancing knowledge transfer. This framework is distinguished by its adaptive integration of modules within transformer architectures, skillfully orchestrating the interactions between existing and new modules to address emerging tasks. Additionally, the implementation of pseudo-experience replay promotes efficient knowledge transfer across these modules. Concurrently, Jin et al. (Jin et al., 2021a) introduce the Continual Learning of Few-Shot Learners (CLIF) challenge, wherein a model accumulates knowledge continuously across a series of NLP tasks. Their investigation delves into the impact of continual learning algorithms on generalization capabilities and advances a novel approach for generating regularized adapters.
Instruction Tuning-based Methods.
Zhao et al. (Zhao et al., 2022) propose the Prompt Conditioned VAE for Lifelong Learning (PCLL) specifically designed for task-oriented dialogue (ToD) systems. PCLL employs a conditional variational autoencoder influenced by natural language prompts to generate high-quality pseudo samples, effectively capturing task-specific distributions. Additionally, a distillation process is integrated to refine past knowledge by reducing noise within pseudo samples. Razdaibiedina et al. (Razdaibiedina et al., 2023) introduce Progressive Prompts (PP), a novel approach to continual learning in language models. PP addresses catastrophic forgetting without resorting to data replay or an excessive proliferation of task-specific parameters. This method involves acquiring a fresh soft prompt for each task, gradually appending it to previously learned prompts while keeping the base model unchanged. The ConTinTin paradigm (Yin et al., 2022) develops a computational framework for sequentially mastering a series of new tasks guided by explicit textual instructions. It synthesizes projected outcomes for new tasks based on instructions, while facilitating knowledge transfer from prior tasks (forward-transfer) and maintaining proficiency in previous tasks (backward-transfer).
In the realm of lifelong in-context instruction learning aimed at enhancing the target PLM’s instance- and task-level generalization performance as it observes more tasks, DYNAINST is devised by Mok et al. (Mok et al., 2023). DYNAINST integrates the principles of parameter regularization and experience replay. The regularization technique employed by DYNAINST is tailored to foster broad local minima within the target PLM. In order to devise a memory- and computation-efficient experience replay mechanism, they introduce Dynamic Instruction Replay, comprising Dynamic Instance Selection (DIS) and Dynamic Task Selection (DTS). DIS and DTS dynamically determine the selection of instances and tasks to be stored and replayed, respectively.
4.2.2. LLMs-based TIL
Recent attention has focused on the convergence of large language models (LLMs) with continual learning methodologies, exemplified by significant contributions such as those by Wang et al. (Wang et al., 2024a) and Peng et al. (PENG et al., 2024). Benefiting from vast corpora and advanced hardware infrastructure, LLMs showcase remarkable capabilities in language comprehension and generation. However, challenges arise in scenarios involving sequential tasks, where LLMs often exhibit a decline in performance known as catastrophic forgetting.
Traditional Methods.
DynaMind, introduced by Du et al. (Du et al., 2023), emerges as a pioneering framework that intricately incorporates memory mechanisms and modular operators, enhancing the precision of LLM outputs. Comprising three essential components, DynaMind includes a memory module dedicated to storing and updating acquired knowledge, a modular operator for processing input data, and a continual learning module responsible for dynamically adjusting LLM parameters in response to new knowledge. Furthermore, Luo et al. (Luo et al., 2023) conduct a thorough investigation into catastrophic forgetting (CF) in Large Language Models (LLMs) during continual fine-tuning. Their experiments across various domains, including domain knowledge, reasoning, and reading comprehension, reveal the prevalence of CF in LLMs ranging from 1b to 7b scale, with severity increasing with model size. Comparative analysis between decoder-only BLOOMZ and encoder-decoder mT0 indicates that BLOOMZ exhibits less forgetting. Additionally, LLMs demonstrate the ability to mitigate language bias during continual fine-tuning. Contrasting ALPACA against LLAMA, the study highlights ALPACA’s superiority in preserving knowledge and capacity, suggesting that general instruction tuning aids in mitigating CF during subsequent fine-tuning phases. This research provides valuable insights into CF dynamics in LLMs, offering strategies for knowledge retention and bias mitigation.
Wang et al. (Wang et al., 2024a) present TRACE, an innovative benchmark meticulously designed to evaluate continual learning capabilities in LLMs. Comprising eight distinct datasets spanning challenging tasks, including domain-specific challenges, multilingual capabilities, code generation, and mathematical reasoning, TRACE serves as a comprehensive evaluation platform. The authors rigorously examine the effectiveness of conventional Continual Learning (CL) methods when applied to LLMs within the TRACE framework. Peng et al. (PENG et al., 2024) propose the Joint Adaptive ReParameterization (JARe) framework, enhanced with Dynamic Task-related Knowledge Retrieval (DTKR), to facilitate adaptive adjustment of language models tailored to specific downstream tasks. This innovative approach leverages task distribution within the vector space, aiming to streamline and optimize the continual learning process seamlessly.
Parameter-Efficient Tuning Methods.
Large language models (LLMs) encounter several substantial challenges that limit their practical applications. These include high computational requirements, significant memory demands, and a tendency toward catastrophic forgetting. Such limitations highlight the need for ongoing research into more efficient and robust approaches to training and deploying these models. Continual Parameter-Efficient Tuning (ConPET) (Song et al., 2023) is designed for the continuous adaptation of LLMs across diverse tasks, leveraging parameter-efficient tuning (PET) strategies to enhance both efficiency and performance. ConPET encompasses two primary modes: Static ConPET and Dynamic ConPET. Static ConPET adapts techniques previously utilized in smaller models for LLMs, thus minimizing tuning costs and reducing the risk of overfitting. Conversely, Dynamic ConPET enhances scalability by employing distinct PET modules for varying tasks, supplemented by a sophisticated selector mechanism.
Moreover, the ELM strategy (Jang et al., 2023) involves initially training a compact expert adapter on the LLM for each specific task, followed by deploying a retrieval method to select the most appropriate expert LLM for each new task. Furthermore, Wang et al. (Wang et al., 2023a) have proposed orthogonal low-rank adaptation (O-LoRA), a straightforward yet efficacious method for facilitating continual learning in language models. O-LoRA mitigates catastrophic forgetting during task acquisition by employing distinct low-rank vector subspaces maintained orthogonally to minimize task interference. This method highlights the potential of orthogonal subspace techniques in improving the adaptability of language models to new tasks without compromising previously acquired knowledge.
Instruction Tuning-based Methods.
Scialom et al. (Scialom et al., 2022a) introduced Continual-T0, an innovative framework aimed at exploring the capabilities of large language models (LLMs) through continual learning, incorporating rehearsal techniques. A central aspect of this approach is the employment of instruction tuning, a key strategy designed to enhance the adaptability and effectiveness of LLMs when encountering novel tasks. Leveraging self-supervised pre-training, Continual-T0 demonstrates exceptional proficiency in mastering new language generation tasks while maintaining high performance across a diverse range of 70 previously encountered datasets. Despite the demonstrated proficiency of LLMs in adhering to instructions, their ability to generalize across underrepresented languages remains suboptimal. In response, InstructAlign (Cahyawijaya et al., 2023) is proposed to address this challenge by aligning newly introduced languages with those previously learned, which possess abundant linguistic resources, thereby mitigating instances of catastrophic forgetting. The core novelty of this approach lies in its advancement of language adaptation methodologies for instruction-tuned LLMs, with particular emphasis on integrating underrepresented languages.
4.2.3. VLMs-based TIL
The long-term sustainability of pre-trained visual-language models (VLMs) is increasingly under scrutiny due to their dependence on continually expanding datasets. Although these models demonstrate robust performance across a diverse range of downstream tasks, the incessant growth of real-world data poses substantial challenges to the sustainability of traditional offline training methodologies.
Traditional Methods.
CTP (Zhu et al., 2023) employs topology preservation and momentum contrast to maintain consistent relationships within sample mini-batches across tasks, thereby preserving the distribution of prior embeddings. CTP also introduces the P9D dataset, comprising over one million image-text pairs across nine domains, aimed at visual language continuous pre-training (VLCP). Zheng et al. (Zheng et al., 2023) address the issue of zero-shot transfer degradation in visual language models by introducing the Zero-Shot Continual Learning (ZSCL) method. This novel approach utilizes a label-free dataset to facilitate distillation in the feature space, coupled with the application of weight regularization within the parameter space. Furthermore, they introduce a new benchmark, the Multi-Domain Task Incremental Learning (MTIL), designed to evaluate incremental learning strategies across various domains, thereby enhancing the assessment of such methods. Moreover, ZSCL has also been adapted for use in the CIL setting, further broadening its applicability
Instruction Tuning-based Methods.
By decoupling prompts and prompt interaction strategies, TRIPLET (Qian et al., 2023) effectively captures complex interactions between modalities. This includes specific designs for visual, textual, and fused prompts, as well as how to interact between different tasks through these prompts and retain crucial information, thereby reducing catastrophic forgetting. Decoupled prompts are designed to separate prompts in terms of multi-modality, layer-wise, and complementary, with each type of prompt containing learnable parameters intended to capture modality-specific knowledge from pre-trained visual-language models and training data. The prompt interaction strategies consist of three main components: the Query-and-Match Strategy, the Modality-Interaction Strategy, and the Task-Interaction Strategy. These components work together to enhance the model’s adaptability to different tasks and its memory for old tasks.
Moreover, COIN (Chen et al., 2024) introduces a new continuous instruction tuning benchmark designed to evaluate the performance of Multimodal Large Language Models (MLLMs) in the sequential instruction fine-tuning paradigm. CoIN includes 10 commonly used data sets covering 8 task categories, ensuring the diversity of instructions and tasks. In addition, the trained model is evaluated from two aspects: instruction following and general knowledge, which respectively evaluate the consistency with human intention and the preservation of knowledge for reasoning. CoIN converts commonly used visual-linguistic datasets into instruction fine-tuning data formats by using different templates to explore the behavior of MLLMs in continuous instruction fine-tuning. This method takes into account the diversity between different tasks and attempts to enhance the model’s adaptability to new and old tasks through diversified instruction templates. To alleviate the catastrophic forgetting problem of MLLMs, CoIN introduces the MoELoRA method. This method reduces forgetting by using different experts to learn knowledge on different tasks and using gate functions to regulate the output of these experts.
Parameter-Efficient Methods.
MoE-Adapters4CL (Yu et al., 2024) introduces a parameter-efficient continual learning method to mitigate long-term forgetting in incremental learning of visual-language models. Their approach involves dynamically extending the pre-trained CLIP model to accommodate new tasks by integrating a Mixture-of-Experts (MoE) adapter. Specifically, MoE consists of several LoRA adapter experts and routers, where the router calculates gating weights and uses the $TopK$ function to select the $k$ most relevant experts for learning the current task. To maintain the zero-shot recognition capabilities of the visual-language model, a Distribution Discriminative Automatic Selector (DDAS) is further introduced, which can automatically route in-distribution and out-of-distribution inputs to the MoE adapters and the original CLIP, respectively. Furthermore, the MoE-Adapters4CL framework has also been adapted for use in the CIL setting.
4.3. Class-Incremental Learning
4.3.1. PLMs-based CIL
Traditional Methods.
The study presented in (Monaikul et al., 2021) introduces ExtendNER, a novel framework for continual learning in Named Entity Recognition (NER) that obviates the need for extensive re-annotation. This framework employs a knowledge distillation (KD) technique, wherein a pre-existing named entity recognition (NER) model, termed the ”teacher”, imparts knowledge to a newly developed model, termed the ”student”. The student model, designed to identify new entity types, progressively learns by emulating the teacher model’s responses to a new dataset. This method allows the student to acquire the ability to recognize new entities while retaining knowledge of previously identified ones. Additionally, Liu et al. (Liu et al., 2019b) propose an innovative method for learning distributed representations of sentences. This method initiates with the configuration of sentence encoders using features independent of any specific corpus, followed by iterative updates through Boolean operations on conceptor matrices. This technique ensures that the encoders maintain their performance on existing datasets while adapting effectively to new data.
(a) IDBR
(b) PLE
(c) Adaptation-CLIP
Figure 7. Frameworks in CIL: IDBR (PLM-based) (Huang et al., 2021), PLE (PLM-based) (Li et al., 2022b), Adaptation-CLIP (VLM-based) (Liu et al., 2023a).
Huang et al. (Huang et al., 2021) introduced an innovative methodology known as Information Disentanglement-based Regularization (IDBR) to address the enduring challenges associated with continual text classification. This method effectively disentangles the hidden spaces of text into task-generic and task-specific representations, employing distinct regularization strategies to enhance knowledge retention and facilitate generalization. Furthermore, the integration of two auxiliary tasks, namely next sentence prediction and task-id prediction, serves to augment the learning process by reinforcing the separation between generic and specific representational spaces.
Instruction Tuning-based Methods.
Introduced by Varshney et al. (Varshney et al., 2022), Prompt Augmented Generative Replay (PAGeR) is a method that enables continual learning in intent detection without retaining past data. PAGeR leverages pre-trained language models to generate intent-specific utterances specific to new intents while preserving accuracy on existing ones. Unlike exemplar replay, which stores specific examples, PAGeR is structured to selectively maintain relevant contexts for each specified intent, which are then employed as generation prompts. This approach combines utterance generation and classification into one model, enhancing knowledge transfer and optimization.
Parameter-Efficient Tuning Methods.
In the domain of task-oriented dialogue systems, Continual Few-shot Intent Detection (CFID) focuses on recognizing new intents with few examples. Li et al. (Li et al., 2022b) propose a Prefix-guided Lightweight Encoder (PLE) to address this by using a parameter-efficient tuning method that combines a Continual Adapter module with a frozen Pre-trained Language Model (PLM) and a Prefix-guided Attention mechanism to reduce forgetting. To further mitigate forgetting, the Pseudo Samples Replay (PSR) strategy reinforces prior knowledge by replaying crucial samples from previous tasks. The Teacher Knowledge Transfer (TKT) strategy uses distillation to transfer task-specific knowledge to maintain performance on new tasks. Additionally, the Dynamic Weighting Replay (DWR) strategy dynamically adjusts weights of previous tasks to balance new knowledge acquisition with the revision of old tasks, navigating the variability and potential negative impacts of prior tasks
The Efficient Parameter Isolation (EPI) method, introduced in Wang et al. (Wang et al., 2023c), assigns unique subsets of private parameters to each task alongside a shared pre-trained model. This approach ensures precise parameter retrieval and has been shown to outperform non-caching methods in continual language learning benchmarks, while remaining competitive with caching methods. Furthermore, EPI employs random static masking to reduce storage requirements, increasing its viability in resource-constrained environments.
In complex system environments, efficient and adaptable machine learning architectures are crucial, especially for classification tasks with sequentially presented data. Wójcik et al. (Wójcik et al., 2023) devise a novel architecture, Domain and Expertise Ensemble (DE&E), comprising a feature extractor, classifier, and gating mechanism. The feature extractor employs a multi-layer neural network to convert input data into embeddings, while the classifier, a mixture of binary class-specific experts, leverages a gating mechanism to select the appropriate expert for the current input dynamically. This Mixture of Experts-based method promotes incremental learning by training experts with class-specific samples and combines their outputs during testing to derive the final classification.
4.3.2. VLMs-based CIL
Traditional Methods.
Kim et al. (Kim et al., 2024) introduce VLM-PL, a novel approach for class incremental object detection that incorporates new object classes into a detection model without forgetting old ones. Utilizing a visual-language model (VLM), this method enhances the pseudo-labeling process to improve the accuracy and performance of object detection in continual learning settings. VLM-PL starts with a pre-trained detector to generate initial pseudo-labels, which are then validated through a visual-language evaluation using a specially designed hint template. Accurate pseudo-labels are retained and combined with ground truth labels to train the model, ensuring it remains proficient in recognizing both new and previously learned categories. Cao et al. (Cao et al., 2024) introduce a framework for a Generative Multi-modal Model (GMM) that leverages large language models for class-incremental learning. This innovative approach entails the generation of labels for images by employing an adapted generative model. Following the production of detailed textual descriptions, a text encoder is utilized to extract salient features from these descriptions. These extracted features are subsequently aligned with existing labels to ascertain the most fitting label for classification predictions.
PROOF (Zhou et al., 2023c) develops a method to enhance model memory retention when adapting to downstream tasks. This method involves a projection technique that maps pre-trained features into a new feature space designed to preserve prior knowledge. Additionally, to effectively utilize cross-modal information, PROOF introduces a fusion module that employs an attention mechanism. This module adjusts both visual and textual features simultaneously, enabling the capture of semantic information with enhanced expressive power. Recently, Jha et al. (Jha et al., 2024) present a new approach for adapting pre-trained vision-language models like CLIP to new tasks without forgetting previous knowledge. It employs a Variational Inference framework to probabilistically model the distribution of visual-guided text features, enhancing fine-tuning reliability by accounting for uncertainties in visual-textual interactions. Key to CLAP is the visual-guided attention (VGA) module, which aligns text and image features to prevent catastrophic forgetting. Additionally, CLAP includes lightweight, task-specific inference modules that learn unique stochastic factors for each task, allowing continuous adaptation and knowledge retention.
Parameter-Efficient Tuning Methods.
Liu et al. (Liu et al., 2023a) introduce Adaptation-CLIP, which employs three strategies for CLIP’s continual learning: linear adapter, self-attention adapter, and prompt tuning. The first two strategies add a linear layer and a self-attention mechanism, respectively, after the image encoder while freezing the remaining architecture. The third, prompt tuning, integrates trained prompts into the text encoder to enhance task comprehension and splices these with prior prompts to maintain continuity. To prevent catastrophic forgetting, a parameter retention strategy preserves significantly altered parameters from $M_{t-1}$ to $M_{t}$, ensuring stability and effective continual learning.
Instruction Tuning-based Methods.
Khan et al. (Khan et al., 2023) introduce two notable advancements: an enhanced prompt pool key query mechanism and category-level language guidance. The key query mechanism uses CLS features to improve prompt selection, featuring key replacement with a fixed CLS tag and dynamic mapping to task-level language representations, thereby enhancing accuracy and robustness across different tasks. Meanwhile, category-level language guidance is implemented in the vision transformer to better align output features with category-specific language representations, significantly improving task handling and category differentiation, leading to improved model performance.
5. Online Continual Learning
5.1. Hard Task Boundary
5.1.1. PLMs-based HTB
The Hard Task Boundary (HTB) setting has been developed to enable continuous knowledge acquisition by learning models from a dynamically changing stream of textual data, without the need for dataset identifiers. For example, Shen et al. (Shen et al., 2019) have implemented HTB in slot filling, Michieli et al. (Michieli et al., 2023) have utilized it in audio classification, and Vander et al. (Vander Eeckt et al., 2023) have explored its use in automatic speech recognition.
Traditional Methods.
Continual learning (CL) methodologies, particularly pertinent to online scenarios, encompass a variety of approaches. These include parameter-isolation-based methods (de Masson D’Autume et al., 2019; Wang et al., 2020), replay-based methods (Holla et al., 2020) and regularization-based methods (Vander Eeckt et al., 2023; Liu et al., 2021b). MBPA++ (de Masson D’Autume et al., 2019) introduces a framework for lifelong language learning, enabling a pre-trained model to learn continually from textual examples without requiring labeled datasets. It employs an episodic memory system with sparse experience replay and local adaptation techniques to prevent catastrophic forgetting. Extending this framework, Meta-MBPA++ (Wang et al., 2020) integrates three core lifelong learning principles, enhancing performance in text classification and question-answering tasks while using only 1% of the typical memory usage. Liu et al. (Liu et al., 2021b) introduce a regularization-based strategy, CID, for lifelong intent detection. This method uses cosine normalization, hierarchical knowledge distillation, and inter-class margin loss to tackle the challenges of data imbalances in the lifelong intent detection (LID) task, aiming to mitigate the negative impacts associated with these imbalances.
(a) MBPA++
(b) CBA
Figure 8. Frameworks in Online Continual Learning: MBPA++ (PLM-based HTB/BTB) (de Masson D’Autume et al., 2019), CBA (VLM-base BTB) (Wang et al., 2023d).
5.2. Blurry Task Boundary
5.2.1. PLMs-based BTB
MBPA++ (de Masson D’Autume et al., 2019) and Meta-MBPA++ (Wang et al., 2020) can also exemplify models capable of adapting to environments with indistinct task boundaries. TPEM (Geng et al., 2021) employs a tripartite approach within an encoder-decoder framework, consisting of pruning, expanding, and masking techniques. Pruning helps preserve essential information from previous tasks, expansion increases the model’s capacity to accommodate new tasks, and masking reduces interference from previous tasks’ weights, thus enhancing learning efficiency. Originally, online meta-learning (OML) (Javed and White, 2019) and a neuromodulatory meta-learning algorithm (ANML) (Beaulieu et al., 2020) were intended to continuously learn new sequences of tasks during the testing phase. Holla et al. (Holla et al., 2020) adapt these algorithms for a conventional continual learning context, where the focus is on evaluating performance on previously encountered tasks. The enhanced versions, named OML-ER and ANML-ER, incorporate an episodic memory module designed for experience replay.
5.2.2. VLMs-based BTB
Traditional Methods.
Cui et al. (Cui et al., 2024) propose the Dynamic Knowledge Rectification (DKR) framework, designed to mitigate the propagation of incorrect information in foundation LMs. The DKR framework operates by initially leveraging an existing model to identify and exclude obsolete or erroneous knowledge when confronted with new data. Subsequently, a rectification process is employed to amend these inaccuracies while ensuring the preservation of valid data associations. This process is especially vital when integrating new data, as it prevents the perpetuation of outdated or incorrect information. In cases where data is inaccessible through existing models, DKR utilizes paired ground-truth labels to support the continuous evolution of knowledge bases, thereby enhancing the model’s accuracy.
Parameter-Efficient Tuning Methods.
Wang et al. (Wang et al., 2023d) introduce the Continual Bias Adaptor (CBA) (Figure 8(b)), a novel method designed to enhance the efficacy of online CL by mitigating catastrophic forgetting. The CBA method dynamically modifies the classifier’s network to adapt to significant shifts in data distribution observed during training, thereby preserving knowledge from earlier tasks. Notably, the CBA module can be deactivated during testing phases, eliminating additional computational burdens or memory demands. This feature highlights the practicality and applicability of the CBA in real-world scenarios.
Instruction Tuning-based Methods.
In the BTB scenario, the Mask and Visual Prompt tuning (MVP) method, as detailed by Moon et al. (Moon et al., 2023), addresses challenges like intra- and inter-task forgetting and class imbalance effectively. MVP features instance-wise logit masking to prevent irrelevant information retention, contrastive visual prompt tuning loss to ensure consistent prompt selection, gradient similarity-based focal loss to focus on overlooked samples, and adaptive feature scaling to balance the integration of new knowledge with existing data retention.
Online Learning for CV
Recent advancements in class-incremental online learning predominantly address computer vision tasks, encapsulating various methodologies such as regularization-based (Yin et al., 2021; Gu et al., 2022; Guo et al., 2022; Bang et al., 2022), replay-based (Aljundi et al., 2019; Shim et al., 2021; Caccia et al., 2021; Lin et al., 2023; Zhang et al., 2022a), distillation-based (He et al., 2020; Fini et al., 2020; Koh et al., 2022), and gradient-based approaches (Jin et al., 2021b; Guo et al., 2023; Chen et al., 2020; Yoon et al., 2021; Ororbia et al., 2022). Among these, Koh et al. (Koh et al., 2021) introduced the Class-Incremental Blurry (CLIB) model, which distinguishes itself through its task-free nature, adaptability to class increments, and prompt response to inference queries, showing superior performance over traditional continual learning methods. In a related vein, Gunasekara et al. (Gunasekara et al., 2023) explore the Online Streaming Continual Learning (OSCL), a hybrid of Stream Learning (SL) and Online Continual Learning (OCL), which integrates aspects of both domains. Furthermore, issues of shortcut learning and bias in online continual learning have been tackled by Wei et al. (Wei et al., 2023) and Chrysalis et al. (Chrysakis and Moens, 2023). Additionally, Semola et al. (Semola et al., 2022) propose Continual-Learning-as-a-Service (CLaaS), a service model that leverages continual learning to monitor shifts in data distribution and update models efficiently. This array of developments highlights the dynamic capabilities of online continual learning frameworks to effectively address complex challenges across a variety of computer vision tasks.
Table 1. The statistics information of the existing CL datasets. #D/T/C means the number of domains/tasks/classes, respectively.
Datasets #Train #Val #Test #Total CL Settings NLP Problems Language #D/T/C
Offline
Progressive Prompts (Razdaibiedina et al., 2023) - - - - TIL Sentiment analysis, topic classification, boolean QA, QA, paraphrase detection, word sense disambiguation, natural language inference English 15 tasks
MeLL (Wang et al., 2021a) 1,430,880 173,781 118,240 1,722,901 TIL Intent classification English 1184 tasks
Continual-T0 (Scialom et al., 2022b) 800,000 - 33,382 833,382 TIL Text Simplification, Headline Generation with Constraint, Haiku Generation, Covid QA, Inquisitive Question Generation, Empathetic Dialogue Generation, Explanation Generation, Twitter Stylometry English 8 tasks
COPR (Zhang et al., 2023b) - - - - TIL QA tasks, summary task, positive file review generation task English 3 tasks
Adaptive Compositional Modules (Zhang et al., 2022b) 50,725 - 27,944 78,669 TIL Natural language generation, SQL query generation, Summarization and Task-oriented dialogue English 8 tasks
CODETASKCL (Yadav et al., 2023) 181,000 9,700 10,000 200,700 TIL Code generation, code translation, code summarization, and code refinement Hybrid 4 tasks
Lifelong Simple Questions (Wang et al., 2019b) - - - - TIL single-relation questions English 20 tasks
Lifelong FewRel (Wang et al., 2019b) - - - - TIL few-shot relation detection English 10 tasks
InstrDialog (Zhang et al., 2023a) 9,500 950 1,900 12,350 TIL Dialogue state tracking, dialogue generation, intent identification English 19 tasks
InstrDialog++ (Zhang et al., 2023a) 3,800 1,900 3,800 9,500 TIL Dialogue Generation, Intent Identification, Dialogue State Tracking, Style Transfer, Sentence Ordering, Word Semantics, Text Categorization, Pos Tagging, Fill in The Blank, Program Execution, Question Generation, Misc, Coherence Classification, Question Answering, Summarization, Commonsense Classification, Wrong Candidate Generation and Toxic Language Detection English 38 tasks
ConTinTin (Yin et al., 2022) - - - - TIL Question generation tasks (QG), answer generation tasks (AG), classification tasks (CF), incorrect answer generation tasks (IAG), minimal modification tasks (MM) and verification tasks (VF) English 61 tasks
Tencent TL (Yuan et al., 2021) - - - - TIL personalized recommendations and profile predictions English 6 tasks
Movielens (Yuan et al., 2021) - - - - TIL personalized recommendations and profile predictions English 3 tasks
NAVER Shopping (Kim et al., 2023) - - - - TIL search query prediction tasks English 6 tasks
TRACE (Wang et al., 2024a) 40,000 - 16,000 56,000 TIL Domain-specific task, Multi-lingual task, Code completion task, Mathematical reasoning task Hybrid 8 tasks
ABSC (Ke and Xu, 2021) 3,452 150 1,120 4,722 DIL Aspect-based sentiment classification English 19 Domains
DecaNLP (Sun et al., 2020a) 169,824 - 32,116 201,940 DIL Question answering, semantic parsing, sentiment analysis, semantic role labeling, and goal-oriented dialogue English 5 domains
Foundational text classification (Sun et al., 2020a) 115,000 - 7,600 122,600 DIL News classification, sentiment analysis, Wikipedia article classification, and question-and-answer categorization English 5 domains
RVAE_LAMOL (Wang et al., 2022a) 15,870 - 5,668 21,538 DIL Oriented dialogue of the restaurant reservation task, semantic role labeling, sentiment classification English 3 domains
COPR (Zhang et al., 2023b) - - - - DIL - English 18 domains
SGD (Zhu et al., 2022) 38,745 5,210 11,349 40,287 DIL Dialogue state tracking English 19 Domains
CPT (Ke et al., 2022) 3,121,926 - - 3,121,926 DIL Domain-adaptive pre-training task English 4 Domains
CKL(Jang et al., 2022b) - - - 30,372 DIL Domain-adaptive pre-training task English 3 Domains
ELLE (Qin et al., 2022) - - - - DIL Domain-adaptive pre-training task English 5 Domains
Domain-incremental Paper Stream (Jin et al., 2022) - - - - DIL Relation extraction and named entity recognition English 4 domains
Chronologically-ordered Tweet Stream (Jin et al., 2022) - - - - DIL multi-label hashtag prediction and single-label emoji prediction English 4 domains
AdapterCL (Madotto et al., 2020) 31,426 4,043 4,818 40,287 DIL Intent classification, Dialogue State Tracking (DST), Natural Language Generation (NLG), end-to-end (E2E) modeling English 37 Domains
DE&E (Wójcik et al., 2023) 28,982 - 12,089 41,071 CIL Text classification English 3 tasks
EPI (Wang et al., 2023c) 12,840 3,524 6,917 23,281 CIL Text classification, topic classification English 13 Classes
PAGeR (Varshney et al., 2022) 59,754 7,115 15,304 82,173 CIL Intent classification English 355 Classes
PLE (Li et al., 2022b) 4,669 4,650 31,642 40,961 CIL Intent classification English 477 Classes
CoNLL-03 (Monaikul et al., 2021) 23,326 5,902 5,613 34,841 CIL Named Entity Recognition English 4 Classes
OntoNotes (Monaikul et al., 2021) 107,169 16,815 10,464 134,448 CIL Named Entity Recognition English 6 Classes
Online
Foundational text classification (de Masson D’Autume et al., 2019) 115,000 - 7,600 122,600 Hard and Blurry News classification, sentiment analysis, Wikipedia article classification, and question-and-answer categorization English 5 tasks
MBPA++ (de Masson D’Autume et al., 2019) 881,000 35,000 38,000 954,000 Hard and Blurry News classification, sentiment analysis, Wikipedia article classification, questions and answers categorization, question answering English 9 tasks
Lifelong FewRel (Holla et al., 2020) - - - - Hard and Blurry few-shot relation detection English 10 tasks
Firehose (Hu et al., 2022) - - - 110,000,000 Blurry Personalized online language learning English 1 tasks
TemporalWiki (Jang et al., 2022a) - - - - - - English -
6. Datasets
6.1. Offline Datasets for NLP
6.1.1. Datasets for Classification.
Text Classification.
The most typical task for continual learning is text classification. The foundational text classification benchmark encompasses five text classification datasets introduced by (Zhang et al., 2015), including AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers (Sun et al., 2020a). Particularly, the AG News dataset has 4 classes for news classification; the Amazon and Yelp dataset has 5 classes for sentiment analysis; the DBpedia dataset has 14 classes for Wikipedia text classification; and the Yahoo dataset has 10 classes for Q&A classification. The text classification benchmark includes 115,000 training and 7,600 test examples for each task, holding out 500 samples per class from the training set for validation. Building upon this, Razdaibiedina et al. (Razdaibiedina et al., 2023) developed a novel continual learning (CL) benchmark. This benchmark not only utilizes the foundational text classification benchmark but also integrates additional datasets from the GLUE benchmark (Wang et al., 2018), SuperGLUE benchmark (Wang et al., 2019a), and the IMDB dataset (Maas et al., 2011). Specifically, the GLUE benchmark datasets included are MNLI, QQP, RTE, and SST2, focusing on tasks like natural language inference, paraphrase detection, and sentiment analysis. Similarly, the SuperGLUE datasets—WiC, CB, COPA, MultiRC, and BoolQ—encompass tasks ranging from word sense disambiguation to question answering. DE&E (Wójcik et al., 2023) uses three common text classification data sets with different characteristics-News-groups, BBC News, and Consumer Finance Complaints2. Such datasets can be used to evaluate the models on tasks with different difficulty levels.
The datasets introduced in (Wang et al., 2023c) further are categorized into two groups based on the domain relevance between tasks: far-domain and near-domain. The far-domain group comprises two text classification tasks, which are foundational benchmarks (Zhang et al., 2015) divided into topic classification (AG News, Yahoo Answers, DBpedia) and sentiment classification (Yelp, Amazon Reviews). In contrast, the near-domain group uses the Web of Science (WOS) (Kowsari et al., 2017) and 20 Newsgroups (Lang, 1995), which are restructured according to their high inter-task relevance. The WOS dataset comprises seven parent classes, each with five closely related sub-classes, while the 20 Newsgroups dataset, containing six news topics, is reorganized into four tasks to maximize inter-task correlation.
Intent Classification.
Some studies focus on intent classification tasks, where the classes are quite different in various domains or scenarios. In the realm of intent classification and detection, several datasets have been specifically designed to advance the field by addressing different challenges and providing diverse environments for model training and evaluation. The dataset, as introduced in PAGeR (Varshney et al., 2022), aims to tackle the lifelong intent detection problem by combining three public intent classification datasets (CLINC150 (Larson et al., 2019), HWU64 (Liu et al., 2021a), BANKING77 (Casanueva et al., 2020)), one text classification dataset (Stackoverflow S20 (Xu et al., 2017)), and two public multidomain dialog intent detection datasets (SGD (Rastogi et al., 2020), MWOZ (Budzianowski et al., 2018)). Moreover, FewRel (Han et al., 2018) is also incorporated to tackle the lifelong relation extraction problem. This integration is intended to simulate real-world applications by encompassing a broad spectrum of domains and query distributions, thereby facilitating the development of more robust and versatile intent detection systems.
Conversely, the dataset compiled in PLE (Li et al., 2022b) consolidates nine well-regarded intent detection datasets, including CLINC150 (Larson et al., 2019) and HWU64 (Liu et al., 2021a), among others, arranged in a fixed random sequence to form a standardized benchmark. This dataset emphasizes the importance of consistency and comparability in performance evaluations across different intent detection models, providing a platform for assessing and enhancing various methodologies. The dataset described by MeLL (Wang et al., 2021a) specifically addresses intent detection within two distinct contexts: task-oriented dialogues (TaskDialog-EUIC) and real-world e-commerce interactions (Hotline-EUIC). TaskDialog-EUIC integrates data from Snips (Coucke et al., 2018), TOP semantic parsing (Gupta et al., 2018), and Facebook’s Multilingual Task Oriented Dataset (Schuster et al., 2018) into 90 tasks with overlapping label sets, amounting to over ten thousand samples. Hotline-EUIC is derived from an e-commerce dialogue system (Li et al., 2017) and the hotline audios are transcribed to text by a high-accuracy industrial Automatic Speech Recognition (ASR) system.
Fine-grained Sentiment Analysis.
Ke et al. (Ke and Xu, 2021) developed a task incremental learning dataset for aspect-based sentiment classification (ABSC). This dataset aggregates reviews from four distinct sources, thereby enhancing its diversity and applicability across multiple domains. The sources include the L5Domains dataset by Hu et al. (Hu and Liu, 2004), which features consumer reviews for five different products; the Liu3Domains dataset by Liu (Liu et al., 2015), comprising reviews pertaining to three products; the Ding9Domains dataset by Ding et al. (Ding et al., 2008), which includes reviews of nine varied products; and the SemEval14 dataset, which is focused on reviews of two specific products—laptops and restaurants.
6.1.2. Datasets for Generation.
In the rapidly advancing field of machine learning, diverse datasets function as crucial benchmarks for exploring various dimensions of language and code generation. These datasets address both universal and task-specific challenges, enabling a comprehensive evaluation of model capabilities. A particularly significant dataset highlighted in the work by Continual-T0 (Scialom et al., 2022b) focuses on English language generation tasks, including text simplification and empathetic dialogue generation, among others (Camburu et al., 2018; Bin Tareaf, 2017). The design of this dataset maintains uniformity in size, facilitating effective comparative analyses of performance across distinct tasks by ensuring a consistent volume of data for training. In a subsequent study, Luo et al. (Luo et al., 2023) conduct an analysis of catastrophic forgetting on Bloomz (Scao et al., 2022) using Continual T0 datasets.
The dataset, introduced in LAMOL (Sun et al., 2020a), integrates elements from both DecaNLP (McCann et al., [n. d.]) and the foundational text classification benchmark (Zhang et al., 2015). This dataset encompasses five distinct NLP tasks originally sourced from DecaNLP: question answering, semantic parsing, sentiment analysis, semantic role labeling, and goal-oriented dialogue. For the purposes of this dataset, all tasks, whether derived from DecaNLP or the foundational text classification benchmark, are restructured into a uniform format, conceptualized under the framework of a question answering task. Moreover, the dataset devised in RVAE_LAMOL (Wang et al., 2022a), employs three tasks from DecaNLP: the English Wizard of Oz (WOZ) for goal-oriented dialogue, QA-SRL for semantic role labeling in a SQuAD-style format, and SST, which is a binary version of the Stanford Sentiment Treebank categorizing sentiments as positive or negative. These tasks are specifically treated as sequence generation tasks.
The dataset introduced in COPR (Zhang et al., 2023b) represents a pioneering effort in applying both Task Incremental Learning (TIL) and Domain Incremental Learning (DIL) within the context of benchmarks that utilize existing human preferences. Specifically, the TIL framework in this dataset mandates that the model sequentially acquires knowledge from three distinct tasks. These include the question-answering task utilizing the HH-RLHF dataset (Bai et al., 2022), the summarization task based on the Reddit TL, DR dataset with human feedback (Volske et al., 2017), and the positive film review generation task using the IMDB dataset (Maas et al., 2011). Meanwhile, the DIL framework requires the model to adapt to three distinct segments from the SHP dataset, as described by Ethayarajh et al. (Ethayarajh et al., 2022).
The dataset described in Adaptive Compositional Modules (Zhang et al., 2022b) explores sequence generation and categorizes tasks into ”similar” and ”dissimilar” groups based on their characteristics. Tasks classified as similar, including E2ENLG (Novikova et al., 2017) and four domains (restaurant, hotel, TV, laptop) from RNNLG (Wen et al., 2015), demonstrate shared patterns and are tested across four sequence orders, comprising a total of five tasks. In contrast, dissimilar tasks such as WikiSQL (SQL query generation) (Zhong et al., 2017), CNN/DailyMail (news article summarization) (See et al., 2017), and MultiWOZ (semantic state sequence generation) (Budzianowski et al., 2018) exhibit significant distributional shifts from previously encountered tasks. The CODETASKCL dataset, explored by Yadav et al. (Yadav et al., 2023), encompasses a diverse array of code-centric tasks, including code generation (Iyer et al., 2018), summarization (Husain et al., 2019), translation (Lu et al., 2021), and refinement (Tufano et al., 2019) across various programming languages. This dataset significantly enhances the breadth of language processing applications within technical fields.
6.1.3. Datasets for Information Extraction.
In the realm of natural language processing (NLP), various datasets are tailored to specific aspects of the task under continual learning paradigms. The dataset introduced in ExtendNER (Monaikul et al., 2021), exemplifies a continual learning approach to Named Entity Recognition (NER). This dataset amalgamates the CoNLL-03 English NER (Sang and De Meulder, 2003) and OntoNotes (Hovy et al., 2006), covering a broad spectrum of entity types and sources. This hybrid dataset is structured to challenge the adaptability and generalization capabilities of NER systems across varied contexts. Unlike the static nature of text in NER tasks, the Schema-Guided Dialog (SGD) (Rastogi et al., 2020) dataset, utilized in C-PT (Zhu et al., 2022), serves the Dialog State Tracking aspect of IE, which involves maintaining the context of a dialog over time. The SGD dataset features 44 services across 19 domains, each treated as a separate task, and is designed to evaluate models on their ability to manage and extract information across conversational turns. Lastly, the lifelong SimpleQuestions and lifelong FewRel datasets, devised in (Wang et al., 2019b) is crafted for the task of relation extraction. It merges elements from the SimpleQuestions (Bordes et al., 2015) and FewRel (Han et al., 2018) to form a lifelong learning benchmark that confronts the challenges of relation detection in a few-shot context.
6.1.4. Datasets for Continual Pre-training.
In the realm of continual pre-training for large language models (LMs), the development and utilization of specialized benchmarks play a pivotal role in evaluating and enhancing the effectiveness of continual learning systems. The dataset, introduced in CPT (Ke et al., 2022), primarily focuses on the continual post-training of LMs across a series of domain-specific, unlabeled datasets. It provides a rigorous test environment by using diverse corpora such as Yelp Restaurant Reviews (Xu et al., 2019), AI and ACL Papers (Lo et al., 2020), and AGNews articles (Zhang et al., 2015). Its main objective is to gauge how well an LM can incrementally integrate domain-specific knowledge without forgetting previously learned information, thereby enhancing its few-shot learning capabilities in these domains. Contrary to the datasets employed in CPT (Ke et al., 2022), which evaluate domain-specific adaptability and incremental learning, the CKL benchmark (Jang et al., 2022b) is meticulously designed to measure the LM’s ability to retain timeless knowledge, update obsolete information, and acquire new knowledge. It comprises subsets like INVARIANTLAMA, UPDATEDLAMA, and NEWLAMA, which are crafted to probe specific types of knowledge that an LM may encounter in its learning trajectory.
Whereas the aforementioned two datasets assess more controlled dimensions of knowledge integration and retention, the dataset introduced in ELLE (Qin et al., 2022) focuses on the dynamic scenario of accumulating streaming data from diverse sources in a lifelong learning context. This dataset mirrors the real-world challenge of a language model (LM) that must continuously adapt to new data inflows from multiple domains, including BOOKCORPUS (WB) (Zhu et al., 2015), NEWS ARTICLES (NS) (Zellers et al., 2019), AMAZON REVIEWS (REV) (He and McAuley, 2016), BIOMEDICAL PAPERS (BIO) (Lo et al., 2020) and COMPUTER SCIENCE PAPERS (CS) (Lo et al., 2020). The benchmark evaluates the LM’s capacity to effectively integrate new information from these varied sources over time, highlighting the essential need for LMs to evolve in response to continual data growth and shifts in data distribution. Jin et al. (Jin et al., 2022) construct data streams to represent two prevalent types of domain shifts observed in practical scenarios. The first, a Domain-incremental Paper Stream, simulates the sequential evolution of research areas within academic papers, encompassing diverse disciplines such as biomedical and computer science. The second, a Chronologically-ordered Tweet Stream, models the temporal progression of tweets over time.
6.1.5. Datasets for Hybrid Tasks.
An increasing number of datasets are adopting a hybrid task approach that integrates multiple learning paradigms and task types, aimed at testing and enhancing the adaptability of models. A notable example is the dataset introduced in AdapterCL (Madotto et al., 2020), which is tailored for task-oriented dialogue systems. This dataset incorporates four task-oriented datasets: TaskMaster 2019 (TM19) (Byrne et al., 2019), TaskMaster 2020 (TM20) (Byrne et al., 2019), Schema Guided Dialogue (SGD) (Rastogi et al., 2020), and MultiWoZ (Budzianowski et al., 2018). These datasets have been pre-processed to form a curriculum encompassing 37 domains, structured under four continual learning settings: INTENT classification, Dialogue State Tracking (DST), Natural Language Generation (NLG), and end-to-end (E2E) modeling.
Continual Instruction Tuning Benchmark (CITB) (Zhang et al., 2023a) extends the concept of continual learning by focusing on instruction-based NLP tasks. Built on the comprehensive SuperNI (Wang et al., 2022c) dataset, it includes over 1,600 tasks across diverse NLP categories. CITB differentiates itself by formulating two distinct streams—InstrDialog and InstrDialog++—to examine how models integrate and retain new dialogue-oriented and varied NLP tasks under continual learning settings. This benchmark suite not only tests task retention and adaptability but also explores how instruction tuning can be optimized for a continual learning framework. The ConTinTin (Yin et al., 2022) is an adaptation of the NATURAL-INSTRUCTIONS dataset, specifically restructured to facilitate a continual learning framework. This adaptation involves decomposing the original crowdsourcing instructions into smaller, distinct sub-tasks to create a new dataset. Additionally, the new dataset incorporates a novel experimental design where tasks are selected randomly to create diverse sequences, enabling the evaluation of a model’s adaptability to novel instructions without prior exposure.
The dataset, used in Conure (Yuan et al., 2021), consists of Tencent TL (TTL) (Yuan et al., 2020) and Movielens (ML). The TTL dataset is designed to address three item recommendation tasks and three user profiling tasks, whereas the ML dataset exclusively focuses on three item recommendation tasks. Both datasets have been pre-processed to facilitate a continual learning framework, simulating environments where models must adapt to evolving data streams. Furthermore, Kim et al. (Kim et al., 2023) introduced the proprietary NAVER Shopping dataset, which builds upon the previously mentioned datasets. The NAVER Shopping dataset features six tasks: two for search query prediction, two for purchased item category prediction, and two for user profiling, all designed to meet real-world industry requirements.
Finally, the TRACE dataset, introduced by Wang et al. (Wang et al., 2024a), is specifically designed to bridge the existing gap in the evaluation of large language models (LLMs) within the continual learning framework, encompassing a wide range of complex and specialized tasks. Distinguished from other datasets, TRACE targets domain-specific tasks that are multilingual and technical, including code completion and mathematical reasoning. This diversity presents a unique set of challenges that span both specialized and broad dimensions. Moreover, TRACE rigorously assesses the models’ ability to sustain performance across tasks that demand different knowledge bases and cognitive skills. This evaluation highlights the essential need for adaptability in LLMs trained to operate under continual learning conditions, underscoring their potential in dynamic real-world applications.
6.2. Online Datasets for NLP
6.2.1. Datasets for Classification.
The foundational text classification benchmark, as introduced by Zhang et al. (Zhang et al., 2015) has traditionally been applied in offline continual learning settings. Recent advancements have adapted this benchmark for online continual learning, notably in studies such as MBPA++ (de Masson D’Autume et al., 2019) and OML-ER (Holla et al., 2020)
6.2.2. Datasets for Generation.
The dataset, used in MBPA++ (de Masson D’Autume et al., 2019), comprises three distinct question-answering collections: SQuAD 1.1 (Rajpurkar et al., 2016), TriviaQA (Joshi et al., 2017), and QuAC (Choi et al., 2018). SQuAD 1.1 is a reading comprehension dataset based on Wikipedia articles, designed to assess the ability to derive answers from structured text. TriviaQA consists of question-answer pairs developed by trivia enthusiasts, accompanied by corroborative evidence sourced from both the web and Wikipedia, testing the model’s capability to handle diverse information sources. QuAC adopts a dialog-style format in which a student queries about information in a Wikipedia article and a teacher responds using text directly from the article, challenging the model’s interactive response generation.
6.2.3. Datasets for Information Extraction.
The lifelong relation extraction benchmark, used in OML-ER (Holla et al., 2020), is structured by Wang et al. (Wang et al., 2019b) based on FewRel. Unlike the original application by Wang et al., the benchmark in OML-ER is adapted for online continuous learning scenarios.
6.2.4. Datasets for Other Tasks.
Hu et al. (Hu et al., 2022) compile the Firehose dataset, consisting of 110 million tweets from over 920,000 users between January 2013 and September 2019. This dataset is split into FIREHOSE 10M and FIREHOSE 100M. TemporalWiki (Jang et al., 2022a) addresses temporal misalignment by serving as a lifelong benchmark that trains and evaluates LMs using consecutive snapshots of Wikipedia and Wikidata. This methodology assists in assessing an LM’s capacity to both retain previously acquired knowledge and assimilate new information over time.
6.3. Offline CL Datasets for Multi-modal Tasks
The P9D dataset (Zhu et al., 2023) consists of over one million image-text pairs from e-commerce data, organized into nine industry sector-based training tasks. It includes 1,014,599 training pairs, 2,846 for cross-modal retrieval tests, and 4,615 query pairs with 46,855 gallery pairs for multi-modal retrieval. Qian et al. (Qian et al., 2023) introduce two novel benchmarks for continual learning, namely CL-TDIUC and CL-VQA2.0, which are derived from the TDIUC (Kafle and Kanan, 2017) and VQA2.0 (Goyal et al., 2017), respectively. These benchmarks are categorized into three scenarios: the Continual Vision Scenario, which deals with new visual scenes; the Continual Language Scenario, focusing on new questions in existing scenes; and the Continual Vision-Language Scenario, addressing changes in both questions and visuals. DKR (Cui et al., 2024) comprises five benchmark datasets: MS-COCO Caption (MS-COCO) (Lin et al., 2014), Flickr30K (Young et al., 2014), IAPR TC-12 (Grubinger et al., 2006), ECommerce-T2I (EC) (Yang et al., 2021), and RSICD (Lu et al., 2017). Furthermore, two experimental scenarios are established. The first scenario involves a sequential processing of the datasets, specifically MS-COCO, Flickr30K, IAPR TC-12, EC, and RSICD, in that order. The second scenario, which builds on the approach proposed by Ni et al. (Ni et al., 2023), partitions the EC dataset into five sub-datasets for the training phase. The model’s performance is subsequently tested on the Flickr30K, MS-COCO, and EC datasets.
6.4. Offline and Online Datasets for Other Tasks
In this paper, we provide a comprehensive review of recent studies in natural language processing (NLP) and multi-modal tasks, with a particular focus on the use of PLMs, LLMs and VLMs. Additionally, we introduce a range of offline continual learning (CL) datasets employed across various applications, including automatic speech recognition (Vander Eeckt et al., 2023; Della Libera et al., 2023), autonomous driving (Verwimp et al., 2023), disease classification (Derakhshani et al., 2022), reinforcement learning (Wołczyk et al., 2021), graph data (Ko et al., 2022), and computer vision (Faber et al., 2023; Lin et al., 2021). Furthermore, some online CL benchmarks are also proposed for computer version tasks, such as continual visual learning (Cai et al., 2021), continual object detection (Wang et al., 2021b).
Figure 9. Illustration of calculating metrics.
7. Metrics
In this section, we review the principal metrics commonly used to evaluate continual learning. These metrics can be categorized into three main types: (1) overall performance, which assesses the algorithm’s effectiveness across all tasks; (2) memory stability, which measures the extent to which an algorithm retains previously acquired knowledge; and (3) learning plasticity, which evaluates the algorithm’s capacity to acquire new skills or knowledge. Each of these metrics provides insights into different aspects of the algorithm’s performance in a continual learning context.
To begin, we establish the notation (Figure 9) used throughout the learning and evaluation phases of the model. Once the model completes a learning task, denoted as $T_{i}$, it evaluates its performance on a test set that encompasses all $N$ tasks, where $N$ is the total number of tasks in the set $T$. This evaluation is represented by a matrix $R\in\mathbb{R}^{N\times N}$, wherein each element $R_{i,j}$ indicates the model’s test classification accuracy on task $T_{j}$ after training on task $T_{i}$.
7.1. Overall Performance.
The metric termed “Last” (Lopez-Paz and Ranzato, 2017; Zheng et al., 2023) evaluates the overall performance of a continual learning (CL) method upon the completion of all tasks. Specifically, it computes the average score from the last row in the performance matrix $R$.
(1) $Last=\frac{1}{N}\sum_{i=1}^{N}R_{N,i}$
Also, Zheng et al. (Zheng et al., 2023) devise the “Avg” score metric, which computes the mean accuracy across all datasets and timestamps.
(2) $Avg=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{N}\sum_{j=1}^{N}R_{i,j}\right)$
In the seminal works of Rebuffi et al. (Rebuffi et al., 2017) and Douillard et al. (Douillard et al., 2020), the concept of Average Incremental Accuracy (AIA) is introduced. This metric is specifically designed to quantify the historical performance across different tasks. It calculates the average performance for each task by considering the lower triangular portion of the matrix $R$, effectively capturing the evolving competence of the system as new tasks are learned.
(3) $AIA=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{i}\sum_{j=1}^{i}R_{i,j}\right)$
The metric, termed Transfer, is derived by computing the average of the performance values for tasks that are represented in the upper-right triangle of matrix $R$. This approach uniformly weights each dataset by averaging their performance across different tasks, thereby assessing the preservation of zero-shot transfer capabilities. Prior to commencing learning on task $T_{i}$, no fine-tuning is performed on tasks that precede $T_{i}$.
(4) $Transfer=\frac{1}{N-1}\sum_{i=2}^{N}\left(\frac{1}{i-1}\sum_{j=1}^{i-1}R_{j,i}\right)$
Moreover, Chaudhry et al. (Chaudhry et al., 2019) devise a metric known as Learning Curve Area (LCA), which quantifies the speed of learning in a model. Qin et al. (Qin et al., 2022) propose two metrics designed to evaluate pre-trained language models (PLMs) based on their performance within learned domains: Average Perplexity ($AP$) and Average Increased Perplexity ($AP^{+}$).
7.2. Memory Stability.
Memory stability is typically evaluated by backward transfer (BWT) (Lopez-Paz and Ranzato, 2017) and forgetting measure (FM) (Chaudhry et al., 2018).
Backward Transfer (BWT) emerges as a pivotal concept extensively documented in the literature, notably by Lopez et al. (Lopez-Paz and Ranzato, 2017) and Wu et al. (Wu et al., 2023). BWT measures the performance degradation on previously mastered tasks after the model is trained on new tasks. This performance degradation phenomenon is often referred to as “forgetting”.
(5) $BWT=\frac{1}{N-1}\sum_{i=1}^{N-1}R_{N,i}-R_{i,i}$
Additionally, Chaudhry et al. (Chaudhry et al., 2018) introduce the Forgetting Measure (FM), a metric designed to quantify the extent of forgetting a model experiences for a specific task. A lower FM indicates better retention of previous tasks. Davari et al. (Davari et al., 2022) propose a method named linear probes (LP) to assess representation forgetting. This approach measures the effectiveness of learned representations via an optimal linear classifier trained on the frozen activations of a base network. Representation forgetting is quantified by evaluating the change in Language Processing (LP) performance before and after the introduction of a new task. Kemker et al. (Kemker et al., 2018) introduce three metrics, where $\Omega_{\text{base}}$ assesses retention of initial learning, $\Omega_{\text{new}}$ measures recall of new tasks, and $\Omega_{\text{all}}$ evaluates overall proficiency in maintaining old knowledge and acquiring new information. Additionally, researchers (Koh et al., 2022) devise a novel metric, termed the Knowledge Loss Ratio (KLR), quantifies knowledge degradation using principles from information theory.
7.3. Learning Plasticity.
Evaluating learning plasticity can be effectively accomplished through two key metrics: forward transfer (FWT) (Lopez-Paz and Ranzato, 2017) and intransigence measure (IM) (Chaudhry et al., 2018).
Forward Transfer (FWT) (Lopez-Paz and Ranzato, 2017) assesses the beneficial effects on the performance of subsequent tasks following a model’s training on prior tasks.
(6) $FWT=\frac{1}{N-1}\sum_{i=2}^{N}R_{i-1,i}-R_{0,i}$
where $R_{0,i}$ denotes the performance metric associated with training on task $i$ independently. Higher values of FWT indicate superior model performance. It is important to note that discussing backward transfer for the initial task is not applicable, as there are no preceding tasks to influence its performance.
Intransigence measure (IM), as defined by Chaudhry et al. (Chaudhry et al., 2018), quantifies a model’s inability to learn new tasks. This measure is calculated by comparing the performance difference of a task when trained jointly with other tasks versus when trained in a continual learning setting. Moreover, Koh et al. (Koh et al., 2022) introduce novel metrics, known as Knowledge Gain Ratio (KGR), which quantifies the capacity to acquire new knowledge through the calculation of knowledge gain.
7.4. Metrics for Continual Pre-training.
CKL (Jang et al., 2022b) introduces a novel metric, named FUAR (FORGOTTEN / (UPDATED + ACQUIRED) RATIO), which quantitatively measures the efficiency of each CKL method. It calculates the number of instances of time-invariant knowledge that a model forgets in order to learn or update one instance of new knowledge. When FUAR is equal to 1.0, it signifies an equilibrium where one time-invariant knowledge instance is forgotten on average to obtain a new or updated knowledge instance.
7.5. Online CL-Specific Metrics.
Near-future accuracy (NFA) (Al Kader Hammoud et al., 2023) is introduced as a novel evaluation metric for OCL problem. Unlike traditional evaluation methods that assess models on immediately subsequent samples, NFA evaluates models on samples slightly further into the future, using a minimal shift $S$. Such operation can mitigate label correlation effects, which can adversely impact the accuracy of model adaptability assessments. The smallest shift $S$ is selected to ensure that the test sample aligns closely with the distribution of recently observed training data. Yogatama et al. (Yogatama et al., 2019) proposed a novel online codelength, inspired by prequential encoding (Blier and Ollivier, 2018), to quantify how quickly an existing model can adapt to a new task.
8. Challenges and Further Work
Autonomous Continual Learning.
Most existing studies in the domain of continual learning assume static datasets with known distributions in a relatively closed environment. Moreover, these studies mainly focus on simple tasks (e.g., text classification, sentiment analysis and intent classification) with clear labels. These assumptions do not hold in real-world applications, where environments continually evolve and introduce novel stimuli. A key challenge is to develop continual learning models that operate effectively in complex, noisy environments where clear labels are not always available, and task domains frequently change. Liu et al. (Liu et al., 2023b) recently proposed the SOLA framework to address these limitations by facilitating autonomous adaptation in AI systems. Despite this progress, significant challenges remain in enabling these systems to independently adjust to new, dynamic environments without ongoing human oversight. Future research should focus on developing algorithms capable of autonomously detecting and adapting to shifts in data distribution, thereby improving the applicability of AI in dynamic real-world scenarios.
Learning Knowledge from Conversation.
Traditional AI systems are typically trained on static data sets, which starkly contrasts with human conversational learning that dynamically updates knowledge through interaction (Liu and Mazumder, 2021). The challenge for AI lies in transitioning from static data learning to more dynamic, conversational engagements. The future direction in this area could involve the development of models that mimic human conversational learning processes, capable of context adaptation, new concept inference, and dynamic knowledge application within ongoing interactions.
Multi-modal Continual Learning.
Continual learning research has predominantly concentrated on natural language processing tasks such as sentiment analysis and text classification. Recent studies have begun exploring basic multi-modal tasks, such as text-to-image retrieval, text-image classification, and visual question answering. The integration of diverse data types—textual, visual, and auditory—poses a substantial challenge. Future studies should expand to more complex multi-modal datasets and strive to devise methodologies that effectively synthesize these varied modalities, thereby enhancing the model’s capability to maintain continuous learning across different sensory inputs.
Privacy Protection in Continual Learning.
Privacy protection in continual learning systems poses a significant challenge, particularly as these systems are designed to continuously update and refine their models based on incoming data streams. Unlike traditional static machine learning models, continual learning systems frequently access and process sensitive data across different contexts and time periods, raising substantial concerns about data confidentiality and user privacy. Effective privacy-preserving mechanisms must be integrated into the architecture of these systems to ensure that they do not inadvertently expose or misuse personal data. Techniques such as differential privacy (Dwork, 2006), federated learning (Zhang et al., 2021), and secure multi-party computation (Goldreich, 1998) offer promising solutions by allowing models to learn from decentralized data sources without needing to access the actual data directly. Future research in continual learning should not only focus on enhancing learning efficiency and adaptability but also prioritize the development of robust frameworks that safeguard user privacy across all phases of data handling and model updating.
Robust Continual Learning.
The existing studies mainly focus on designing a continual learning model to improve the performance of forgetting and transferring with various metrics while the robustness of continual learning systems is not well studied. It is critical, especially in applications where safety and reliability are paramount. The main challenges include evaluating the robustness of these systems against adversarial attacks or when faced with drastically changing environments. Future research could focus on developing evaluation metrics for robustness in continual learning and designing systems that maintain performance reliability over time despite environmental changes.
Large-Scale and High-Quality Datasets and Benchmarks.
As discussed in Section 6, most of the datasets are constructed by merging the existing datasets. This often results in datasets that lack diversity and real-world complexity, which hampers the development of robust and adaptable continual learning models. The creation of large-scale, high-quality datasets that accurately reflect real-world complexities represents a critical challenge. Moving forward, the development of such datasets and benchmarks will be essential not only for assessing the efficacy of continual learning algorithms but also for pushing the limits of what these algorithms can achieve in practical settings.
9. Conclusions
This survey provides an in-depth exploration of continual learning (CL) methodologies tailored for foundation language models (LMs), such as pre-trained language models (PLMs), large language models (LLMs), and vision-language models (VLMs). By integrating the dynamic adaptability of CL with the robust foundational capabilities of LMs, this field promises to significantly advance the state of artificial intelligence. We categorize existing research into offline and online continual learning paradigms, offering a clear distinction between the settings and methodologies used within these frameworks. Offline CL is discussed in terms of domain-incremental, task-incremental, and class-incremental learning. Meanwhile, online CL is analyzed with a focus on the delineation between hard and blurry task boundaries, providing insights into how these approaches handle real-time data streams. Our review of the literature not only clarifies the current landscape of CL approaches for foundation LMs but also emphasizes the innovative integration of continual pre-training, parameter-efficient tuning, and instruction tuning methods that are specifically designed to leverage the vast capabilities of foundation LMs. Furthermore, we highlight the main characteristics of datasets used in this domain and the metrics that effectively measure both the mitigation of catastrophic forgetting and the enhancement of knowledge transfer. This work hopes to inspire further research that will ultimately lead to more robust, efficient, and intelligent systems capable of lifelong learning.
  • (1)
  • Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al. 2023. Gpt-4 technical report. arXiv preprint arXiv:2303.08774 (2023).
  • Al Kader Hammoud et al. (2023) Hasan Abed Al Kader Hammoud, Ameya Prabhu, Ser-Nam Lim, Philip HS Torr, Adel Bibi, and Bernard Ghanem. 2023. Rapid Adaptation in Online Continual Learning: Are We Evaluating It Right?. In ICCV. 18852–18861.
  • Aljundi et al. (2019) Rahaf Aljundi, Eugene Belilovsky, Tinne Tuytelaars, Laurent Charlin, Massimo Caccia, Min Lin, and Lucas Page-Caccia. 2019. Online Continual Learning with Maximal Interfered Retrieval. In NeurIPS, Vol. 32.
  • Bai et al. (2022) Yuntao Bai, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, Nova DasSarma, Dawn Drain, Stanislav Fort, Deep Ganguli, Tom Henighan, et al. 2022. Training a helpful and harmless assistant with reinforcement learning from human feedback. arXiv preprint arXiv:2204.05862 (2022).
  • Bang et al. (2022) Jihwan Bang, Hyunseo Koh, Seulki Park, Hwanjun Song, Jung-Woo Ha, and Jonghyun Choi. 2022. Online continual learning on a contaminated data stream with blurry task boundaries. In CVPR. 9275–9284.
  • Beaulieu et al. (2020) Shawn Beaulieu, Lapo Frati, Thomas Miconi, Joel Lehman, Kenneth O Stanley, Jeff Clune, and Nick Cheney. 2020. Learning to continually learn. arXiv preprint arXiv:2002.09571 (2020).
  • Belouadah et al. (2021) Eden Belouadah, Adrian Popescu, and Ioannis Kanellos. 2021. A comprehensive study of class incremental learning algorithms for visual tasks. Neural Networks 135 (2021), 38–54.
  • Biesialska et al. (2020) Magdalena Biesialska, Katarzyna Biesialska, and Marta R Costa-jussà. 2020. Continual Lifelong Learning in Natural Language Processing: A Survey. In COLING. 6523–6541.
  • Bin Tareaf (2017) Raad Bin Tareaf. 2017. Tweets Dataset - Top 20 most followed users in Twitter social platform.
  • Blier and Ollivier (2018) Léonard Blier and Yann Ollivier. 2018. The description length of deep learning models. NeurIPS 31 (2018).
  • Bordes et al. (2015) Antoine Bordes, Nicolas Usunier, Sumit Chopra, and Jason Weston. 2015. Large-scale simple question answering with memory networks. arXiv preprint arXiv:1506.02075 (2015).
  • Budzianowski et al. (2018) Paweł Budzianowski, Tsung-Hsien Wen, Bo-Hsiang Tseng, Inigo Casanueva, Stefan Ultes, Osman Ramadan, and Milica Gašić. 2018. MultiWOZ–a large-scale multi-domain wizard-of-oz dataset for task-oriented dialogue modelling. arXiv preprint arXiv:1810.00278 (2018).
  • Byrne et al. (2019) Bill Byrne, Karthik Krishnamoorthi, Chinnadhurai Sankar, Arvind Neelakantan, Daniel Duckworth, Semih Yavuz, Ben Goodrich, Amit Dubey, Andy Cedilnik, and Kyu-Young Kim. 2019. Taskmaster-1: Toward a realistic and diverse dialog dataset. arXiv preprint arXiv:1909.05358 (2019).
  • Caccia et al. (2021) Lucas Caccia, Rahaf Aljundi, Nader Asadi, Tinne Tuytelaars, Joelle Pineau, and Eugene Belilovsky. 2021. New insights on reducing abrupt representation change in online continual learning. arXiv preprint arXiv:2104.05025 (2021).
  • Cahyawijaya et al. (2023) Samuel Cahyawijaya, Holy Lovenia, Tiezheng Yu, Willy Chung, and Pascale Fung. 2023. InstructAlign: High-and-Low Resource Language Alignment via Continual Crosslingual Instruction Tuning. In the First Workshop in South East Asian Language Processing. 55–78.
  • Cai et al. (2021) Zhipeng Cai, Ozan Sener, and Vladlen Koltun. 2021. Online continual learning with natural distribution shifts: An empirical study with visual data. In ICCV. 8281–8290.
  • Camburu et al. (2018) Oana-Maria Camburu, Tim Rocktäschel, Thomas Lukasiewicz, and Phil Blunsom. 2018. e-snli: Natural language inference with natural language explanations. NeurIPS 31 (2018).
  • Cao et al. (2024) Xusheng Cao, Haori Lu, Linlan Huang, Xialei Liu, and Ming-Ming Cheng. 2024. Generative Multi-modal Models are Good Class-Incremental Learners. arXiv preprint arXiv:2403.18383 (2024).
  • Casanueva et al. (2020) Iñigo Casanueva, Tadas Temčinas, Daniela Gerz, Matthew Henderson, and Ivan Vulić. 2020. Efficient intent detection with dual sentence encoders. arXiv preprint arXiv:2003.04807 (2020).
  • Castellucci et al. (2021) Giuseppe Castellucci, Simone Filice, Danilo Croce, and Roberto Basili. 2021. Learning to Solve NLP Tasks in an Incremental Number of Languages. In ACL. 837–847.
  • Chaudhry et al. (2018) Arslan Chaudhry, Puneet K Dokania, Thalaiyasingam Ajanthan, and Philip HS Torr. 2018. Riemannian walk for incremental learning: Understanding forgetting and intransigence. In ECCV. 532–547.
  • Chaudhry et al. (2019) Arslan Chaudhry, Marc’Aurelio Ranzato, Marcus Rohrbach, and Mohamed Elhoseiny. 2019. Efficient Lifelong Learning with A-GEM. In ICLR.
  • Chen et al. (2024) Cheng Chen, Junchen Zhu, Xu Luo, Hengtao Shen, Lianli Gao, and Jingkuan Song. 2024. CoIN: A Benchmark of Continual Instruction tuNing for Multimodel Large Language Model. arXiv preprint arXiv:2403.08350 (2024).
  • Chen et al. (2020) Hung-Jen Chen, An-Chieh Cheng, Da-Cheng Juan, Wei Wei, and Min Sun. 2020. Mitigating forgetting in online continual learning via instance-aware parameterization. NeurIPS 33 (2020), 17466–17477.
  • Chen et al. (2023) Wuyang Chen, Yanqi Zhou, Nan Du, Yanping Huang, James Laudon, Zhifeng Chen, and Claire Cui. 2023. Lifelong language pretraining with distribution-specialized experts. In ICML. PMLR, 5383–5395.
  • Cheng et al. (2023a) Daixuan Cheng, Shaohan Huang, and Furu Wei. 2023a. Adapting large language models via reading comprehension. arXiv preprint arXiv:2309.09530 (2023).
  • Cheng et al. (2023b) Xin Cheng, Yankai Lin, Xiuying Chen, Dongyan Zhao, and Rui Yan. 2023b. Decouple knowledge from paramters for plug-and-play language modeling. In ACL. 14288–14308.
  • Choi et al. (2018) Eunsol Choi, He He, Mohit Iyyer, Mark Yatskar, Wen-tau Yih, Yejin Choi, Percy Liang, and Luke Zettlemoyer. 2018. QuAC: Question Answering in Context. In EMNLP. 2174–2184.
  • Chrysakis and Moens (2023) Aristotelis Chrysakis and Marie-Francine Moens. 2023. Online bias correction for task-free continual learning. ICLR (2023).
  • Cossu et al. (2022) Andrea Cossu, Tinne Tuytelaars, Antonio Carta, Lucia Passaro, Vincenzo Lomonaco, and Davide Bacciu. 2022. Continual pre-training mitigates forgetting in language and vision. arXiv preprint arXiv:2205.09357 (2022).
  • Coucke et al. (2018) Alice Coucke, Alaa Saade, Adrien Ball, Théodore Bluche, Alexandre Caulier, David Leroy, Clément Doumouro, Thibault Gisselbrecht, Francesco Caltagirone, Thibaut Lavril, et al. 2018. Snips voice platform: an embedded spoken language understanding system for private-by-design voice interfaces. arXiv preprint arXiv:1805.10190 (2018).
  • Cui et al. (2024) Zhenyu Cui, Yuxin Peng, Xun Wang, Manyu Zhu, and Jiahuan Zhou. 2024. Continual Vision-Language Retrieval via Dynamic Knowledge Rectification. In AAAI, Vol. 38. 11704–11712.
  • Davari et al. (2022) MohammadReza Davari, Nader Asadi, Sudhir Mudur, Rahaf Aljundi, and Eugene Belilovsky. 2022. Probing representation forgetting in supervised and unsupervised continual learning. In CVPR. 16712–16721.
  • De Lange et al. (2021) Matthias De Lange, Rahaf Aljundi, Marc Masana, Sarah Parisot, Xu Jia, Aleš Leonardis, Gregory Slabaugh, and Tinne Tuytelaars. 2021. A continual learning survey: Defying forgetting in classification tasks. TPAMI 44, 7 (2021), 3366–3385.
  • de Masson D’Autume et al. (2019) Cyprien de Masson D’Autume, Sebastian Ruder, Lingpeng Kong, and Dani Yogatama. 2019. Episodic memory in lifelong language learning. NeurIPS 32 (2019).
  • Della Libera et al. (2023) Luca Della Libera, Pooneh Mousavi, Salah Zaiem, Cem Subakan, and Mirco Ravanelli. 2023. CL-MASR: A Continual Learning Benchmark for Multilingual ASR. arXiv preprint arXiv:2310.16931 (2023).
  • Derakhshani et al. (2022) Mohammad Mahdi Derakhshani, Ivona Najdenkoska, Tom van Sonsbeek, Xiantong Zhen, Dwarikanath Mahapatra, Marcel Worring, and Cees GM Snoek. 2022. Lifelonger: A benchmark for continual disease classification. In International Conference on Medical Image Computing and Computer-Assisted Intervention. 314–324.
  • Ding et al. (2008) Xiaowen Ding, Bing Liu, and Philip S Yu. 2008. A holistic lexicon-based approach to opinion mining. In WSDM. 231–240.
  • Dong et al. (2022) Qingxiu Dong, Lei Li, Damai Dai, Ce Zheng, Zhiyong Wu, Baobao Chang, Xu Sun, Jingjing Xu, and Zhifang Sui. 2022. A survey on in-context learning. arXiv preprint arXiv:2301.00234 (2022).
  • Douillard et al. (2020) Arthur Douillard, Matthieu Cord, Charles Ollion, Thomas Robert, and Eduardo Valle. 2020. Podnet: Pooled outputs distillation for small-tasks incremental learning. In ECCV. 86–102.
  • Du et al. (2023) Mingzhe Du, Anh Tuan Luu, Bin Ji, and See-kiong Ng. 2023. From Static to Dynamic: A Continual Learning Framework for Large Language Models. arXiv preprint arXiv:2310.14248 (2023).
  • Du et al. (2022) Yifan Du, Zikang Liu, Junyi Li, and Wayne Xin Zhao. 2022. A Survey of Vision-Language Pre-Trained Models. In IJCAI. 5436–5443.
  • Dwork (2006) Cynthia Dwork. 2006. Differential privacy. In International colloquium on automata, languages, and programming. 1–12.
  • Ethayarajh et al. (2022) Kawin Ethayarajh, Yejin Choi, and Swabha Swayamdipta. 2022. Understanding dataset difficulty with V-usable information. In ICML. 5988–6008.
  • Faber et al. (2023) Kamil Faber, Dominik Zurek, Marcin Pietron, Nathalie Japkowicz, Antonio Vergari, and Roberto Corizzo. 2023. From MNIST to ImageNet and Back: Benchmarking Continual Curriculum Learning. arXiv preprint arXiv:2303.11076 (2023).
  • Fini et al. (2020) Enrico Fini, Stéphane Lathuiliere, Enver Sangineto, Moin Nabi, and Elisa Ricci. 2020. Online continual learning under extreme memory constraints. In ECCV 2020. 720–735.
  • Geng et al. (2021) Binzong Geng, Fajie Yuan, Qiancheng Xu, Ying Shen, Ruifeng Xu, and Min Yang. 2021. Continual Learning for Task-oriented Dialogue System with Iterative Network Pruning, Expanding and Masking. In ACL. 517–523.
  • Gogoulou et al. (2023) Evangelia Gogoulou, Timothée Lesort, Magnus Boman, and Joakim Nivre. 2023. A study of continual learning under language shift. arXiv preprint arXiv:2311.01200 (2023).
  • Goldreich (1998) Oded Goldreich. 1998. Secure multi-party computation. Manuscript. Preliminary version 78, 110 (1998), 1–108.
  • Goyal et al. (2017) Yash Goyal, Tejas Khot, Douglas Summers-Stay, Dhruv Batra, and Devi Parikh. 2017. Making the v in vqa matter: Elevating the role of image understanding in visual question answering. In CVPR. 6904–6913.
  • Grubinger et al. (2006) Michael Grubinger, Paul Clough, Henning Müller, and Thomas Deselaers. 2006. The iapr tc-12 benchmark: A new evaluation resource for visual information systems. In International workshop ontoImage, Vol. 2.
  • Gu et al. (2022) Yanan Gu, Xu Yang, Kun Wei, and Cheng Deng. 2022. Not just selection, but exploration: Online class-incremental continual learning via dual view consistency. In CVPR. 7442–7451.
  • Gunasekara et al. (2023) Nuwan Gunasekara, Bernhard Pfahringer, Heitor Murilo Gomes, and Albert Bifet. 2023. Survey on Online Streaming Continual Learning. In IJCAI.
  • Guo et al. (2022) Yiduo Guo, Bing Liu, and Dongyan Zhao. 2022. Online continual learning through mutual information maximization. In ICML. 8109–8126.
  • Guo et al. (2023) Yiduo Guo, Bing Liu, and Dongyan Zhao. 2023. Dealing with Cross-Task Class Discrimination in Online Continual Learning. In CVPR. 11878–11887.
  • Gupta et al. (2018) Sonal Gupta, Rushin Shah, Mrinal Mohit, Anuj Kumar, and Mike Lewis. 2018. Semantic parsing for task oriented dialog using hierarchical representations. arXiv preprint arXiv:1810.07942 (2018).
  • Gururangan et al. (2022) Suchin Gururangan, Mike Lewis, Ari Holtzman, Noah A Smith, and Luke Zettlemoyer. 2022. DEMix Layers: Disentangling Domains for Modular Language Modeling. In NAACL. 5557–5576.
  • Han et al. (2018) Xu Han, Hao Zhu, Pengfei Yu, Ziyun Wang, Yuan Yao, Zhiyuan Liu, and Maosong Sun. 2018. FewRel: A large-scale supervised few-shot relation classification dataset with state-of-the-art evaluation. arXiv preprint arXiv:1810.10147 (2018).
  • Han et al. (2024) Zeyu Han, Chao Gao, Jinyang Liu, Sai Qian Zhang, et al. 2024. Parameter-Efficient Fine-Tuning for Large Models: A Comprehensive Survey. arXiv preprint arXiv:2403.14608 (2024).
  • Hayes et al. (2021) Tyler L Hayes, Giri P Krishnan, Maxim Bazhenov, Hava T Siegelmann, Terrence J Sejnowski, and Christopher Kanan. 2021. Replay in deep learning: Current approaches and missing biological elements. Neural computation 33, 11 (2021), 2908–2950.
  • He et al. (2020) Jiangpeng He, Runyu Mao, Zeman Shao, and Fengqing Zhu. 2020. Incremental learning in online scenario. In CVPR. 13926–13935.
  • He and McAuley (2016) Ruining He and Julian McAuley. 2016. Ups and downs: Modeling the visual evolution of fashion trends with one-class collaborative filtering. In WWW. 507–517.
  • Holla et al. (2020) Nithin Holla, Pushkar Mishra, Helen Yannakoudakis, and Ekaterina Shutova. 2020. Meta-learning with sparse experience replay for lifelong language learning. arXiv preprint arXiv:2009.04891 (2020).
  • Houlsby et al. (2019) Neil Houlsby, Andrei Giurgiu, Stanislaw Jastrzebski, Bruna Morrone, Quentin De Laroussilhe, Andrea Gesmundo, Mona Attariyan, and Sylvain Gelly. 2019. Parameter-efficient transfer learning for NLP. In ICML. 2790–2799.
  • Hovy et al. (2006) Eduard Hovy, Mitchell Marcus, Martha Palmer, Lance Ramshaw, and Ralph Weischedel. 2006. OntoNotes: The 90% Solution. In NAACL. 57–60.
  • Hu et al. (2022) Hexiang Hu, Ozan Sener, Fei Sha, and Vladlen Koltun. 2022. Drinking from a firehose: Continual learning with web-scale natural language. TPAMI 45, 5 (2022), 5684–5696.
  • Hu and Liu (2004) Minqing Hu and Bing Liu. 2004. Mining and summarizing customer reviews. In SIGKDD. 168–177.
  • Huang et al. (2021) Yufan Huang, Yanzhe Zhang, Jiaao Chen, Xuezhi Wang, and Diyi Yang. 2021. Continual Learning for Text Classification with Information Disentanglement Based Regularization. In ACL. 2736–2746.
  • Husain et al. (2019) Hamel Husain, Ho-Hsiang Wu, Tiferet Gazit, Miltiadis Allamanis, and Marc Brockschmidt. 2019. Codesearchnet challenge: Evaluating the state of semantic code search. arXiv preprint arXiv:1909.09436 (2019).
  • Iyer et al. (2018) Srinivasan Iyer, Ioannis Konstas, Alvin Cheung, and Luke Zettlemoyer. 2018. Mapping language to code in programmatic context. arXiv preprint arXiv:1808.09588 (2018).
  • Jang et al. (2023) Joel Jang, Seungone Kim, Seonghyeon Ye, Doyoung Kim, Lajanugen Logeswaran, Moontae Lee, Kyungjae Lee, and Minjoon Seo. 2023. Exploring the benefits of training expert language models over instruction tuning. In ICML. 14702–14729.
  • Jang et al. (2022a) Joel Jang, Seonghyeon Ye, Changho Lee, Sohee Yang, Joongbo Shin, Janghoon Han, Gyeonghun Kim, and Minjoon Seo. 2022a. TemporalWiki: A Lifelong Benchmark for Training and Evaluating Ever-Evolving Language Models. In EMNLP. 6237–6250.
  • Jang et al. (2022b) Joel Jang, Seonghyeon Ye, Sohee Yang, Joongbo Shin, Janghoon Han, Gyeonghun Kim, Stanley Jungkyu Choi, and Minjoon Seo. 2022b. Towards Continual Knowledge Learning of Language Models. In ICLR.
  • Javed and White (2019) Khurram Javed and Martha White. 2019. Meta-learning representations for continual learning. NeurIPS 32 (2019).
  • Jha et al. (2024) Saurav Jha, Dong Gong, and Lina Yao. 2024. CLAP4CLIP: Continual Learning with Probabilistic Finetuning for Vision-Language Models. arXiv preprint arXiv:2403.19137 (2024).
  • Jin et al. (2021a) Xisen Jin, Bill Yuchen Lin, Mohammad Rostami, and Xiang Ren. 2021a. Learn Continually, Generalize Rapidly: Lifelong Knowledge Accumulation for Few-shot Learning. In EMNLP. 714–729.
  • Jin et al. (2021b) Xisen Jin, Arka Sadhu, Junyi Du, and Xiang Ren. 2021b. Gradient-based editing of memory examples for online task-free continual learning. NeurIPS 34 (2021), 29193–29205.
  • Jin et al. (2022) Xisen Jin, Dejiao Zhang, Henghui Zhu, Wei Xiao, Shang-Wen Li, Xiaokai Wei, Andrew Arnold, and Xiang Ren. 2022. Lifelong Pretraining: Continually Adapting Language Models to Emerging Corpora. In NAACL. 4764–4780.
  • Joshi et al. (2017) Mandar Joshi, Eunsol Choi, Daniel Weld, and Luke Zettlemoyer. 2017. TriviaQA: A Large Scale Distantly Supervised Challenge Dataset for Reading Comprehension. In ACL. 1601–1611.
  • Kafle and Kanan (2017) Kushal Kafle and Christopher Kanan. 2017. An analysis of visual question answering algorithms. In ICCV. 1965–1973.
  • Ke et al. (2022) Zixuan Ke, Haowei Lin, Yijia Shao, Hu Xu, Lei Shu, and Bing Liu. 2022. Continual Training of Language Models for Few-Shot Learning. In EMNLP. 10205–10216.
  • Ke and Liu (2022) Zixuan Ke and Bing Liu. 2022. Continual learning of natural language processing tasks: A survey. arXiv preprint arXiv:2211.12701 (2022).
  • Ke et al. (2021a) Zixuan Ke, Bing Liu, Nianzu Ma, Hu Xu, and Lei Shu. 2021a. Achieving Forgetting Prevention and Knowledge Transfer in Continual Learning. In NeurIPS, Vol. 34. 22443–22456.
  • Ke et al. (2021b) Zixuan Ke, Bing Liu, Hu Xu, and Lei Shu. 2021b. CLASSIC: Continual and Contrastive Learning of Aspect Sentiment Classification Tasks. In EMNLP.
  • Ke et al. (2023) Zixuan Ke, Yijia Shao, Haowei Lin, Tatsuya Konishi, Gyuhak Kim, and Bing Liu. 2023. Continual Pre-training of Language Models. In ICLR.
  • Ke and Xu (2021) Zixuan Ke and Hu Xu. 2021. Adapting BERT for Continual Learning of a Sequence of Aspect Sentiment Classification Tasks. In NAACL.
  • Kemker et al. (2018) Ronald Kemker, Marc McClure, Angelina Abitino, Tyler Hayes, and Christopher Kanan. 2018. Measuring catastrophic forgetting in neural networks. In AAAI, Vol. 32.
  • Kenton and Toutanova (2019) Jacob Devlin Ming-Wei Chang Kenton and Lee Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. In NAACL. 4171–4186.
  • Khan et al. (2023) Muhammad Gul Zain Ali Khan, Muhammad Ferjad Naeem, Luc Van Gool, Didier Stricker, Federico Tombari, and Muhammad Zeshan Afzal. 2023. Introducing language guidance in prompt-based continual learning. In ICCV. 11463–11473.
  • Kim et al. (2024) Junsu Kim, Yunhoe Ku, Jihyeon Kim, Junuk Cha, and Seungryul Baek. 2024. VLM-PL: Advanced Pseudo Labeling approach Class Incremental Object Detection with Vision-Language Model. arXiv preprint arXiv:2403.05346 (2024).
  • Kim et al. (2023) Sein Kim, Namkyeong Lee, Donghyun Kim, Minchul Yang, and Chanyoung Park. 2023. Task Relation-aware Continual User Representation Learning. In SIGKDD. 1107–1119.
  • Kirkpatrick et al. (2017) James Kirkpatrick, Razvan Pascanu, Neil Rabinowitz, Joel Veness, Guillaume Desjardins, Andrei A Rusu, Kieran Milan, John Quan, Tiago Ramalho, Agnieszka Grabska-Barwinska, et al. 2017. Overcoming catastrophic forgetting in neural networks. the national academy of sciences 114, 13 (2017), 3521–3526.
  • Ko et al. (2022) Jihoon Ko, Shinhwan Kang, and Kijung Shin. 2022. BeGin: Extensive Benchmark Scenarios and An Easy-to-use Framework for Graph Continual Learning. arXiv preprint arXiv:2211.14568 (2022).
  • Koh et al. (2021) Hyunseo Koh, Dahyun Kim, Jung-Woo Ha, and Jonghyun Choi. 2021. Online continual learning on class incremental blurry task configuration with anytime inference. arXiv preprint arXiv:2110.10031 (2021).
  • Koh et al. (2022) Hyunseo Koh, Minhyuk Seo, Jihwan Bang, Hwanjun Song, Deokki Hong, Seulki Park, Jung-Woo Ha, and Jonghyun Choi. 2022. Online Boundary-Free Continual Learning by Scheduled Data Prior. In ICLR.
  • Kowsari et al. (2017) Kamran Kowsari, Donald E Brown, Mojtaba Heidarysafa, Kiana Jafari Meimandi, Matthew S Gerber, and Laura E Barnes. 2017. Hdltex: Hierarchical deep learning for text classification. In ICMLA. 364–371.
  • Lang (1995) Ken Lang. 1995. Newsweeder: Learning to filter netnews. In Machine learning proceedings. 331–339.
  • Larson et al. (2019) Stefan Larson, Anish Mahendran, Joseph J Peper, Christopher Clarke, Andrew Lee, Parker Hill, Jonathan K Kummerfeld, Kevin Leach, Michael A Laurenzano, Lingjia Tang, et al. 2019. An evaluation dataset for intent classification and out-of-scope prediction. arXiv preprint arXiv:1909.02027 (2019).
  • Lee et al. (2023) Kuan-Ying Lee, Yuanyi Zhong, and Yu-Xiong Wang. 2023. Do pre-trained models benefit equally in continual learning?. In WCACV. 6485–6493.
  • Lee (2017) Sungjin Lee. 2017. Toward continual learning for conversational agents. arXiv preprint arXiv:1712.09943 (2017).
  • Lesort et al. (2020) Timothée Lesort, Vincenzo Lomonaco, Andrei Stoian, Davide Maltoni, David Filliat, and Natalia Díaz-Rodríguez. 2020. Continual learning for robotics: Definition, framework, learning strategies, opportunities and challenges. Information fusion 58 (2020), 52–68.
  • Lewis et al. (2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020. BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension. In ACL.
  • Li et al. (2022a) Dingcheng Li, Zheng Chen, Eunah Cho, Jie Hao, Xiaohu Liu, Fan Xing, Chenlei Guo, and Yang Liu. 2022a. Overcoming catastrophic forgetting during domain adaptation of seq2seq language generation. In NAACL. 5441–5454.
  • Li et al. (2017) Feng-Lin Li, Minghui Qiu, Haiqing Chen, Xiongwei Wang, Xing Gao, Jun Huang, Juwei Ren, Zhongzhou Zhao, Weipeng Zhao, Lei Wang, et al. 2017. Alime assist: An intelligent assistant for creating an innovative e-commerce experience. In CIKM. 2495–2498.
  • Li et al. (2022b) Guodun Li, Yuchen Zhai, Qianglong Chen, Xing Gao, Ji Zhang, and Yin Zhang. 2022b. Continual few-shot intent detection. In COLING. 333–343.
  • Li et al. (2019) Liunian Harold Li, Mark Yatskar, Da Yin, Cho-Jui Hsieh, and Kai-Wei Chang. 2019. Visualbert: A simple and performant baseline for vision and language. arXiv preprint arXiv:1908.03557 (2019).
  • Lin et al. (2023) Huiwei Lin, Baoquan Zhang, Shanshan Feng, Xutao Li, and Yunming Ye. 2023. PCR: Proxy-based Contrastive Replay for Online Class-Incremental Continual Learning. In CVPR. 24246–24255.
  • Lin et al. (2014) Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C Lawrence Zitnick. 2014. Microsoft coco: Common objects in context. In ECCV. 740–755.
  • Lin et al. (2021) Zhiqiu Lin, Jia Shi, Deepak Pathak, and Deva Ramanan. 2021. The clear benchmark: Continual learning on real-world imagery. In NeurIPS.
  • Liu (2020) Bing Liu. 2020. Learning on the job: Online lifelong and continual learning. In AAAI, Vol. 34. 13544–13549.
  • Liu and Mazumder (2021) Bing Liu and Sahisnu Mazumder. 2021. Lifelong and continual learning dialogue systems: learning during conversation. In AAAI, Vol. 35. 15058–15063.
  • Liu et al. (2023b) Bing Liu, Sahisnu Mazumder, Eric Robertson, and Scott Grigsby. 2023b. AI Autonomy: Self-initiated Open-world Continual Learning and Adaptation. AI Magazine (2023).
  • Liu et al. (2024) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2024. Visual instruction tuning. NeurIPS 36 (2024).
  • Liu et al. (2015) Qian Liu, Zhiqiang Gao, Bing Liu, and Yuanlin Zhang. 2015. Automated rule selection for aspect extraction in opinion mining. In IJCAI.
  • Liu et al. (2021b) Qingbin Liu, Xiaoyan Yu, Shizhu He, Kang Liu, and Jun Zhao. 2021b. Lifelong intent detection via multi-strategy rebalancing. arXiv preprint arXiv:2108.04445 (2021).
  • Liu et al. (2019b) Tianlin Liu, Lyle Ungar, and João Sedoc. 2019b. Continual Learning for Sentence Representations Using Conceptors. In NAACL. 3274–3279.
  • Liu et al. (2023a) Xialei Liu, Xusheng Cao, Haori Lu, Jia-wen Xiao, Andrew D Bagdanov, and Ming-Ming Cheng. 2023a. Class Incremental Learning with Pre-trained Vision-Language Models. arXiv preprint arXiv:2310.20348 (2023).
  • Liu et al. (2021a) Xingkun Liu, Arash Eshghi, Pawel Swietojanski, and Verena Rieser. 2021a. Benchmarking natural language understanding services for building conversational agents. In International Workshop on Spoken Dialogue Systems. 165–183.
  • Liu et al. (2022) Xiao Liu, Kaixuan Ji, Yicheng Fu, Weng Tam, Zhengxiao Du, Zhilin Yang, and Jie Tang. 2022. P-Tuning: Prompt Tuning Can Be Comparable to Fine-tuning Across Scales and Tasks. In ACL. 61–68.
  • Liu et al. (2019a) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019a. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692 (2019).
  • Lo et al. (2020) Kyle Lo, Lucy Lu Wang, Mark Neumann, Rodney Kinney, and Daniel Weld. 2020. S2ORC: The Semantic Scholar Open Research Corpus. In ACL. 4969–4983.
  • Lopez-Paz and Ranzato (2017) David Lopez-Paz and Marc’Aurelio Ranzato. 2017. Gradient episodic memory for continual learning. NeurIPS 30 (2017).
  • Lu et al. (2021) Shuai Lu, Daya Guo, Shuo Ren, Junjie Huang, Alexey Svyatkovskiy, Ambrosio Blanco, Colin Clement, Dawn Drain, Daxin Jiang, Duyu Tang, et al. 2021. Codexglue: A machine learning benchmark dataset for code understanding and generation. arXiv preprint arXiv:2102.04664 (2021).
  • Lu et al. (2017) Xiaoqiang Lu, Binqiang Wang, Xiangtao Zheng, and Xuelong Li. 2017. Exploring models and data for remote sensing image caption generation. IEEE Transactions on Geoscience and Remote Sensing 56, 4 (2017), 2183–2195.
  • Luo et al. (2023) Yun Luo, Zhen Yang, Fandong Meng, Yafu Li, Jie Zhou, and Yue Zhang. 2023. An Empirical Study of Catastrophic Forgetting in Large Language Models During Continual Fine-tuning. arXiv e-prints (2023).
  • Ma et al. (2023) Shirong Ma, Shen Huang, Shulin Huang, Xiaobin Wang, Yangning Li, Hai-Tao Zheng, Pengjun Xie, Fei Huang, and Yong Jiang. 2023. EcomGPT-CT: Continual pre-training of e-commerce large language models with semi-structured data. arXiv preprint arXiv:2312.15696 (2023).
  • Maas et al. (2011) Andrew L. Maas, Raymond E. Daly, Peter T. Pham, Dan Huang, Andrew Y. Ng, and Christopher Potts. 2011. Learning Word Vectors for Sentiment Analysis. In ACL. 142–150.
  • Madotto et al. (2020) Andrea Madotto, Zhaojiang Lin, Zhenpeng Zhou, Seungwhan Moon, Paul Crook, Bing Liu, Zhou Yu, Eunjoon Cho, and Zhiguang Wang. 2020. Continual learning in task-oriented dialogue systems. arXiv preprint arXiv:2012.15504 (2020).
  • Maekawa et al. (2023) Aru Maekawa, Hidetaka Kamigaito, Kotaro Funakoshi, and Manabu Okumura. 2023. Generative Replay Inspired by Hippocampal Memory Indexing for Continual Language Learning. In EACL. 930–942.
  • Mai et al. (2022) Zheda Mai, Ruiwen Li, Jihwan Jeong, David Quispe, Hyunwoo Kim, and Scott Sanner. 2022. Online continual learning in image classification: An empirical survey. Neurocomputing 469 (2022), 28–51.
  • Masana et al. (2022) Marc Masana, Xialei Liu, Bartłomiej Twardowski, Mikel Menta, Andrew D Bagdanov, and Joost Van De Weijer. 2022. Class-incremental learning: survey and performance evaluation on image classification. TPAMI 45, 5 (2022), 5513–5533.
  • Mazumder and Liu (2024) Sahisnu Mazumder and Bing Liu. 2024. Lifelong and Continual Learning Dialogue Systems. Springer Nature.
  • McCann et al. ([n. d.]) Bryan McCann, Nitish Shirish Keskar, Caiming Xiong, and Richard Socher. [n. d.]. The natural language decathlon: Multitask learning as question answering. arXiv 2018. arXiv preprint arXiv:1806.08730 ([n. d.]).
  • Mehta et al. (2021) Sanket Vaibhav Mehta, Darshan Patil, Sarath Chandar, and Emma Strubell. 2021. An Empirical Investigation of the Role of Pre-training in Lifelong Learning. (2021).
  • Michieli et al. (2023) Umberto Michieli, Pablo Peso Parada, and Mete Ozay. 2023. Online Continual Learning in Keyword Spotting for Low-Resource Devices via Pooling High-Order Temporal Statistics. arXiv preprint arXiv:2307.12660 (2023).
  • Min et al. (2023) Bonan Min, Hayley Ross, Elior Sulem, Amir Pouran Ben Veyseh, Thien Huu Nguyen, Oscar Sainz, Eneko Agirre, Ilana Heintz, and Dan Roth. 2023. Recent advances in natural language processing via large pre-trained language models: A survey. Comput. Surveys 56, 2 (2023), 1–40.
  • Mok et al. (2023) Jisoo Mok, Jaeyoung Do, Sungjin Lee, Tara Taghavi, Seunghak Yu, and Sungroh Yoon. 2023. Large-scale Lifelong Learning of In-context Instructions and How to Tackle It. In ACL. 12573–12589.
  • Monaikul et al. (2021) Natawut Monaikul, Giuseppe Castellucci, Simone Filice, and Oleg Rokhlenko. 2021. Continual learning for named entity recognition. In AAAI, Vol. 35. 13570–13577.
  • Moon et al. (2023) Jun-Yeong Moon, Keon-Hee Park, Jung Uk Kim, and Gyeong-Moon Park. 2023. Online Class Incremental Learning on Stochastic Blurry Task Boundary via Mask and Visual Prompt Tuning. In ICCV. 11731–11741.
  • Mundra et al. (2024) Nandini Mundra, Sumanth Doddapaneni, Raj Dabre, Anoop Kunchukuttan, Ratish Puduppully, and Mitesh M Khapra. 2024. A Comprehensive Analysis of Adapter Efficiency. In IKDD. 136–154.
  • Ni et al. (2023) Zixuan Ni, Longhui Wei, Siliang Tang, Yueting Zhuang, and Qi Tian. 2023. Continual vision-language representation learning with off-diagonal information. In ICML. 26129–26149.
  • Novikova et al. (2017) Jekaterina Novikova, Ondřej Dušek, and Verena Rieser. 2017. The E2E dataset: New challenges for end-to-end generation. arXiv preprint arXiv:1706.09254 (2017).
  • Ororbia et al. (2022) Alex Ororbia, Ankur Mali, C Lee Giles, and Daniel Kifer. 2022. Lifelong neural predictive coding: Learning cumulatively online without forgetting. NeurIPS 35 (2022), 5867–5881.
  • Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. NeurIPS 35 (2022), 27730–27744.
  • Parisi et al. (2019) German I Parisi, Ronald Kemker, Jose L Part, Christopher Kanan, and Stefan Wermter. 2019. Continual lifelong learning with neural networks: A review. Neural networks 113 (2019), 54–71.
  • PENG et al. (2024) Bohao PENG, Zhuotao Tian, Shu Liu, Ming-Chang Yang, and Jiaya Jia. 2024. Scalable Language Model with Generalized Continual Learning. In ICLR.
  • Poth et al. (2023) Clifton Poth, Hannah Sterz, Indraneil Paul, Sukannya Purkayastha, Leon Engländer, Timo Imhof, Ivan Vulić, Sebastian Ruder, Iryna Gurevych, and Jonas Pfeiffer. 2023. Adapters: A Unified Library for Parameter-Efficient and Modular Transfer Learning. In EMNLP. 149–160.
  • Qian et al. (2023) Zi Qian, Xin Wang, Xuguang Duan, Pengda Qin, Yuhong Li, and Wenwu Zhu. 2023. Decouple before interact: Multi-modal prompt learning for continual visual question answering. In ICCV. 2953–2962.
  • Qin and Joty (2022a) Chengwei Qin and Shafiq Joty. 2022a. Continual Few-shot Relation Learning via Embedding Space Regularization and Data Augmentation. In ACL. 2776–2789.
  • Qin and Joty (2022b) Chengwei Qin and Shafiq Joty. 2022b. LFPT5: A Unified Framework for Lifelong Few-shot Language Learning Based on Prompt Tuning of T5. In ICLR.
  • Qin et al. (2023) Yujia Qin, Cheng Qian, Xu Han, Yankai Lin, Huadong Wang, Ruobing Xie, Zhiyuan Liu, Maosong Sun, and Jie Zhou. 2023. Recyclable Tuning for Continual Pre-training. In ACL. 11403–11426.
  • Qin et al. (2022) Yujia Qin, Jiajie Zhang, Yankai Lin, Zhiyuan Liu, Peng Li, Maosong Sun, and Jie Zhou. 2022. ELLE: Efficient Lifelong Pre-training for Emerging Data. In ACL. 2789–2810.
  • Qu et al. (2021) Haoxuan Qu, Hossein Rahmani, Li Xu, Bryan Williams, and Jun Liu. 2021. Recent advances of continual learning in computer vision: An overview. arXiv preprint arXiv:2109.11369 (2021).
  • Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In ICML. PMLR, 8748–8763.
  • Rajpurkar et al. (2016) Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016. SQuAD: 100,000+ Questions for Machine Comprehension of Text. In EMNLP. 2383–2392.
  • Ramesh et al. (2021) Aditya Ramesh, Mikhail Pavlov, Gabriel Goh, Scott Gray, Chelsea Voss, Alec Radford, Mark Chen, and Ilya Sutskever. 2021. Zero-shot text-to-image generation. In ICML. Pmlr, 8821–8831.
  • Rastogi et al. (2020) Abhinav Rastogi, Xiaoxue Zang, Srinivas Sunkara, Raghav Gupta, and Pranav Khaitan. 2020. Towards scalable multi-domain conversational agents: The schema-guided dialogue dataset. In AAAI, Vol. 34. 8689–8696.
  • Razdaibiedina et al. (2023) Anastasia Razdaibiedina, Yuning Mao, Rui Hou, Madian Khabsa, Mike Lewis, and Amjad Almahairi. 2023. Progressive Prompts: Continual Learning for Language Models. In ICLR.
  • Rebuffi et al. (2017) Sylvestre-Alvise Rebuffi, Alexander Kolesnikov, Georg Sperl, and Christoph H. Lampert. 2017. iCaRL: Incremental Classifier and Representation Learning. In CVPR.
  • Sang and De Meulder (2003) Erik F Sang and Fien De Meulder. 2003. Introduction to the CoNLL-2003 shared task: Language-independent named entity recognition. arXiv preprint cs/0306050 (2003).
  • Scao et al. (2022) Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, Matthias Gallé, et al. 2022. Bloom: A 176b-parameter open-access multilingual language model. arXiv preprint arXiv:2211.05100 (2022).
  • Schuster et al. (2018) Sebastian Schuster, Sonal Gupta, Rushin Shah, and Mike Lewis. 2018. Cross-lingual transfer learning for multilingual task oriented dialog. arXiv preprint arXiv:1810.13327 (2018).
  • Scialom et al. (2022a) Thomas Scialom, Tuhin Chakrabarty, and Smaranda Muresan. 2022a. Fine-tuned Language Models are Continual Learners. In EMNLP. 6107–6122.
  • Scialom et al. (2022b) Thomas Scialom, Tuhin Chakrabarty, and Smaranda Muresan. 2022b. Fine-tuned language models are continual learners. In EMNLP. 6107–6122.
  • See et al. (2017) Abigail See, Peter J Liu, and Christopher D Manning. 2017. Get to the point: Summarization with pointer-generator networks. arXiv preprint arXiv:1704.04368 (2017).
  • Semola et al. (2022) Rudy Semola, Vincenzo Lomonaco, and Davide Bacciu. 2022. Continual-learning-as-a-service (claas): On-demand efficient adaptation of predictive models. arXiv preprint arXiv:2206.06957 (2022).
  • Shaheen et al. (2022) Khadija Shaheen, Muhammad Abdullah Hanif, Osman Hasan, and Muhammad Shafique. 2022. Continual learning for real-world autonomous systems: Algorithms, challenges and frameworks. Journal of Intelligent & Robotic Systems 105, 1 (2022), 9.
  • Shen et al. (2019) Yilin Shen, Xiangyu Zeng, and Hongxia Jin. 2019. A progressive model to enable continual learning for semantic slot filling. In EMNLP. 1279–1284.
  • Shim et al. (2021) Dongsub Shim, Zheda Mai, Jihwan Jeong, Scott Sanner, Hyunwoo Kim, and Jongseong Jang. 2021. Online class-incremental continual learning with adversarial shapley value. In AAAI, Vol. 35. 9630–9638.
  • Song et al. (2023) Chenyang Song, Xu Han, Zheni Zeng, Kuai Li, Chen Chen, Zhiyuan Liu, Maosong Sun, and Tao Yang. 2023. ConPET: Continual Parameter-Efficient Tuning for Large Language Models. arXiv preprint arXiv:2309.14763 (2023).
  • Sun et al. (2020a) Fan-Keng Sun, Cheng-Hao Ho, and Hung-Yi Lee. 2020a. LAMOL: LAnguage MOdeling for Lifelong Language Learning. In ICLR. OpenReview.net.
  • Sun et al. (2020b) Yu Sun, Shuohuan Wang, Yukun Li, Shikun Feng, Hao Tian, Hua Wu, and Haifeng Wang. 2020b. Ernie 2.0: A continual pre-training framework for language understanding. In AAAI, Vol. 34. 8968–8975.
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971 (2023).
  • Tufano et al. (2019) Michele Tufano, Cody Watson, Gabriele Bavota, Massimiliano Di Penta, Martin White, and Denys Poshyvanyk. 2019. An empirical study on learning bug-fixing patches in the wild via neural machine translation. TOSEM 28, 4 (2019), 1–29.
  • Van de Ven et al. (2022) Gido M Van de Ven, Tinne Tuytelaars, and Andreas S Tolias. 2022. Three types of incremental learning. Nature Machine Intelligence 4, 12 (2022), 1185–1197.
  • Vander Eeckt et al. (2023) Steven Vander Eeckt et al. 2023. Rehearsal-Free Online Continual Learning for Automatic Speech Recognition. arXiv e-prints (2023), arXiv–2306.
  • Varshney et al. (2022) Vaibhav Varshney, Mayur Patidar, Rajat Kumar, Lovekesh Vig, and Gautam Shroff. 2022. Prompt augmented generative replay via supervised contrastive learning for lifelong intent detection. In NAACL. 1113–1127.
  • Verwimp et al. (2023) Eli Verwimp, Kuo Yang, Sarah Parisot, Lanqing Hong, Steven McDonagh, Eduardo Pérez-Pellitero, Matthias De Lange, and Tinne Tuytelaars. 2023. Clad: A realistic continual learning benchmark for autonomous driving. Neural Networks 161 (2023), 659–669.
  • Volske et al. (2017) Michael Volske, Martin Potthast, Shahbaz Syed, and Benno Stein. 2017. TL;DR: Mining Reddit to Learn Automatic Summarization. In the Workshop on New Frontiers in Summarization. 59–63.
  • Wang et al. (2019a) Alex Wang, Yada Pruksachatkun, Nikita Nangia, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel Bowman. 2019a. Superglue: A stickier benchmark for general-purpose language understanding systems. NeurIPS 32 (2019).
  • Wang et al. (2018) Alex Wang, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel Bowman. 2018. GLUE: A Multi-Task Benchmark and Analysis Platform for Natural Language Understanding. In EMNLP. 353–355.
  • Wang et al. (2021a) Chengyu Wang, Haojie Pan, Yuan Liu, Kehan Chen, Minghui Qiu, Wei Zhou, Jun Huang, Haiqing Chen, Wei Lin, and Deng Cai. 2021a. Mell: Large-scale extensible user intent classification for dialogue systems with meta lifelong learning. In SIGKDD. 3649–3659.
  • Wang et al. (2022a) Han Wang, Ruiliu Fu, Xuejun Zhang, and Jun Zhou. 2022a. RVAE-LAMOL: Residual Variational Autoencoder to Enhance Lifelong Language Learning. In IJCNN. IEEE, 1–9.
  • Wang et al. (2019b) Hong Wang, Wenhan Xiong, Mo Yu, Xiaoxiao Guo, Shiyu Chang, and William Yang Wang. 2019b. Sentence Embedding Alignment for Lifelong Relation Extraction. In NAACL. 796–806.
  • Wang et al. (2021b) Jianren Wang, Xin Wang, Yue Shang-Guan, and Abhinav Gupta. 2021b. Wanderlust: Online continual object detection in the real world. In ICCV. 10829–10838.
  • Wang et al. (2024b) Liyuan Wang, Xingxing Zhang, Hang Su, and Jun Zhu. 2024b. A comprehensive survey of continual learning: Theory, method and application. TPAMI (2024).
  • Wang et al. (2023d) Quanziang Wang, Renzhen Wang, Yichen Wu, Xixi Jia, and Deyu Meng. 2023d. CBA: Improving Online Continual Learning via Continual Bias Adaptor. In ICCV. 19082–19092.
  • Wang et al. (2023b) Xiao Wang, Guangyao Chen, Guangwu Qian, Pengcheng Gao, Xiao-Yong Wei, Yaowei Wang, Yonghong Tian, and Wen Gao. 2023b. Large-scale multi-modal pre-trained models: A comprehensive survey. Machine Intelligence Research 20, 4 (2023), 447–482.
  • Wang et al. (2023a) Xiao Wang, Tianze Chen, Qiming Ge, Han Xia, Rong Bao, Rui Zheng, Qi Zhang, Tao Gui, and Xuanjing Huang. 2023a. Orthogonal Subspace Learning for Language Model Continual Learning. In EMNLP. 10658–10671.
  • Wang et al. (2024a) Xiao Wang, Yuansen Zhang, Tianze Chen, Songyang Gao, Senjie Jin, Xianjun Yang, Zhiheng Xi, Rui Zheng, Yicheng Zou, Tao Gui, Qi Zhang, and Xuanjing Huang. 2024a. TRACE: A Comprehensive Benchmark for Continual Learning in Large Language Models.
  • Wang et al. (2022b) Yabin Wang, Zhiwu Huang, and Xiaopeng Hong. 2022b. S-prompts learning with pre-trained transformers: An occam’s razor for domain incremental learning. NeurIPS 35 (2022), 5682–5695.
  • Wang et al. (2022c) Yizhong Wang, Swaroop Mishra, Pegah Alipoormolabashi, Yeganeh Kordi, Amirreza Mirzaei, Anjana Arunkumar, Arjun Ashok, Arut Selvan Dhanasekaran, Atharva Naik, David Stap, et al. 2022c. Super-naturalinstructions: Generalization via declarative instructions on 1600+ nlp tasks. arXiv preprint arXiv:2204.07705 (2022).
  • Wang et al. (2023c) Zhicheng Wang, Yufang Liu, Tao Ji, Xiaoling Wang, Yuanbin Wu, Congcong Jiang, Ye Chao, Zhencong Han, Ling Wang, Xu Shao, et al. 2023c. Rehearsal-free Continual Language Learning via Efficient Parameter Isolation. In ACL. 10933–10946.
  • Wang et al. (2020) Zirui Wang, Sanket Vaibhav Mehta, Barnabas Poczos, and Jaime Carbonell. 2020. Efficient Meta Lifelong-Learning with Limited Memory. In EMNLP. 535–548.
  • Wei et al. (2023) Yujie Wei, Jiaxin Ye, Zhizhong Huang, Junping Zhang, and Hongming Shan. 2023. Online Prototype Learning for Online Continual Learning. In ICCV. 18764–18774.
  • Wen et al. (2015) Tsung-Hsien Wen, Milica Gasic, Nikola Mrksic, Pei-Hao Su, David Vandyke, and Steve Young. 2015. Semantically conditioned lstm-based natural language generation for spoken dialogue systems. arXiv preprint arXiv:1508.01745 (2015).
  • Wójcik et al. (2023) Mateusz Wójcik, Witold Kościukiewicz, Mateusz Baran, Tomasz Kajdanowicz, and Adam Gonczarek. 2023. Domain-Agnostic Neural Architecture for Class Incremental Continual Learning in Document Processing Platform. In ACL. 527–537.
[…]
  • Wu et al. (2023) Yuhao Wu, Tongjun Shi, Karthick Sharma, Chun Wei Seah, and Shuhao Zhang. 2023. Online Continual Knowledge Learning for Language Models. arXiv preprint arXiv:2311.09632 (2023).
  • Xia et al. (2021) Congying Xia, Wenpeng Yin, Yihao Feng, and Philip Yu. 2021. Incremental few-shot text classification with multi-round new classes: Formulation, dataset and system. arXiv preprint arXiv:2104.11882 (2021).
  • Xie et al. (2023) Yong Xie, Karan Aggarwal, and Aitzaz Ahmad. 2023. Efficient continual pre-training for building domain specific large language models. arXiv preprint arXiv:2311.08545 (2023).
  • Xu et al. (2019) Hu Xu, Bing Liu, Lei Shu, and Philip S Yu. 2019. BERT post-training for review reading comprehension and aspect-based sentiment analysis. arXiv preprint arXiv:1904.02232 (2019).
  • Xu et al. (2017) Jiaming Xu, Bo Xu, Peng Wang, Suncong Zheng, Guanhua Tian, and Jun Zhao. 2017. Self-taught convolutional neural networks for short text clustering. Neural Networks 88 (2017), 22–31.
  • Yadav et al. (2023) Prateek Yadav, Qing Sun, Hantian Ding, Xiaopeng Li, Dejiao Zhang, Ming Tan, Xiaofei Ma, Parminder Bhatia, Ramesh Nallapati, Murali Krishna Ramanathan, et al. 2023. Exploring continual learning for code generation models. arXiv preprint arXiv:2307.02435 (2023).
  • Yang et al. (2021) An Yang, Junyang Lin, Rui Men, Chang Zhou, Le Jiang, Xianyan Jia, Ang Wang, Jie Zhang, Jiamang Wang, Yong Li, et al. 2021. M6-t: Exploring sparse expert models and beyond. arXiv preprint arXiv:2105.15082 (2021).
  • Yi et al. (2023) Huahui Yi, Ziyuan Qin, Qicheng Lao, Wei Xu, Zekun Jiang, Dequan Wang, Shaoting Zhang, and Kang Li. 2023. Towards General Purpose Medical AI: Continual Learning Medical Foundation Model. arXiv preprint arXiv:2303.06580 (2023).
  • Yin et al. (2021) Haiyan Yin, Ping Li, et al. 2021. Mitigating forgetting in online continual learning with neuron calibration. NeurIPS 34 (2021), 10260–10272.
  • Yin et al. (2022) Wenpeng Yin, Jia Li, and Caiming Xiong. 2022. ConTinTin: Continual Learning from Task Instructions. In ACL. 3062–3072.
  • Yogatama et al. (2019) Dani Yogatama, Cyprien de Masson d’Autume, Jerome Connor, Tomas Kocisky, Mike Chrzanowski, Lingpeng Kong, Angeliki Lazaridou, Wang Ling, Lei Yu, Chris Dyer, et al. 2019. Learning and evaluating general linguistic intelligence. arXiv preprint arXiv:1901.11373 (2019).
  • Yoon et al. (2021) Jaehong Yoon, Divyam Madaan, Eunho Yang, and Sung Ju Hwang. 2021. Online coreset selection for rehearsal-based continual learning. arXiv preprint arXiv:2106.01085 (2021).
  • Young et al. (2014) Peter Young, Alice Lai, Micah Hodosh, and Julia Hockenmaier. 2014. From image descriptions to visual denotations: New similarity metrics for semantic inference over event descriptions. TACL 2 (2014), 67–78.
  • Yu et al. (2024) Jiazuo Yu, Yunzhi Zhuge, Lu Zhang, Ping Hu, Dong Wang, Huchuan Lu, and You He. 2024. Boosting Continual Learning of Vision-Language Models via Mixture-of-Experts Adapters. In CVPR.
  • Yuan et al. (2020) Fajie Yuan, Xiangnan He, Alexandros Karatzoglou, and Liguang Zhang. 2020. Parameter-efficient transfer from sequential behaviors for user modeling and recommendation. In SIGIR. 1469–1478.
  • Yuan et al. (2021) Fajie Yuan, Guoxiao Zhang, Alexandros Karatzoglou, Joemon Jose, Beibei Kong, and Yudong Li. 2021. One Person, One Model, One World: Learning Continual User Representation without Forgetting. 696–705.
  • Zellers et al. (2019) Rowan Zellers, Ari Holtzman, Hannah Rashkin, Yonatan Bisk, Ali Farhadi, Franziska Roesner, and Yejin Choi. 2019. Defending against neural fake news. NeurIPS 32 (2019).
  • Zhang et al. (2021) Chen Zhang, Yu Xie, Hang Bai, Bin Yu, Weihong Li, and Yuan Gao. 2021. A survey on federated learning. Knowledge-Based Systems 216 (2021), 106775.
  • Zhang et al. (2023b) Han Zhang, Lin Gui, Yuanzhao Zhai, Hui Wang, Yu Lei, and Ruifeng Xu. 2023b. Copf: Continual learning human preference through optimal policy fitting. arXiv preprint arXiv:2310.15694 (2023).
  • Zhang et al. (2024) Han Zhang, Yu Lei, Lin Gui, Min Yang, Yulan He, Hui Wang, and Ruifeng Xu. 2024. CPPO: Continual Learning for Reinforcement Learning with Human Feedback. In ICLR.
  • Zhang and Kim (2023) Peiyan Zhang and Sunghun Kim. 2023. A Survey on Incremental Update for Neural Recommender Systems. arXiv preprint arXiv:2303.02851 (2023).
  • Zhang et al. (2023d) Xi Zhang, Feifei Zhang, and Changsheng Xu. 2023d. Vqacl: A novel visual question answering continual learning setting. In CVPR. 19102–19112.
  • Zhang et al. (2015) Xiang Zhang, Junbo Zhao, and Yann LeCun. 2015. Character-level convolutional networks for text classification. NeurIPS 28 (2015).
  • Zhang et al. (2022a) Yaqian Zhang, Bernhard Pfahringer, Eibe Frank, Albert Bifet, Nick Jin Sean Lim, and Yunzhe Jia. 2022a. A simple but strong baseline for online continual learning: Repeated augmented rehearsal. NeurIPS 35 (2022), 14771–14783.
  • Zhang et al. (2022b) Yanzhe Zhang, Xuezhi Wang, and Diyi Yang. 2022b. Continual Sequence Generation with Adaptive Compositional Modules. In ACL. 3653–3667.
  • Zhang et al. (2023c) Yating Zhang, Yexiang Wang, Fei Cheng, Sadao Kurohashi, et al. 2023c. Reformulating Domain Adaptation of Large Language Models as Adapt-Retrieve-Revise. arXiv preprint arXiv:2310.03328 (2023).
  • Zhang et al. (2023a) Zihan Zhang, Meng Fang, Ling Chen, and Mohammad-Reza Namazi-Rad. 2023a. Citb: A benchmark for continual instruction tuning. arXiv preprint arXiv:2310.14510 (2023).
  • Zhao et al. (2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023. A survey of large language models. arXiv preprint arXiv:2303.18223 (2023).
  • Zhao et al. (2022) Yingxiu Zhao, Yinhe Zheng, Zhiliang Tian, Chang Gao, Jian Sun, and Nevin L. Zhang. 2022. Prompt Conditioned VAE: Enhancing Generative Replay for Lifelong Learning in Task-Oriented Dialogue. In EMNLP. 11153–11169.
  • Zheng et al. (2023) Zangwei Zheng, Mingyuan Ma, Kai Wang, Ziheng Qin, Xiangyu Yue, and Yang You. 2023. Preventing zero-shot transfer degradation in continual learning of vision-language models. In ICCV. 19125–19136.
  • Zhong et al. (2017) Victor Zhong, Caiming Xiong, and Richard Socher. 2017. Seq2sql: Generating structured queries from natural language using reinforcement learning. arXiv preprint arXiv:1709.00103 (2017).
  • Zhou et al. (2023b) Da-Wei Zhou, Qi-Wei Wang, Zhi-Hong Qi, Han-Jia Ye, De-Chuan Zhan, and Ziwei Liu. 2023b. Deep class-incremental learning: A survey. arXiv preprint arXiv:2302.03648 (2023).
  • Zhou et al. (2023c) Da-Wei Zhou, Yuanhan Zhang, Jingyi Ning, Han-Jia Ye, De-Chuan Zhan, and Ziwei Liu. 2023c. Learning without forgetting for vision-language models. arXiv preprint arXiv:2305.19270 (2023).
  • Zhou et al. (2023a) Jie Zhou, Pei Ke, Xipeng Qiu, Minlie Huang, and Junping Zhang. 2023a. ChatGPT: potential, prospects, and limitations. Frontiers of Information Technology & Electronic Engineering (2023), 1–6.
  • Zhu et al. (2023) Hongguang Zhu, Yunchao Wei, Xiaodan Liang, Chunjie Zhang, and Yao Zhao. 2023. Ctp: Towards vision-language continual pretraining via compatible momentum contrast and topology preservation. In ICCV. 22257–22267.
  • Zhu et al. (2022) Qi Zhu, Bing Li, Fei Mi, Xiaoyan Zhu, and Minlie Huang. 2022. Continual Prompt Tuning for Dialog State Tracking. In ACL. 1124–1137.
  • Zhu et al. (2015) Yukun Zhu, Ryan Kiros, Rich Zemel, Ruslan Salakhutdinov, Raquel Urtasun, Antonio Torralba, and Sanja Fidler. 2015. Aligning books and movies: Towards story-like visual explanations by watching movies and reading books. In ICCV. 19–27.
Appendix A Details of Metrics
A.1. Overall Performance.
Moreover, Chaudhry et al. (Chaudhry et al., 2019) devise a metric known as Learning Curve Area (LCA), which quantifies the speed of learning in a model. It first defines an average $b$-shot performance, where $b$ represents the number of mini-batches, subsequent to the completion of training across all $T$ tasks as follows:
(7) $Z_{b}=\frac{1}{N}\sum_{i=1}^{N}R_{N,i}$
$LCA$ at $\beta$ is the area of the convergence curve $Z_{b}$ as a function of $b\in[0,\beta]$:
[…]
The Learning Curve Area (LCA) provides insights into model learning dynamics. $LCA_{0}$ measures the average 00-shot performance, similar to forward transfer ((Lopez-Paz and Ranzato, 2017)). $LCA_{\beta}$, quantifying the area under the $Z_{b}$ curve, evaluates both average 00-shot performance and learning speed. Although two models may achieve similar $Z_{b}$ or $A_{T}$ values, they can differ significantly in $LCA_{\beta}$ due to variations in learning rates. This metric is crucial for identifying models that quickly learn from few examples, particularly when $\beta$ is small.
Qin et al. (Qin et al., 2022) propose two metrics designed to evaluate pre-trained language models (PLMs) based on their performance within learned domains: Average Perplexity ($AP$) and Average Increased Perplexity ($AP^{+}$). The aforementioned metrics are utilized to assess key capabilities of PLMs, such as instruction following and safety, as discussed in Wang et al. (Wang et al., 2024a).
A.2. Memory Stability.
Chaudhry et al. (Chaudhry et al., 2018) introduce the Forgetting Measure (FM), a metric designed to quantify the extent of forgetting a model experiences for a specific task. The forgetting for a given task $T_{j}$ after sequential training on tasks up to $T_{N}$ is quantified as the difference between the highest proficiency ($max(R_{l,j})$) achieved on task $T_{j}$ during initial training and its proficiency ($R_{N,j}$) after subsequent learning phases:
(9) $f_{j}=\max_{l\in\{1,\ldots,N-1\}}(R_{l,j}-R_{N,j}),\quad\forall j<N.$
For the purpose of quantifying forgetting in previous tasks, the function $f_{j}$ is defined within the interval $[-1,1]$ for $j<N$.
Furthermore, to account for the number of tasks previously encountered, the Forgetting Measure (FM) at the $N$-th task represents the mean level of forgetting across all preceding tasks:
(10) $FM=\frac{1}{N-1}\sum_{j=1}^{N-1}f_{j}$
A lower FM indicates better retention of previous tasks. Here, the expansion or $R_{j,j}$ serves as a more effective quantifier of retained knowledge concerning past tasks, as opposed to using max. Nonetheless, max remains a valuable estimator for assessing the extent of forgetting that occurs throughout the learning process.
Davari et al. (Davari et al., 2022) propose a method named linear probes (LP) to assess representation forgetting. This approach measures the effectiveness of learned representations via an optimal linear classifier trained on the frozen activations of a base network. Representation forgetting is quantified by evaluating the change in Language Processing (LP) performance before and after the introduction of a new task. Formally, for each model ($f_{\theta_{i}}$) at time step $i$ of a task sequence, the classifier ($W^{*}_{i}$) is optimized as: $W^{*}_{i}=\arg\min_{W_{i}}\mathcal{L}(W_{i};f_{\theta_{i}}(X_{i}),Y_{i})$, where $\mathcal{L}$, $X_{i}$, and $Y_{i}$ represent the objective function, input data, and labels for task $i$, respectively. The degree of representational forgetting between two model states, $\theta_{a}$ and $\theta_{b}$, where $\theta_{b}$ is derived later in the sequence, is evaluated by calculating the difference in scores: $Score(W_{a}f_{\theta_{a}}(X_{a}),Y_{a})-Score(W_{b}f_{\theta_{b}}(X_{a}),Y_{a})$, where $Score$ represents the performance metric, such as accuracy, on the task.
Kemker et al. (Kemker et al., 2018) introduce three metrics designed to CF: $\Omega_{\text{base}}$, $\Omega_{\text{new}}$, and $\Omega_{\text{all}}$. $\Omega_{\text{base}}$ assesses retention of initial learning, $\Omega_{\text{new}}$ measures recall of new tasks, and $\Omega_{\text{all}}$ evaluates overall proficiency in maintaining old knowledge and acquiring new information.
(11) $\Omega_{\text{base}}=\frac{1}{N-1}\sum_{i=2}^{N}\frac{\alpha_{\text{base},i}}{\alpha_{\text{ideal}}}$
(12) $\Omega_{\text{new}}=\frac{1}{N-1}\sum_{i=2}^{N}\frac{\alpha_{\text{new},i}}{\alpha_{\text{ideal}}}$
(13) $\Omega_{\text{all}}=\frac{1}{N-1}\sum_{i=2}^{N}\frac{\alpha_{\text{all},i}}{\alpha_{\text{ideal}}}$
where $N$ represents the total number of sessions, $\alpha_{\text{new},i}$ is the test accuracy after learning session $i$, $\alpha_{\text{base},i}$ denotes the accuracy on the initial session after $i$ sessions, and $\alpha_{\text{all},i}$ refers to the test accuracy across all test data for classes encountered up to point $i$. The ideal performance ($\alpha_{\text{ideal}}$) is defined as the offline MLP accuracy on the base set. To facilitate comparative analysis across different datasets, $\Omega_{\text{base}}$ and $\Omega_{\text{all}}$ are normalized by $\alpha_{\text{ideal}}$. Consequently, unless a model surpasses $\alpha_{\text{ideal}}$, normalized results will range from 0 to 1, enabling consistent cross-dataset comparisons.
Additionally, researchers (Koh et al., 2022) devise a novel metric, termed the Knowledge Loss Ratio (KLR), quantifies knowledge degradation using principles from information theory.
A.3. Learning Plasticity.
Intransigence measure (IM), as defined by Chaudhry et al. (Chaudhry et al., 2018), quantifies a model’s inability to learn new tasks. This measure is calculated by comparing the performance difference of a task when trained jointly with other tasks versus when trained in a continual learning setting. Then the intransigence for the $N$-th task can be defined as:
(14) $IM=R^{*}_{N}-R_{N,N},$
where $R^{*}_{N}$ represents the accuracy achieved on the held-out dataset of the $N$-th task, $R_{N,N}$ indicates the accuracy on the $N$-th task upon completion of training in an incremental sequence up to and including task $N$. Note, $IM_{N}\in[-1,1]$, and lower values indicate superior performance.
A.4. Metrics for Continual Pre-training.
CKL (Jang et al., 2022b) introduces a novel metric, named FUAR (FORGOTTEN / (UPDATED + ACQUIRED) RATIO), which quantitatively measures the efficiency of each CKL method. It calculates the number of instances of time-invariant knowledge that a model forgets in order to learn or update one instance of new knowledge. When FUAR is equal to 1.0, it signifies an equilibrium where one time-invariant knowledge instance is forgotten on average to obtain a new or updated knowledge instance. Formally, FUAR is defined as:
(15) $Eq_{1}=\sum_{i=0}^{N-1}\max(0,\text{Gap}(T_{i}^{F},D_{i},D_{N}))\mathbbm{1}_{\{T_{i}^{F}\neq n.d.\}}$
(16) $\begin{split}Eq_{2}=\sum_{i=0}^{N-1}\max(0,\text{Gap}(T_{B}^{U},D_{N},D_{i}))\mathbbm{1}_{\{T_{i}^{F}\neq n.d.\}}\\ +\max(0,\text{Gap}(T_{N}^{A},D_{N},D_{i}))\mathbbm{1}_{\{T_{i}^{F}\neq n.d.\}}\end{split}$
(17) $FUAR(\mathbbm{T}^{F},T_{N}^{U},T_{N}^{A})=\left\{\begin{aligned} &\frac{Eq_{1}}{Eq_{2}}\quad&&\text{if }denominator>0\\ &no\;gain\quad&&\text{otherwise }\end{aligned}\right.$
where $T$ represents an arbitrary task, and $(D_{i})_{i=0}^{N}$ is a sequence of corpora for LM pretraining. Gap($T,D_{a},D_{b}$) is $\text{Score}(T)$ of $LM_{a}$ $\text{Score}(T)$ of $LM_{b}$, where $LM_{a}$ is pretrained on $D_{a}$. $\mathbbm{T}^{F}=(T_{i}^{F})_{i=0}^{N-1}$ measures forgetting of invariant-knowledge from $(D_{i})_{i=0}^{N-1}$. If no task is from $D_{i}$, $T_{i}^{F}$ is ”n.d.” (not defined). $T_{N}^{U}$ and $T_{N}^{A}$ from $D_{N}$ measure update and acquisition of new knowledge, respectively.
A.5. Online CL-Specific Metrics.
Near-future accuracy (NFA) (Al Kader Hammoud et al., 2023) is introduced as a novel evaluation metric for OCL problem. Unlike traditional evaluation methods that assess models on immediately subsequent samples, NFA evaluates models on samples slightly further into the future, using a minimal shift $S$. Such operation can mitigate label correlation effects, which can adversely impact the accuracy of model adaptability assessments. The smallest shift $S$ is selected to ensure that the test sample aligns closely with the distribution of recently observed training data. The calculation of NFA involves first checking if the model correctly predicts the label of a future sample, which can be expressed as $a_{t}=\mathbbm{1}\{f_{\theta_{t}}(x_{t+1+S})=y_{t+1+S}\}$. Subsequently, the running average is updated using the formula $A^{RA}t=\frac{1}{t}(A^{RA}{t-1}\cdot(t-1)+a_{t})$.
Yogatama et al. (Yogatama et al., 2019) proposed a novel online codelength ($\ell(D)$), inspired by prequential encoding (Blier and Ollivier, 2018), to quantify how quickly an existing model can adapt to a new task.
(18) $\ell(D)=\log_{2}|y|-\sum_{i=2}^{N}\log_{2}p(y_{i}|x_{i};\theta_{D_{i-1}})$
where $|Y|$ is the number of possible labels (classes), and $\theta_{D_{i}}$ represents a particular subset of the dataset $D$. Similar to the approach in Latent Contextual Allocation (LCA) (Chaudhry et al., 2019), the concept of online codelength is associated with the area under the learning curve.

### v2

@@
 Unlike traditional neural network models, foundation LMs obtain a great ability for transfer learning by acquiring rich commonsense knowledge through pre-training on extensive unsupervised datasets with a vast number of parameters.
-However, they still can not emulate human-like continuous learning due to catastrophic forgetting.
-Consequently, various continual learning (CL)-based methodologies have been developed to refine LMs, enabling them to adapt to new tasks without forgetting previous knowledge.
-However, a systematic taxonomy of existing approaches and a comparison of their performance are still lacking, which is the gap that our survey aims to fill.
-We delve into a comprehensive review, summarization, and classification of the existing literature on CL-based approaches applied to foundation language models, such as pre-trained language models (PLMs), large language models (LLMs) and vision-language models (VLMs).
-We divide these studies into offline CL and online CL, which consist of traditional methods, parameter-efficient-based methods, instruction tuning-based methods and continual pre-training methods.
-Offline CL encompasses domain-incremental learning, task-incremental learning, and class-incremental learning, while online CL is subdivided into hard task boundary and blurry task boundary settings.
+Despite these capabilities, LMs still struggle with catastrophic forgetting, hindering their ability to learn continuously like humans.
+To address this, continual learning (CL) methodologies have been introduced, allowing LMs to adapt to new tasks while retaining learned knowledge.
+However, a systematic taxonomy of existing approaches and a comparison of their performance are still lacking.
+In this paper, we delve into a comprehensive review, summarization, and classification of the existing literature on CL-based approaches applied to foundation language models, such as pre-trained language models (PLMs), large language models (LLMs) and vision-language models (VLMs).
+We divide these studies into offline and online CL, which consist of traditional methods, parameter-efficient-based methods, instruction tuning-based methods and continual pre-training methods.
 Additionally, we outline the typical datasets and metrics employed in CL research and provide a detailed analysis of the challenges and future work for LMs-based continual learning.
 Continual Learning, Foundation Language Models, Pre-trained Language Models, Large Language Models, Vision-Language Models, Survey
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Scene understandingccs: Computing methodologies Cognitive roboticsccs: Computing methodologies Cognitive scienceccs: Computing methodologies Intelligent agents
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
 1. Introduction
-Recent advancements in foundation language models (LMs) have set new benchmarks in both natural language processing (NLP) (Zhao et al., 2023; Min et al., 2023; Zhou et al., 2023a) and computer vision (CV) (Wang et al., 2023b).
+Recent advancements in foundation language models (LMs) have set new benchmarks in both natural language processing (NLP) (Zhao et al., 2023; Min et al., 2023; Zhou et al., 2023a) and computer vision (CV) (Wang et al., 2023b; Tiwari et al., 2022; Ramasesh et al., 2022).
 Foundation LMs encompass three primary categories: Pre-trained Language Models (PLMs) (Min et al., 2023), Large Language Models (LLMs) (Zhao et al., 2023), and Vision-Language Models (VLMs) (Du et al., 2022).
-PLMs such as BERT (Kenton and Toutanova, 2019), RoBERTa (Liu et al., 2019a), and BART (Lewis et al., 2020) focus on text-based tasks and are crucial for understanding and generating language by leveraging tasks like masked language modeling during pre-training.
-LLMs, including models like GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), extend the capabilities of PLMs by increasing the scale of model architecture and training data, thus enhancing their generality and adaptability across a broader range of tasks.
-VLMs, represented by VisualBERT (Li et al., 2019), CLIP (Radford et al., 2021), LLaVA (Liu et al., 2024) and DALL-E (Ramesh et al., 2021), integrate text and image modalities to enable complicated interactions between visual and textual information.
-The underlying paradigm of these models involves pre-training on extensive, often unlabeled datasets to capture rich semantic information, which is subsequently fine-tuned for specific tasks or domains.
-This methodology not only boosts performance across various applications but also significantly enhances the models’ flexibility and task adaptability.
+These models are pre-trained on large, unlabeled datasets to capture rich semantic information, which is then fine-tuned for specific tasks or domains.
+This strategy not only enhances performance across various applications but also significantly improves the flexibility and adaptability of these models.
+However, despite their strengths, foundation LMs face challenges in dynamic environments where tasks evolve over time.
+A key issue is “catastrophic forgetting” (Kirkpatrick et al., 2017), where models lose previously learned knowledge when adapting to new information.
+Unlike human learning, which is inherently continuous and adaptive (De Lange et al., 2021), foundation LMs generally require retraining to incorporate new data.
+Effective learning in such environments demands not only the ability to accelerate learning on new tasks (forward transfer) but also to improve performance on previous tasks by integrating newly acquired knowledge (backward transfer).
+While multi-task learning (MTL) and transfer learning (TL) offer potential solutions, MTL requires all task data to be available upfront, and TL focuses on limited tasks, making both approaches impractical for dynamic, real-world applications.
 Figure 1. Comparison between traditional CL and Foundation language models (LMs)-Based CL.
-However, these foundation models often demonstrate limitations in dynamic environments with a sequence of tasks, primarily due to their fixed parameters once training is completed.
-These models generally lack the capability to integrate new data or concepts without undergoing a retraining process.
-A significant challenge associated with training on a sequence of tasks is “catastrophic forgetting” (Kirkpatrick et al., 2017), a phenomenon where a model loses previously acquired knowledge upon learning new information.
-This is in stark contrast to human learning processes, which are inherently continuous and adaptive.
-Despite the successes of multi-task learning (MTL) and transfer learning (TL) in certain applications, they have limitations in real-world scenarios.
-MTL necessitates having all tasks and their data available upfront, which poses a challenge when launching a new service as the model must be retrained with all the data.
-Furthermore, TL is typically done with only two tasks, i.e., the source and the target, rendering it impractical for real-world online platforms with multiple target tasks.
-To address these challenges, it is crucial for models to process and learn the continuously expanding and diversifying datasets.
-This requires mechanisms that allow models to adapt to new linguistic phenomena and trends without compromising the accuracy and sensitivity towards historical data.
-Consequently, continual learning (CL) (Wang et al., 2024b; Van de Ven et al., 2022), also referred to as lifelong learning (Parisi et al., 2019) or incremental learning (Zhou et al., 2023b), is a crucial area in artificial intelligence that seeks to develop systems capable of continuously updating themselves and acquiring new knowledge, without forgetting previously learned information, similar to human learning (De Lange et al., 2021).
-This paradigm is particularly relevant in the context of foundation language models (LMs), which are challenged by specific issues such as catastrophic forgetting (CF) and cross-task knowledge transfer (KT).
-Catastrophic forgetting represents a significant challenge, where a model tends to lose previously acquired knowledge upon learning new information.
-To address this, language models must maintain a robust grasp of past language data while adapting to new linguistic trends.
-Furthermore, cross-task knowledge transfer is essential for enhancing the continual learning process.
-Effective KT not only accelerates the learning curve for new tasks (forward transfer) but also enhances the model’s performance on prior tasks via the feedback of new knowledge (backward transfer).
+Continual learning (CL) (Wang et al., 2024c; Van de Ven et al., 2022), also known as lifelong learning (Parisi et al., 2019) or incremental learning (Zhou et al., 2023b), offers an effective solution to these challenges.
+It aims to develop systems capable of continuously learning and updating without forgetting past knowledge.
+Recent advancements in CL methodologies have substantially enhanced the adaptability and knowledge retention capabilities of foundation LMs (Mehta et al., 2021; Cossu et al., 2022; Lee et al., 2023).
+Notable successes have been documented in diverse downstream tasks, such as aspect-based sentiment analysis (Ke et al., 2021b), dialogue generation (Scialom et al., 2022a), text classification (Razdaibiedina et al., 2023), visual question answering (Zhang et al., 2023d; Qian et al., 2023) and so on.
+Luo et al.
+(Luo et al., 2023) conduct an empirical study on catastrophic forgetting (CF) in large language models (LLMs) during continual instruction tuning.
+The aforementioned works underscore the potential of continual learning to significantly boost the performance of foundation LMs.
 {forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.4em,font=,, where level=2text width=8.4em,font=,, where level=3text width=10.0em,font=,, where level=4text width=10.5em,font=,, [ Foundation LMs-based CL, ver [ Offline Continual
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.4em,font=,, where level=2text width=8.4em,font=,, where level=3text width=10.5em,font=,, where level=4text width=10.5em,font=,, [ Foundation LMs-based CL, ver [ Offline Continual
 Learning  (§5) [ Domain-Incremental
 Learning  (§5.1) [ PLMs-based DIL  (§5.1.1) [ LFPT5 (Qin and Joty, 2022b), B-CL (Ke and Xu, 2021), ELLE (Qin et al., 2022), AdapterCL (Madotto et al., 2020), RMR_ DSE (Li et al., 2022a), DEMIX (Geng et al., 2021) , CLASSIC (Ke et al., 2021b),
-CPT (Ke et al., 2022), C-PT (Zhu et al., 2022), CL-KD (Castellucci et al., 2021), PlugLM (Cheng et al., 2023b), Pretr (Cossu et al., 2022), AEWC (Lee, 2017), Continual DAP-training (Ke et al., 2023) , leaf, text width=44.6em ] ] [ LLM-based DIL  (§4.1.2) [ COPR (Zhang et al., 2023b), LAMOL (Sun et al., 2020a), RVAE_LAMOL (Wang et al., 2022a), Adapt-Retrieve-Revise (Zhang et al., 2023c), Lifelong-MoE (Chen et al., 2023),
-CKL (Jang et al., 2022b), DACP (Xie et al., 2023), CPPO (Zhang et al., 2024), EcomGPT-CT (Ma et al., 2023) , leaf, text width=44.6em ] ] [ VLMs-based DIL  (§4.1.3) [ S-Prompt (Wang et al., 2022b), Medical AI (Yi et al., 2023), VQACL (Zhang et al., 2023d) , leaf, text width=44.6em ] ] ] [ Task-Incremental
-Learning  (§4.2) [ PLMs-based TIL  (§4.2.1) [ PP (Razdaibiedina et al., 2023), CTR (Ke et al., 2021a), MeLL (Wang et al., 2021a), LINC (Liu and Mazumder, 2021), ERDA (Qin and Joty, 2022a), PCLL (Zhao et al., 2022), CLIF (Jin et al., 2021a), ConTinTin (Yin et al., 2022),
-HMI (Maekawa et al., 2023), Adaptive Compositional Modules (Zhang et al., 2022b), DYNAINST (Mok et al., 2023), Conure (Yuan et al., 2021), TERACON (Kim et al., 2023),
-ERNIE 2.0 (Sun et al., 2020b), RecyclableTuning (Qin et al., 2023) , leaf, text width=44.6em ] ] [ LLMs-based TIL  (§4.2.2) [ Conpet (Song et al., 2023), InstructAlign (Cahyawijaya et al., 2023), Continual-T0 (Scialom et al., 2022a), DynaMind (Du et al., 2023), ELM (Jang et al., 2023), O-LoRA (Wang et al., 2023a), JARe (PENG et al., 2024) , leaf, text width=44.6em ] ] [ VLMs-based TIL  (§4.2.3) [ Medical AI (Yi et al., 2023), CTP (Zhu et al., 2023), ZSCL (Zheng et al., 2023), MoE-Adapters4CL (Yu et al., 2024), TRIPLET (Qian et al., 2023) , leaf, text width=44.6em ] ] ] [ Class-Incremental
+Pretr (Cossu et al., 2022), CPT (Ke et al., 2022), C-PT (Zhu et al., 2022), CL-KD (Castellucci et al., 2021), PlugLM (Cheng et al., 2023b), AEWC (Lee, 2017), DAS (Ke et al., 2023) , leaf, text width=44.6em ] ] [ LLM-based DIL  (§5.1.2) [ COPF (Zhang et al., 2023b), LAMOL (Sun et al., 2020a), RVAE_LAMOL (Wang et al., 2022a), Adapt-Retrieve-Revise (Zhang et al., 2023c), Lifelong-MoE (Chen et al., 2023),
+DACP (Xie et al., 2023), CPPO (Zhang et al., 2024b), EcomGPT-CT (Ma et al., 2023), LLM-CL (Ding et al., 2024), AMA (Lin et al., 2024) , leaf, text width=44.6em ] ] [ VLMs-based DIL  (§5.1.3) [ S-Prompt (Wang et al., 2022b), VQACL (Zhang et al., 2023d), SC-MLLM (Liu et al., 2024), DIKI (Tang et al., 2024) , leaf, text width=44.6em ] ] ] [ Task-Incremental
+Learning  (§5.2) [ PLMs-based TIL  (§5.2.1) [ PP (Razdaibiedina et al., 2023), CTR (Ke et al., 2021a), MeLL (Wang et al., 2021), LINC (Liu and Mazumder, 2021), ERDA (Qin and Joty, 2022a), PCLL (Zhao et al., 2022), BiHNet-Reg (Jin et al., 2021), ConTinTin (Yin et al., 2022),
+HMI (Maekawa et al., 2023), ACM (Zhang et al., 2022), DYNAINST (Mok et al., 2023), Conure (Yuan et al., 2021), TERACON (Kim et al., 2023),
+EMR (Wang et al., 2019b), ERNIE 2.0 (Sun et al., 2020b), RecyclableTuning (Qin et al., 2023), RecAdam (Chen et al., 2020) , leaf, text width=44.6em ] ] [ LLMs-based TIL  (§5.2.2) [ Conpet (Song et al., 2023), InstructAlign (Cahyawijaya et al., 2023), Continual-T0 (Scialom et al., 2022b), DynaMind (Du et al., 2023), ELM (Jang et al., 2023), O-LoRA (Wang et al., 2023a), JARe (PENG et al., 2024),
+Robocoder (Li et al., 2024), Eureka (Ma et al., 2024), COPF (Zhang et al., 2024a) , leaf, text width=44.6em ] ] [ VLMs-based TIL  (§5.2.3) [ Medical AI (Yi et al., 2023), CTP (Zhu et al., 2023), ZSCL (Zheng et al., 2023), MoE-Adapters4CL (Yu et al., 2024b), TRIPLET (Qian et al., 2023), AwoForget (Zheng et al., [n.
+d.]),
+SND (Yu et al., 2024a) , leaf, text width=44.6em ] ] ] [ Class-Incremental
 Learning  (§5.3) [ PLMs-based CIL  (§5.3.1) [ EPI (Wang et al., 2023c), IDBR (Huang et al., 2021), PAGeR (Varshney et al., 2022), ENTAILMENT (Xia et al., 2021), ExtendNER (Monaikul et al., 2021), PLE (Li et al., 2022b), DE&E (Wójcik et al., 2023),
-SRC (Liu et al., 2019b) , leaf, text width=44.6em ] ] [ VLMs-based CIL  (§4.3.2) [ MoE-Adapters4CL (Yu et al., 2024), VLM-PL (Kim et al., 2024), Adaptation-CLIP (Liu et al., 2023a), PROOF (Zhou et al., 2023c), LGCL (Khan et al., 2023), ZSCL (Zheng et al., 2023),
-CLAP (Jha et al., 2024), GMM (Cao et al., 2024) , leaf, text width=44.6em ] ] ] ] [ Online Continual
+SRC (Liu et al., 2019b) , leaf, text width=44.6em ] ] [ VLMs-based CIL  (§5.3.2) [ MoE-Adapters4CL (Yu et al., 2024b), VLM-PL (Kim et al., 2024), Adaptation-CLIP (Liu et al., 2023a), PROOF (Zhou et al., 2023c), LGCL (Khan et al., 2023), ZSCL (Zheng et al., 2023),
+CLAP (Jha et al., 2024), GMM (Cao et al., 2024), RAPF (Huang et al., 2024), STAR-Prompt (Menabue et al., 2024), DIKI (Tang et al., 2024), SND (Yu et al., 2024a), AwoForget (Zheng et al., [n.
+d.]) , leaf, text width=44.6em ] ] ] ] [ Online Continual
 Learning  (§6) [ Hard Task
-Boundary  (§5.1) [ PLMs-based HTB  (§5.1.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b), ProgModel (Shen et al., 2019) , leaf, text width=44.6em ] ] ] [ Blurry Task
-Boundary  (§5.2) [ PLMs-based BTB  (§5.2.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b) , leaf, text width=44.6em ] ] [ VLMs-based BTB  (§5.2.2) [ CBA (Wang et al., 2023d), MVP (Moon et al., 2023), DKR (Cui et al., 2024) , leaf, text width=44.6em ] ] ] ] ]
+Boundary  (§6.1) [ PLMs-based HTB  (§6.1.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b), ProgModel (Shen et al., 2019) , leaf, text width=44.6em ] ] [ VLMs-based HTB  (§6.1.2) [ PEGP (Qiao et al., 2024) , leaf, text width=44.6em ] ] ] [ Blurry Task
+Boundary  (§6.2) [ PLMs-based BTB  (§6.2.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b), S6 (Li et al., 2023) , leaf, text width=44.6em ] ] [ VLMs-based BTB  (§6.2.2) [ DKR (Cui et al., 2024), SIT (Wang et al., 2024a), OLiVia-Nav (Narasimhan et al., 2024), G-NoCL (Seo et al., 2024) , leaf, text width=44.6em ] ] ] ] ]
 Figure 2. Taxonomy of foundation language models for continual learning.
-Recent advancements in continual learning methodologies have substantially enhanced the adaptability and knowledge retention capabilities of foundational language models (LMs).
-These developments are crucial for addressing complex challenges previously observed in CL.
-Researchers have formulated innovative strategies to mitigate these challenges, thereby enabling LMs to maintain high performance across a variety of tasks while continually integrating new knowledge (Mehta et al., 2021; Cossu et al., 2022; Lee et al., 2023).
-Notable successes have been documented in diverse downstream tasks, such as aspect-based sentiment analysis, where continual learning enables dynamic adaptation to evolving aspects and sentiments (Ke et al., 2021b).
-Similarly, in dialogue generation, the novel technologies assist models in refining and expanding their conversational capabilities through ongoing interactions (Scialom et al., 2022b).
-In text classification, continual learning facilitates the incorporation of new categories and adjustments to shifts in text distributions without the need for complete retraining (Razdaibiedina et al., 2023).
-Moreover, in the realm of visual question answering, continual learning is essential for updating the models’ capabilities to process and respond to new types of visual content and queries (Zhang et al., 2023d; Qian et al., 2023).
-The aforementioned works underscore the potential of continual learning to significantly boost the performance of foundation LMs.
-In the domain of continual learning, there has been a significant paradigm shift from traditional methodologies to those that integrate foundation LMs (See Figure 1).
-First, foundation LMs demonstrate enhanced generalization and transfer learning abilities across diverse tasks owing to their broad pre-training on large-scale datasets.
-The model has specialized transfer capability to quickly adapt to downstream tasks with only a few samples.
-Consequently, it is crucial to mitigate the degradation of both the zero-shot transfer and history task abilities in LMs while facilitating the acquisition of new skills.
+In the domain of continual learning, there has been a shift from traditional methods to those incorporating foundation LMs (Figure 1).
+First, foundation LMs has specialized transfer capability to quickly adapt to downstream tasks with only a few samples.
+Consequently, it is crucial to mitigate the degradation of both the zero-shot transfer and history task abilities while facilitating the acquisition of new skills.
 Second, due to the substantial number of parameters in foundation LMs, it is crucial to employ parameter-efficient techniques (Han et al., 2024), such as prompt tuning (Liu et al., 2022) and adapters (Mundra et al., 2024), to update parameters without comprehensive retraining.
 Third, the foundation LMs possess the capability to follow instructions through instructional learning (Dong et al., 2022; Ouyang et al., 2022), enabling more dynamic and context-aware interactions.
-This review systematically categorizes these strategies and technologies into two core areas: offline continual learning and online continual learning (Figure 2).
-We first give detailed definitions and scenarios to format the setting of offline and online CL, where offline CL includes domain-incremental, task-incremental and class-incremental CL, and online CL includes hard task boundary and blurry task boundary.
-These learning strategies are further subdivided into methods based on Pre-trained Language Models (PLMs), Large Language Models (LLMs), and Vision-Language Models (VLMs).
-Then, we summarize the related papers about traditional methods, continual pre-training methods, parameter-efficient tuning methods and instruction-based methods.
-Finally, we static the main datasets from various perspectives and review the key metrics to evaluate the forgetting and transferring of the models.
-The main contributions of this survey paper can be summarized as follows.
+This review systematically organizes continual learning strategies and technologies into two main categories: offline continual learning and online continual learning (Figure 2).
+We begin by defining and explaining the different settings for these two types of continual learning.
+Offline continual learning includes domain/task/class-incremental CL, while online continual learning is further divided into methods that address hard task boundaries and those that manage blurry task boundaries (Section 4).
+To further clarify the relationship between these strategies and model architectures, we group the methods based on three major model types: Pre-trained Language Models (PLMs), Large Language Models (LLMs), and Vision-Language Models (VLMs).
+These categories reflect the distinct requirements for continual learning in each architecture and emphasize the role these models play in shaping learning strategies.
+Additionally, we classify the methods into four key approaches: traditional continual learning methods, continual pre-training methods, parameter-efficient tuning methods, and instruction-based methods (Section 5, 6).
+This categorization highlights the different techniques used to address challenges like catastrophic forgetting and to improve knowledge transfer in continual learning scenarios.
+Finally, we static the main datasets from various perspectives (Section7) and review the key metrics to evaluate the forgetting and transferring of the models (Section 8).
+The main contributions of this survey paper can be summarized as follows:
   • We thoroughly review the existing literature on foundation LMs-based CL approaches, which integrate foundation LMs with CL to learn new knowledge without retraining the models.
@@
   • We provide the characters of existing datasets for CL and present the main metrics to evaluate the performance of preventing forgetting and knowledge transfer.
-  • We discuss the most challenging problems of foundation LMs-based CL and point out promising future research directions in this field.
-The paper is organized as follows.
-In Section 2, we review the mainly related surveys about continual learning.
-Then, we introduce the base settings and learning modes of continual learning in Section 3, including the definitions and scenarios of CL.
-Furthermore, we present the related studies about offline continual learning, which can be divided into domain-incremental learning, task-incremental learning and class-incremental learning in Section 4.
-In Section 5, we focus on online continual learning, including hard task boundary and blurry task boundary settings.
-The typical datasets and metrics are provided in Section 6 and 7.
-Finally, we analyze the challenge and further work in Section 8 and give the conclusion in Section 9.
+Furthermore, we discuss the most challenging problems of foundation LMs-based CL and point out promising future research directions in this field.
 2. Related Surveys
 2.1. Continual Learning
-Early examinations of Continual Learning (CL) have provided broad coverage, as observed in surveys such as Parisi et al.
+Early examinations have provided broad coverage, as observed in surveys such as Parisi et al.
 (Parisi et al., 2019).
 Recently, Wang et al.
-(Wang et al., 2024b) conduct a comprehensive survey that categorizes five key strategies in CL: regularization-based, replay-based, optimization-based, representation-based, and architecture-based approaches.
+(Wang et al., 2024c) conduct a comprehensive survey that categorizes five key strategies in CL: regularization-based, replay-based, optimization-based, representation-based, and architecture-based approaches.
 This survey reflects an effort to organize and understand the diverse methodologies employed in the field.
-Notably, there is a growing focus on class-incremental setting (Zhou et al., 2023b; Masana et al., 2022; Belouadah et al., 2021) and replay-based approaches (Hayes et al., 2021), reflecting the increasing granularity of research interests within the CL domain.
+Notably, there is also a growing focus on class-incremental setting (Zhou et al., 2023b; Masana et al., 2022; Belouadah et al., 2021) and replay-based approaches (Hayes et al., 2021).
+Recent surveys (Wu et al., 2024; Shi et al., 2024; Zheng et al., 2024) provide comprehensive overviews of continual pre-training, continual domain-adaptive pre-training, continual instruction tuning, and continual alignment.
+In particular, Shi et al.
+(Shi et al., 2024) provides an in-depth overview of CL, specifically in LLMs, introducing vertical CL (adapting from general to specific domains) and horizontal CL (adapting across time and domains).
+Additionally, one notable approach in continual alignment is Reinforcement Learning from Human Feedback (RLHF) (Wu et al., 2022), which incorporates human feedback to optimize model outputs for better task adaptation.
 2.2. Continual Learning for Computer Vision
-In the realm of computer vision, De et al.
-(De Lange et al., 2021) address the pressing challenge of continual learning, specifically in the task-incremental setting, where tasks arrive sequentially with clear boundaries.
-They introduce a novel stability-plasticity trade-off framework tailored for continual learners and undertake a comprehensive experimental analysis, comparing the efficacy of 11 methodologies across three benchmarks.
+In the realm of computer vision (CV), De et al.
+(De Lange et al., 2021) concentrate on task incremental classification and offer significant contributions including a taxonomy, a new framework for balancing stability and plasticity, and an extensive comparison of continual learning methods against baselines.
 Qu et al.
 (Qu et al., 2021) present a comprehensive examination of continual learning, highlighting its vital role in the accumulation of knowledge from sequential data streams.
-This research investigates a range of approaches, encompassing regularization, knowledge distillation, memory-based methods, and more.
-These approaches are systematically categorized by their characteristics and applications in computer vision.
+This research explores multiple methods including regularization, knowledge distillation, memory-based approaches, and more, categorizing them by their characteristics and applications in CV.
 Moreover, Mai et al.
-(Mai et al., 2022) focus on the realm of online continual learning within image classification, addressing catastrophic forgetting.
-This study evaluates the efficacy of state-of-the-art methodologies across diverse memory and data configurations.
+(Mai et al., 2022) focus on the realm of online continual learning within the image classification task.
+This study evaluates the efficacy of state-of-the-art methods across diverse memory and data configurations.
 Masana et al.
-(Masana et al., 2022) present a comprehensive performance evaluation of class-incremental methods applied to image classification.
-The experiments span a range of scenarios, incorporating large-scale datasets and varied network architectures.
+(Masana et al., 2022) conduct a comprehensive evaluation of class-incremental methods for image classification, involving large-scale datasets and various network architectures.
 Belouadah et al.
-(Belouadah et al., 2021) pay more attention to class-incremental learning algorithms for visual tasks.
-This study defines the essential properties of incremental learning algorithms, offers a unified formalization of the class-incremental learning problem, and provides an evaluation framework for detailed analysis.
+(Belouadah et al., 2021) focus more on class-incremental learning algorithms specifically for visual tasks, defining key properties of these algorithms, formalizing the class-incremental learning problem, and providing an evaluation framework for thorough analysis.
 2.3. Continual Learning for NLP
@@
 (Biesialska et al., 2020) address the challenge of continual learning within Natural Language Processing (NLP), wherein conventional architectures struggle to accommodate new tasks without compromising previously acquired knowledge.
-The research presents an extensive review of methods, including rehearsal, regularization, and architectural approaches, all designed to mitigate the aforementioned challenge.
 In a similar vein, Ke et al.
@@
 This work presents an in-depth analysis of state-of-the-art approaches and extends original CL settings to be more general and up-to-date.
-Additionally, it emphasizes the significance of knowledge transfer within NLP and the challenges posed by inter-task class separation.
+Gogoulou et al.
+(Gogoulou et al., 2023) study the pros and cons of updating a language model when new data comes from new languages – the case of CL under language shift.
+They feed various languages into the model to examine the impact of pre-training sequence and linguistic characteristics on both forward and backward transfer effects across three distinct model sizes.
 2.4. Continual Learning for Other Domains
-Recent surveys, surveys like (Zhang and Kim, 2023; Shaheen et al., 2022; Lesort et al., 2020) investigate the advancements in incremental learning for neural recommendation systems and continual learning (CL) in robotics, respectively.
+Recent surveys, surveys like (Zhang and Kim, 2023; Shaheen et al., 2022; Lesort et al., 2020) explore advancements in incremental learning across other domains.
 Zhang et al.
-(Zhang and Kim, 2023) make a notable contribution to narrow the gap between academic research and industrial applications through the introduced Incremental Update Recommendation Systems (IURS).
-They highlight the imperative for real-time updates with streaming data and focus on the distinctive challenges posed by IURS in contrast to traditional Batch Update Recommendation Systems (BURS) and offer a thorough examination of existing literature and evaluation methodologies in this domain.
+(Zhang and Kim, 2023) focus on neural recommendation systems, particularly introducing the Incremental Update Recommendation Systems (IURS) to bridge the gap between academic research and industrial applications.
+They discuss the unique challenges of IURS compared to traditional Batch Update Recommendation Systems (BURS) and offer a detailed review of existing literature and evaluation methods.
 Shaheen et al.
-(Shaheen et al., 2022) provide a comprehensive overview of contemporary approaches for CL within real-world contexts.
-Their analysis focuses on learning algorithms that efficiently handle large sequential datasets within computational and memory constraints.
-The survey also explores challenges in applying CL to autonomous systems, comparing methods across metrics such as computational efficiency, memory utilization, and network complexity.
-In the field of robotics, it is essential for agents to adapt and interact with their environment using a continuous stream of observations.
-Thus, Lesort et al.
-(Lesort et al., 2020) explore CL within this domain, defining CL as a paradigm where both data distribution and learning objectives evolve dynamically.
-They emphasize the challenges in evaluating CL algorithms in robotic applications and introduce a novel framework alongside metrics tailored to effectively present and assess CL methodologies.
-This paper centers on the crucial advancements in CL as applied to foundational language models, which have obtained significant success in the fields of NLP and multimodal.
+(Shaheen et al., 2022) offer a comprehensive overview of the CL methods in real-world contexts, highlighting efficient learning algorithms for handling large sequential datasets within resource constraints and exploring the applicability of CL to autonomous systems.
+Lesort et al.
+(Lesort et al., 2020) investigate CL in robotics, defining it as a dynamic learning paradigm where both data distribution and objectives evolve.
+They also emphasize the challenges in evaluating CL algorithms and introduce a new framework with tailored metrics for assessing CL methods.
+Continual learning across domains aims to mitigate catastrophic forgetting by integrating new tasks while preserving knowledge from prior ones (Monaikul et al., 2021; Li et al., 2022b).
+However, domains like NLP, CV, and robotics exhibit significant differences.
+CV methods (Wang et al., 2022d, e) typically focus on classification, while NLP involves more complex tasks such as information extraction and text generation.
+Robotics tasks evolve sequentially, relying on decision-making and state transitions (Kirkpatrick et al., 2017; Aljundi et al., 2018).
+NLP models, primarily based on PLMs or LLMs, excel in few-shot and transfer learning, whereas CV tasks commonly utilize traditional neural networks (e.g., ResNet).
+Recently, Vision-Language Models (VLMs) have been introduced in CV to enhance continual learning.
+Additionally, the pre-trained transformers in NLP enables parameter-efficient strategies like adapter and prompt tuning, allowing adaptation to new tasks with minimal retraining (Ke et al., 2022; Madotto et al., 2020).
+This paper centers on the advancements in CL as applied to foundation LMs, which have obtained success in the fields of NLP and multimodal.
 We categorize existing works into offline and online CL based on PLMs, LLMs, and VLMs.
@@
 The samples with different classes (domains) are marked with various shapes (colors).
+3. Typical Foundation Language Model
+Pre-trained Language Models (PLMs).
+PLMs are trained on large-scale textual data using techniques such as Masked Language Modeling (MLM) and Next Sentence Prediction (NSP) for various tasks, including text classification, sentiment analysis, and named entity recognition.
+Notable examples of PLMs include ELMo (Sarzynska-Wawer et al., 2021), BERT (Devlin, 2018), early versions of the GPT series (Radford et al., 2018), and RoBERTa (Liu et al., 2019a).
+PLM-based methods aim to mitigate catastrophic forgetting when learning new tasks or domains within a closed-world setting.
+Continual learning for PLMs focuses on parameter-efficient tuning and replay-based strategies to reduce resource consumption and prevent catastrophic forgetting.
+Approaches like adapter-based tuning, prompt-based methods, and knowledge distillation are popular because they update the model with few parameters.
+This ensures that PLMs remain feasible to deploy on standard hardware while keeping computational costs low.
+Large Language Models (LLMs).
+LLMs are significantly larger in scale than PLMs, often containing billions or even trillions of parameters.
+These models are trained on vast datasets using language modeling techniques and excel in tasks such as generation, instruction learning, and in-context learning.
+LLMs are typically exposed to a broader range of complex tasks, from domain-specific corpora to open-domain challenges.
+Prominent examples include OpenAI’s GPT-3 (Brown, 2020), GPT-4 (Achiam et al., 2023), LLaMA (Touvron et al., 2023), and others.
+Due to the substantial memory and computational demands associated with LLMs, continual learning strategies for LLMs often focus on optimizing memory efficiency (e.g., orthogonal low-rank adaptation, recyclable tuning) and selectively updating parameters (e.g., Mixture-of-Experts, dynamic task selection), as retraining or fine-tuning the entire model is typically cost-prohibitive.
+Additionally, LLMs require techniques to mitigate catastrophic forgetting while managing task diversity and retaining foundation knowledge.
+Vision-Language Models (VLMs).
+VLMs are designed to process and integrate multimodal information, such as text and vision, by building on PLMs or LLMs for tasks like image captioning, visual question answering, and visual reasoning.
+Notable examples include CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021).
+VLMs must effectively handle multimodal data, which necessitates continual learning methods that can efficiently manage cross-modal learning.
+Continual learning for VLMs often emphasizes alignment across multiple modalities using specialized modules such as cross-modal fusion techniques, attention mechanisms, and lightweight adapters.
+Additionally, methods that focus on retaining modality-specific knowledge and enabling parameter-efficient fine-tuning (e.g., Adaptation-CLIP (Liu et al., 2023a), MoE-Adapters4CL (Yu et al., 2024b)) are essential due to the high costs associated with continual learning across both vision and language domains.
 4. Settings and Learning Modes of CL
@@
 4.2.1. Offline Continual Learning
-Offline CL (Figure 3) comprises three principal scenarios, each distinguished by distinct characteristics: Domain-Incremental Learning, Task-Incremental Learning, and Class-Incremental Learning.
+Offline CL (Figure 3) comprises three principal scenarios, each distinguished by distinct characteristics: Domain-Incremental Learning (e.g., LFPT5 (Qin and Joty, 2022b), LAMOL (Sun et al., 2020a), S-Prompt (Wang et al., 2022b)), Task-Incremental Learning (e.g., PP (Razdaibiedina et al., 2023), Conpet (Song et al., 2023), ZSCL (Zheng et al., 2023)), and Class-Incremental Learning (e.g., EPI (Wang et al., 2023c), IDBR (Huang et al., 2021), GMM (Cao et al., 2024)).
   • Domain-Incremental Learning (DIL): The model aims to process diverse data distributions.
@@
 Moreover, the task IDs are only available during training.
-In summary, Domain-Incremental Learning concentrates on adapting the model to the shifts in input data distributions while maintaining consistency in tasks and classes.
-Task-Incremental Learning necessitates the model’s ability to learn and retain task-specific knowledge over successive tasks.
-On the other hand, Class-Incremental Learning highlights the gradual integration of new classes into the model’s recognition capabilities without compromising knowledge of previously learned classes.
+In summary, DIL concentrates on adapting the model to the shifts in input data distributions while maintaining consistency in tasks and classes.
+TIL necessitates the model’s ability to learn and retain task-specific knowledge over successive tasks.
+On the other hand, CIL highlights the gradual integration of new classes into the model’s recognition capabilities without compromising knowledge of previously learned classes.
 4.2.2. Online Continual Learning
-In online continual learning (See Figure 4), existing research is categorized into two configurations based on the arrival pattern of tasks: ”Hard Task Boundary” and ”Blurry Task Boundary”:
+In online continual learning (Figure 4), the existing researches are categorized into two configurations based on the arrival pattern of tasks: ”Hard Task Boundary” (e.g., MBPA++ (de Masson D’Autume et al., 2019), PEGP (Qiao et al., 2024)) and ”Blurry Task Boundary” (e.g., SIT (Wang et al., 2024a), G-NoCL (Seo et al., 2024)):
   • Hard Task Boundary: The arrival of tasks follows a strictly structured and sequential process.
@@
 Traditional Methods.
-Continual Learning methodologies are frequently employed in the context of Pre-trained Language Models (PLMs), encompassing approaches such as replay-based, regularization-based, and parameter-isolation-based algorithms.
-Li et al.
-(Li et al., 2022a) present a regularization-centric framework for Lifelong Learning (LLL) termed RMR-DSE, specifically tailored for sequential operation across multiple domains.
-Unlike conventional strategies requiring incremental memory allocation, RMR-DSE employs a recall optimization mechanism, driven by regularization, to selectively retain important parameters from preceding tasks.
-Furthermore, it incorporates a domain drift estimation algorithm to address embedding space shifts.
-Castellucci et al.
-(Castellucci et al., 2021) propose a Knowledge Distillation-based Continual Learning (CL-KD) approach with a Teacher-Student framework.
-When the student model is trained on a new language, the teacher model also transfers its knowledge of supported languages to the student model.
-Lee et al.
-(Lee, 2017) introduce a domain-agnostic framework that employs a two-phase training process, initially using synthetic data to learn general conversational patterns and subsequently using human-computer dialogs in customer support.
-Moreover, the Adaptive Elastic Weight Consolidation algorithm is applied to adjust the loss function, ensuring the balance between acquiring new knowledge and retaining previously learned information.
-Gururangan et al.
-(Gururangan et al., 2022) develop a parameter-isolation-based method, named DEMIX layers, which consists of a collection of experts.
-The dynamic addition or removal of these experts can enhance the model’s ability to adapt to new domains while maintaining robust performance in previously established ones.
-PlugLM (Cheng et al., 2023b) is a pre-training model equipped with a differentiable plug-in memory (DPM) designed for domain-adaptive continual training.
-The core concept behind this approach is to separate the knowledge storage from model parameters using an adaptable key-value memory structure.
-It enables the explicit retrieval and utilization of knowledge stored within the DPM.
-(a) CLASSIC
+Continual Learning methods are increasingly used with Pre-trained Language Models (PLMs), adopting various techniques such as replay, regularization, and parameter-isolation.
+Driven by regularization, RMR_DSE (Li et al., 2022a) optimizes recall and uses domain drift estimation to manage embedding space shifts across domains, effectively reducing catastrophic forgetting in seq2seq tasks under memory constraints.
+CL-KD (Castellucci et al., 2021) uses a Teacher-Student model to transfer language knowledge efficiently.
+AEWC (Lee, 2017) is an improved EWC algorithm, where the importance can be computed in an online manner.
+DEMIX (Gururangan et al., 2022) is a parameter-isolation approach, utilizing a collection of domain-specific experts for dynamic domain adaptation.
+It offers superior performance in both in-domain and out-of-domain tasks, providing strong generalization to unseen domains and promising improved adaptability and scalability in multi-domain settings.
+Lastly, PlugLM (Cheng et al., 2023b) (Figure 5(a)) features a differentiable plug-in memory (DPM) for domain-adaptive training, with a flexible key-value structure that decouples knowledge storage from model parameters, enabling more effective knowledge management in domain-adaptive training.
+(a) PlugLM
 (b) Lifelong-MoE
 (c) S-Prompts
-Figure 5. Frameworks in DIL: CLASSIC (PLM-based) (Ke et al., 2021b), Lifelong-MoE (LLM-based) (Chen et al., 2023), S-Prompts (VLM-based) (Wang et al., 2022b).
+Figure 5. Frameworks in DIL: PlugLM (PLM-based) (Cheng et al., 2023b), Lifelong-MoE (LLM-based) (Chen et al., 2023), S-Prompts (VLM-based) (Wang et al., 2022b).
 Continual Pre-training Methods.
-Continual domain-adaptive pre-training (DAP-training) (Ke et al., 2023) is based on two main ideas: (1) the general knowledge in the LM and the knowledge gained from prior domains are crucial to mitigating CF and enhancing cross-task knowledge transfer.
-This is achieved through soft-masking units based on their importance, and (2) the model is designed to develop complementary representations of both the current domain and prior domains, thereby facilitating the integration of knowledge.
-The key novelty of Continual DAP-training is a soft-masking mechanism that directly controls the update to the LM.
-Cossu et al.
-(Cossu et al., 2022) formalize and explore the dynamics of continual pre-training (Pretr) scenarios across language and vision domains.
-In this framework, models undergo continuous pre-training on a sequential stream of data before subsequent fine-tuning for various downstream tasks.
+Pretr (Cossu et al., 2022) formalizes continual pre-training, where models are continuously pre-trained on sequential data streams before fine-tuning for downstream tasks.
+DAS (Ke et al., 2023) offers an effective solution for continual domain-adaptive pre-training (DAP-training) of LMs.
+By integrating soft-masking, KL-divergence loss, and contrastive learning, it mitigates catastrophic forgetting and enhances knowledge transfer between domains.
 Parameter-Efficient Tuning Methods.
@@
 The adapter architecture incorporates a skip-connection to minimize the number of parameters.
-A notable exemplar of this approach is AdapterCL (Madotto et al., 2020), which employs residual adapters tailored specifically for task-oriented dialogue systems.
-This framework, comprising 37 domains, is structured to facilitate continual learning across four important aspects: intent recognition, state tracking, natural language generation, and end-to-end processing.
-In a related vein, Ke et al.
-(Ke and Xu, 2021) introduce B-CL model to tackle critical challenges in CL for Aspect-Based Sentiment Classification.
-B-CL integrates continual learning adapters within capsule network architectures.
-Aiming at mitigating catastrophic forgetting during fine-tuning, the CLASSIC model (Ke et al., 2021b) presents an innovative solution by deploying adapters to tap into BERT’s capabilities (Figure 5(a)).
-A novel contrastive continual learning strategy is used to facilitate the transfer of knowledge across tasks and distill insights from previous tasks to subsequent ones.
-It also effectively eliminates the necessity for task identifiers during testing.
-Furthermore, Continual PostTraining (CPT) (Ke et al., 2022) introduces two continual learning plug-in modules, termed CL-plugins, embedded within each transformer layer of RoBERTa.
-Prompt tuning (Liu et al., 2022), or P-tuning, introduces trainable continuous prompts into the sequence of input word embeddings, while the language model remains frozen.
-To tackle the challenge of CL under limited labeled data, Qin et al.
-(Qin and Joty, 2022b) propose a Lifelong Few-shot Language Learning framework (LFPT5).
-In this framework, prompt tuning, replay and regularization strategies are leveraged.
-When presented with a new task, the model generates pseudo-labeled samples representative of prior domains.
-The training process then incorporates these pseudo-labeled samples alongside new task-specific data.
-Additionally, the KL divergence loss is employed to maintain label consistency between the previous and the current model.
-Furthermore, Zhu et al.
-(Zhu et al., 2022) introduced Continual Prompt Tuning (C-PT) as a methodology to address the challenges of continual learning within dialogue systems.
-C-PT facilitates knowledge transfer between tasks through continual prompt initialization, query fusion, memory replay, and a memory-guided technique.
+A notable exemplar of this approach is AdapterCL (Madotto et al., 2020), which employs residual adapters tailored for task-oriented dialogue systems.
+This makes it ideal for real-world deployment with frequent domain updates without the need for full retraining.
+In a related vein, B-CL (Ke and Xu, 2021) sets a new standard in Aspect-Based Sentiment Classification by utilizing CL adapters within capsule network architectures, demonstrating that backward knowledge transfer significantly boosts performance while preserving task-specific information.
+CLASSIC (Ke et al., 2021b) addresses catastrophic forgetting by leveraging adapters in BERT, employing a contrastive CL strategy that facilitates knowledge transfer across tasks without requiring task identifiers during testing.
+Furthermore, CPT (Ke et al., 2022) incorporates CL-plugins in RoBERTa, effectively improving end-task performance by managing domain-specific knowledge transfer, particularly in scenarios with limited labeled data.
+Prompt tuning (Liu et al., 2022), or P-tuning, introduces trainable prompts into the sequence of input word embeddings, while the LM remains frozen.
+LFPT5 (Qin and Joty, 2022b) integrates prompt tuning, pseudo-labeled samples, and KL divergence to excel in lifelong few-shot language learning (LFLL).
+The use of pseudo-labeled samples and KL divergence helps maintain label consistency between models, though task transferability requires careful attention.
+Similarly, C-PT (Zhu et al., 2022) enhances knowledge transfer between tasks in dialogue systems through prompt initialization, query fusion, and memory replay, outperforming baselines like EWC, Replay, and AdapterCL, while being efficient in parameter usage and memory.
 Instruction Tuning-based Methods.
 Instruction tuning-based methods involve transforming a given task into natural language instructions.
-Qin et al.
-(Qin et al., 2022) propose ELLE, a novel approach aimed at effectively incorporating continuously expanding streaming data into pre-trained language models (PLMs).
-It consists of two fundamental components: (1) function-preserved model expansion, which enhances knowledge acquisition efficiency by changing the width and depth of an existing PLM, and (2) pre-trained domain prompts, which significantly enhance the adaptation for downstream tasks by effectively segregating the diverse knowledge acquired during pre-training phases.
+ELLE (Qin et al., 2022) incorporates expanding streaming data into PLMs with two key components: (1) function-preserved model expansion, which enhances knowledge acquisition efficiency by changing the width and depth of PLM, and (2) pre-trained domain prompts, which enhance downstream task adaptation by utilizing domain-specific knowledge from the pre-training phase.
 5.1.2. LLMs-based DIL
 Traditional Methods.
-In many practical scenarios, retraining Language Models (LMs) is challenging due to resource constraints and data privacy concerns.
-Zhang et al.
-(Zhang et al., 2024) introduce Continual Proximal Policy Optimization (CPPO) to address this issue.
-CPPO integrates sample-wise weighting into the Proximal Policy Optimization (PPO) algorithm, effectively balancing policy learning and knowledge retention.
-Zhang et al.
-(Zhang et al., 2023b) propose Continual Optimal Policy Regularization (COPR), which calculates the optimal policy distribution without the partition function and uses the previous optimal policy to regularize the current policy.
-Sun et al.
-(Sun et al., 2020a) introduce LAMOL, which generates pseudo-samples from previous tasks while training on a new task.
-It effectively mitigates knowledge loss without requiring additional memory or computational resources.
-Building on this framework, Wang et al.
-(Wang et al., 2022a) developed RVAE_LAMOL, which integrates a residual variational autoencoder (RVAE) to encode input data into a unified semantic space, thereby enhancing task representation.
-This model also incorporates an identity task to enhance the model’s discriminative ability for task identification.
-To enhance training efficacy, the Alternate Lag Training (ALT) is devised to segment the training process into multiple phases.
-To reduce hallucinations in specialized domains such as the Chinese legal domain, Zhang et al.
-(Zhang et al., 2023c) propose a novel domain adaptation framework, named Adapt-Retrieve-Revise (ARR).
+In Lifelong Language Learning (LLL), a key challenge is training models on sequential NLP tasks while preserving knowledge from previous tasks.
+LAMOL (Sun et al., 2020a) tackles this by generating pseudo-samples from past tasks during new task training, effectively reducing knowledge loss without additional memory or computational overhead.
+RVAE_LAMOL (Wang et al., 2022a) builds on this by utilizing a variational autoencoder (RVAE) to map tasks into a unified semantic space, while introducing Alternate Lag Training (ALT) and an identity task to improve stability and task-specific sample generation.
+Efficiently and stably aligning with dynamic human preferences is challenging, especially when minimizing retraining costs.
+CPPO (Zhang et al., 2024b) addresses this issue by integrating sample-wise weighting into the PPO algorithm, effectively balancing policy learning and knowledge retention, particularly in adaptive tasks like summarization and question-answering.
+Similarly, COPF (Zhang et al., 2023b) reduces reliance on human feedback by regularizing the current policy with past optimal policies.
+It is well-suited for handling unlabeled data, making it ideal for real-world applications with limited labeled datasets.
+Furthermore, AMA (Lin et al., 2024) addresses the ”alignment tax”—the trade-off between maximizing alignment rewards and mitigating forgetting—by interpolating between pre- and post-RLHF model weights.
+This interpolation identifies optimal layer-wise weight combinations, balancing ”alignment tax” across various RLHF algorithms.
+Continual Pre-training Methods
+Large language models (LLMs) excel in open-domain tasks but face significant challenges in domain-specific applications, such as insufficient domain knowledge, limited capacity to leverage that knowledge, and inadequate adaptation to specialized data formats.
+To overcome these limitations, continual pre-training (Xie et al., 2023; Ma et al., 2023; Cheng et al., 2023a) has emerged as a promising approach, enabling LLMs to better fit domain-specific tasks.
+For example, AdaptLLM (Cheng et al., 2023a) transforms raw corpora into reading comprehension texts, but it was found that while pre-training on raw data enhances domain knowledge, it hampers the model’s ability to answer questions effectively.
+As an alternative to full retraining, Domain-Adaptive Continual Pre-training (DACP) leverages large domain-specific corpora to continually train models, although it involves high computational costs.
+DACP (Xie et al., 2023) introduces Efficient Task-Specific (ETS-DACP) and Task-Agnostic (ETA-DACP) strategies to optimize performance by focusing either on task-specific foundational models or by selecting the most informative domain samples.
+Given the high cost of training from scratch and limited annotated data, EcomGPT-CT (Ma et al., 2023) explores continual pre-training on unlabeled general and E-commerce corpora, employing a data-mixing strategy to better handle semi-structured data.
+Experimental results across multiple tasks show continual pre-training and data-mixing strategies improve few-shot and zero-shot performance.
+Parameter-Efficient Tuning Methods.
+Lifelong-MoE (Chen et al., 2023) (Figure 5(b)) uses a Mixture-of-Experts (MoE) architecture, expanding capacity by adding new experts while freezing prior experts and gating mechanisms.
+It surpasses existing methods with better few-shot performance on 19 NLP tasks.
+Instruction Tuning-based Methods.
+Adapt-Retrieve-Revise (ARR)(Zhang et al., 2023c) aims at reducing hallucinations in specialized domains such as the Chinese legal domain.
 It consists of three steps: adapting a 7-billion-parameter language model for initial responses, retrieving corroborative evidence from an external knowledge base, and integrating these to refine the final response with GPT-4.
-Gogoulou et al.
-(Gogoulou et al., 2023) study the pros and cons of updating a language model when new data comes from new languages – the case of continual learning under language shift.
-They feed various languages into the model to examine the impact of pre-training sequence and linguistic characteristics on both forward and backward transfer effects across three distinct model sizes.
-A new continual learning (CL) problem, named Continual Knowledge Learning (CKL), is introduced by Jang et al.
-(Jang et al., 2022b).
-To assess CKL approaches, the authors establish a benchmark and metric measuring knowledge retention, updating, and acquisition.
-Continual Pre-training Methods
-LLMs have demonstrated remarkable proficiency in tackling open-domain tasks.
-However, their application in specific domains faces notable challenges, encompassing the lack of domain-specific knowledge, limited capacity to utilize such knowledge, and inadequate adaptation to domain-specific data formats.
-To address these issues, researchers have explored a novel approach known as continual pre-training, aiming to adapt LLMs to specific domains (Xie et al., 2023; Ma et al., 2023; Cheng et al., 2023a).
-Among these studies, Cheng et al.
-(Cheng et al., 2023a) draw inspiration from human learning patterns to develop a novel method that transforms raw corpora into reading comprehension texts.
-Furthermore, they discover that while training directly on raw data enhances the model’s domain knowledge, it significantly hurts the question-answering capability of the model.
-Domain-adaptive Continual Pre-training (DACP) uses a large domain corpus, leading to high costs.
-To reduce these, Xie et al.
-(Xie et al., 2023) propose two strategies: Efficient Task-Similar Domain-Adaptive Continual Pre-training (ETS-DACP) and Efficient Task-Agnostic Domain-Adaptive Continual Pre-training (ETA-DACP).
-ETS-DACP is tailored to improve performance on specific tasks, building task-specific foundational LLMs.
-Conversely, ETA-DACP selects the most informative samples across the domain.
-Given the high cost of training LLMs from scratch and limited annotated data in certain domains, Ma et al.
-(Ma et al., 2023) propose a novel model called EcomGPT-CT.
-It employs a fusion strategy to exploit semi-structured E-commerce data.
-Moreover, multiple tasks are designed to assess LLMs’ few-shot in-context learning ability and zero-shot performance after fine-tuning.
-Parameter-Efficient Tuning Methods.
-Chen et al.
-(Chen et al., 2023) present an innovative Lifelong Learning framework, termed Lifelong-MoE, which leverages a Mixture-of-Experts (MoE) architecture (Figure 5(b)).
-This architecture enhances the model’s capacity by incorporating new experts, where previously trained experts and gating mechanisms are frozen.
+Furthermore, LLM-CL (Ding et al., 2024), designed for aspect-based sentiment analysis, employs task-specific prompts and independently learns both domain-shared and domain-specific knowledge.
 5.1.3. VLMs-based DIL
-Vision-language models (VLMs) have demonstrated superiority in domain-incremental learning contexts.
-Yi et al.
-(Yi et al., 2023) integrate VLMS with continual learning methodologies to develop a general-purpose medical AI.
-Moreover, their study highlights the significance of data-efficient adaptation algorithms in minimizing the necessity for extensive labeling when transitioning to new domains or tasks.
-Furthermore, the prompt text is utilized to master the pre-trained knowledge embedded within VLMs.
-Aiming to independently learn prompts across disparate domains by using pre-trained VLMs, S-Prompt (Wang et al., 2022b) is devised (Figure 5(c)).
-This method encompasses techniques for acquiring image prompts and introduces an innovative methodology for language-image prompt acquisition.
-Prompt learning is conducted separately, utilizing a unified cross-entropy loss function during training.
-During inference, a K-NN (k-nearest neighbors) technique is employed to discern the domain.
-In the domain of visual question answering, Zhang et al.
-(Zhang et al., 2023d) introduce VQACL, a novel framework designed to effectively integrate data from both visual and linguistic modalities.
-This integration is achieved through a dual-level task sequence that enhances the model’s performance on complex multimodal tasks.
-Central to VQACL is a compositionality test that evaluates the model’s ability to generalize new skill and concept combinations.
-The framework also incorporates a novel representation learning strategy that differentiates between sample-specific (SS) and sample-invariant (SI) features.
-SS features capture distinctive attributes of individual inputs, enhancing output uniqueness, while SI features, derived from category prototypes, ensure essential characteristics are retained.
+Vision-language models (VLMs) excel in domain-incremental learning, particularly through methods like S-Prompt (Wang et al., 2022b) (Figure 5(c)), which independently learns prompts across diverse domains using pre-trained VLMs.
+This approach introduces innovative image and language-image prompt acquisition techniques, utilizing a unified cross-entropy loss during training and a k-nearest neighbors (K-NN) method to identify domains during inference.
+In visual question answering, VQACL (Zhang et al., 2023d) enhances multimodal tasks through a dual-level task sequence.
+This framework uses a compositionality test to assess generalization to new skill-concept combinations and employs a representation learning strategy that differentiates sample-specific (SS) features, capturing unique input attributes, from sample-invariant (SI) features, which retain essential characteristics.
+DIKI (Tang et al., 2024) preserves pre-trained knowledge by using a residual attention mechanism to inject new information into a frozen backbone, minimizing interference.
+A distribution-aware calibration scheme ensures smooth integration of new knowledge while maintaining the VLM’s zero-shot capabilities.
+SC-MLLM (Liu et al., 2024) is a robotic manipulation framework that utilizes an exponential moving average (EMA) for continuous learning, reducing the risk of catastrophic forgetting.
+5.2. Task-Incremental Learning
+5.2.1. PLMs-based TIL
+Traditional Methods.
+Drawing inspiration from neurobiological mechanisms, HMI (Maekawa et al., 2023) (Figure 6(a)) integrates compressed representations of prior training instances, providing selective guidance for generating training samples.
+Unlike traditional Continual Relation Learning (CRL) that rely on large labeled datasets.
+Continual Few-Shot Relation Learning (CFRL) focuses on learning new relational patterns with minimal labeled data.
+ERDA (Qin and Joty, 2022a) tackles this challenge by selecting informative samples from unlabeled data and enforcing relational constraints in the embedding space.
+As for Lifelong Relation Extraction, EMR (Wang et al., 2019b) utilizes a working memory mechanism to selectively replay stored samples, helping the model retain prior knowledge while learning new tasks.
+Learning user representations is fundamental for personalized recommender systems, but current approaches often train separate models for each task, which leads to inefficiency.
+To improve this, Conure (Yuan et al., 2021) introduces a framework that manages multiple tasks by pruning less critical parameters to accommodate new, task-specific ones.
+This method allows for positive transfer learning and preserves essential parameters to avoid knowledge loss.
+However, learning task-specific user representations for each task is impractical.
+Recent studies focus on universal user representations, which capture generalized user traits relevant to multiple tasks.
+TERACON (Kim et al., 2023) addresses this challenge by employing task-specific soft masks to isolate parameters and ensure effective knowledge retention.
 (a) HMI
@@
 Figure 6. Frameworks in TIL: HMI (PLM-based) (Maekawa et al., 2023), DynaMind (LLM-based) (Du et al., 2023), TRIPLET (VLM-based) (Qian et al., 2023).
-4.2. Task-Incremental Learning
-4.2.1. PLMs-based TIL
-Traditional Methods.
-Drawing inspiration from neurobiological mechanisms, Maekawa et al.
-(Maekawa et al., 2023) present an inventive approach known as Hippocampal Memory Indexing (HMI) to augment the generative replay technique.
-HMI leverages hippocampal memory indexing to integrate compressed representations of prior training instances, facilitating selective guidance for the generation of training samples.
-This methodological refinement contributes to heightened specificity, balance, and overall quality of the replayed samples.
-In tackling the Continual Few-Shot Relation Learning (CFRL) challenge, Qin et al.
-(Qin and Joty, 2022a) propose ERDA as a solution, drawing upon replay- and regularization-based techniques.
-The ERDA framework integrates embedding space regularization and data augmentation strategies to effectively confront the task of acquiring new relational patterns from scarce labeled instances while mitigating the risk of catastrophic forgetting.
-Wang et al.
-(Wang et al., 2019b) introduced a memory-based approach to continual learning, termed Episodic Memory Replay (EMR).
-This method leverages working memory by selectively replaying stored samples during each iteration of learning new tasks, thereby facilitating the integration of new knowledge while preserving previously acquired information.
-Conure (Yuan et al., 2021) is a framework that effectively manages multiple tasks by leveraging the redundancy of parameters in deep user representation models.
-Initially, it prunes less critical parameters to make room for new, task-specific ones.
-Subsequently, it incorporates these new parameters while retaining key parameters from prior tasks, facilitating positive transfer learning.
-To prevent the loss of previously acquired knowledge, it maintains these essential parameters in a fixed state.
-Notably, TERACON (Kim et al., 2023) utilizes task-specific soft masks to isolate parameters, which not only targets parameter updates during training but also clarifies the relationships between tasks.
-This method includes a novel knowledge retention module that utilizes pseudo-labeling to mitigate the well-known problem of catastrophic forgetting.
-Ke et al.
-(Ke et al., 2021a) introduced a model known as CTR, which employs innovative techniques such as CL-plugin and task masking to tackle the issue of catastrophic forgetting and to facilitate knowledge transfer across tasks.
-These strategies are particularly effective when utilized in conjunction with pre-trained language models, such as BERT, enhancing their adaptability and efficacy.
-In the specific context of User Intent Classification (UIC) within large-scale industrial applications, Wang et al.
-(Wang et al., 2021a) introduce a novel methodology, MeLL, which utilizes a BERT-based text encoder to generate robust, dynamically updated text representations for continual learning.
-MeLL combines global and local memory networks to preserve prototype representations across tasks, acting as a meta-learner that rapidly adapts to new challenges.
-It employs a Least Recently Used (LRU) policy for efficient global memory management and minimizes parameter growth.
-Recent advancements in conversational AI have concentrated on mitigating the limitations of traditional chatbots, which are dependent on static knowledge bases and extensive manual data annotation.
-The Lifelong INteractive learning in Conversation (LINC) methodology represents a significant stride in this direction, embodying a dynamic learning framework that mimics human cognitive processes during interactions (Liu and Mazumder, 2021; Mazumder and Liu, 2024; Liu, 2020).
-Structured around an Interaction Module, a Task Learner, and a Knowledge Store, LINC enables chatbots to dynamically integrate and utilize knowledge within conversational contexts (Figure LABEL:fig:LINC).
-This approach not only facilitates real-time information extraction and learning from user interactions but also addresses challenges such as managing erroneous inputs, refining conversational skills, and maintaining user engagement, thereby enhancing the chatbots’ linguistic and interactive capabilities.
+Given that PLMs often suffers catastrophic forgetting when applied to CL scenarios, CTR (Ke et al., 2021a) tackles this by utilizing innovative techniques such as CL-plugins and task masking, which significantly improve knowledge transfer while reducing forgetting.
+This approach proves particularly effective on challenging benchmarks, such as the 20News dataset, where task dissimilarity often causes severe forgetting in other models.
+RecAdam (Chen et al., 2020) extends the concepts of EWC by applying a regularization term to the pre-trained weights, utilizing an annealing coefficient to progressively incorporate the importance of prior knowledge.
+The diversity of text distributions and intents across domains poses challenges for scaling user intent detection (UIC) in industrial applications.
+MeLL (Wang et al., 2021) proposes a solution by leveraging global and local memory networks to maintain cross-task representations, which allows the model to continually adapt to new tasks without significant loss of performance.
+MeLL also employs an LRU policy for efficient memory management and controls parameter growth, making it scalable for real-world applications.
+Recent advancements in conversational AI focus on overcoming the limitations of traditional chatbots, which rely on static knowledge bases and extensive manual data annotation.
+Lifelong INteractive learning in Conversation (LINC) introduces a dynamic learning framework that mirrors human cognitive processes, enabling chatbots to integrate and utilize knowledge in real-time (Liu and Mazumder, 2021; Mazumder and Liu, 2024; Liu, 2020).
+Structured with an Interaction Module, Task Learner, and Knowledge Store, LINC enhances chatbots by improving real-time information extraction, handling erroneous inputs, and refining conversational skills, thereby boosting linguistic and interactive capabilities.
 Continual Pre-training Methods.
-Continual pre-training represents a paradigm wherein pre-trained language models (PLMs) are progressively enhanced through the assimilation of new knowledge from expanding datasets.
-Sun et al.
-(Sun et al., 2020b) introduced a framework called ERNIE 2.0, which incrementally constructs pre-training tasks, enabling the acquisition of complex lexical, syntactic, and semantic nuances embedded within the training corpora.
-This model deviates from traditional fixed-task training, instead employing continual multi-task learning to refine its capabilities continually.
-Further advancing the field of continual pre-training, RecyclableTuning (Qin et al., 2023) introduces a novel concept of recyclable tuning that features two distinct strategies: initialization-based and distillation-based methods.
-The former uses fine-tuned weights of existing PLMs as the basis for further enhancements, capitalizing on the established parametric relationships.
-Conversely, the distillation-based approach harnesses outdated weights to maintain knowledge continuity and efficiency in successor models.
+Continual pre-training represents a paradigm where PLMs are progressively enhanced by assimilating new knowledge from expanding datasets.
+ERNIE 2.0 (Sun et al., 2020b) illustrates this approach by incrementally constructing pre-training tasks, enabling the model to capture increasingly complex lexical, syntactic, and semantic nuances.
+Unlike traditional fixed-task training, it employs a continual multi-task learning framework to enhance its capabilities.
+Advancing this field further, RecyclableTuning (Qin et al., 2023) introduces two strategies: initialization-based and distillation-based.
+The former builds on fine-tuned weights from existing PLMs, while the latter reuses outdated weights to preserve knowledge and improve efficiency in future models.
 Parameter-Efficient Tuning Methods.
-Expounding upon the crucial need for more efficacious knowledge integration, Zhang et al.
-(Zhang et al., 2022b) propose a pioneering strategy that entails the integration of Adaptive Compositional Modules alongside a replay mechanism.
-These modules are designed to dynamically adjust to new tasks and are supplemented by pseudo-experience replay, significantly enhancing knowledge transfer.
-This framework is distinguished by its adaptive integration of modules within transformer architectures, skillfully orchestrating the interactions between existing and new modules to address emerging tasks.
-Additionally, the implementation of pseudo-experience replay promotes efficient knowledge transfer across these modules.
-Concurrently, Jin et al.
-(Jin et al., 2021a) introduce the Continual Learning of Few-Shot Learners (CLIF) challenge, wherein a model accumulates knowledge continuously across a series of NLP tasks.
-Their investigation delves into the impact of continual learning algorithms on generalization capabilities and advances a novel approach for generating regularized adapters.
+Expounding upon the crucial need for more efficacious knowledge integration, ACM (Zhang et al., 2022) dynamically adjusts transformer architectures and leverages pseudo-experience replay to enhance knowledge transfer.
+Continual Learning of Few-Shot Learners (CLIF) addresses a gap between rapid generalization methods (such as few-shot learning) and traditional continual learning approaches, which are not inherently designed for such generalization.
+BiHNet-Reg (Jin et al., 2021) provides a solution by generating task-specific adapter weights through bi-level task representations and incorporating regularization.
 Instruction Tuning-based Methods.
-Zhao et al.
-(Zhao et al., 2022) propose the Prompt Conditioned VAE for Lifelong Learning (PCLL) specifically designed for task-oriented dialogue (ToD) systems.
-PCLL employs a conditional variational autoencoder influenced by natural language prompts to generate high-quality pseudo samples, effectively capturing task-specific distributions.
-Additionally, a distillation process is integrated to refine past knowledge by reducing noise within pseudo samples.
-Razdaibiedina et al.
-(Razdaibiedina et al., 2023) introduce Progressive Prompts (PP), a novel approach to continual learning in language models.
-PP addresses catastrophic forgetting without resorting to data replay or an excessive proliferation of task-specific parameters.
-This method involves acquiring a fresh soft prompt for each task, gradually appending it to previously learned prompts while keeping the base model unchanged.
-The ConTinTin paradigm (Yin et al., 2022) develops a computational framework for sequentially mastering a series of new tasks guided by explicit textual instructions.
-It synthesizes projected outcomes for new tasks based on instructions, while facilitating knowledge transfer from prior tasks (forward-transfer) and maintaining proficiency in previous tasks (backward-transfer).
-In the realm of lifelong in-context instruction learning aimed at enhancing the target PLM’s instance- and task-level generalization performance as it observes more tasks, DYNAINST is devised by Mok et al.
-(Mok et al., 2023).
-DYNAINST integrates the principles of parameter regularization and experience replay.
-The regularization technique employed by DYNAINST is tailored to foster broad local minima within the target PLM.
-In order to devise a memory- and computation-efficient experience replay mechanism, they introduce Dynamic Instruction Replay, comprising Dynamic Instance Selection (DIS) and Dynamic Task Selection (DTS).
-DIS and DTS dynamically determine the selection of instances and tasks to be stored and replayed, respectively.
+PCLL (Zhao et al., 2022), designed for task-oriented dialogue systems, uses a conditional variational autoencoder with prompts to generate pseudo samples that capture task-specific distributions.
+It incorporates distillation to reduce noise in these samples, outperforming baselines like Experience Replay and HAT in intent detection and slot filling.
+PP (Razdaibiedina et al., 2023) mitigates catastrophic forgetting and enables forward transfer without data replay or excessive task-specific parameters.
+This method appends new soft prompts to previously learned ones, keeping the base model unchanged.
+It surpasses LPT5, IDBR, and MBPA++ on both standard and custom continual learning benchmarks.
+Furthermore, ConTinTin (Yin et al., 2022) provides a framework for sequentially mastering tasks through textual instructions, supporting both forward and backward transfer of knowledge.
+DYNAINST (Mok et al., 2023), designed for lifelong in-context instruction learning, enhances a PLM’s generalization by combining parameter regularization with experience replay.
+Its Dynamic Instruction Replay, using Dynamic Instance Selection (DIS) and Dynamic Task Selection (DTS), optimizes memory and computation by selectively replaying relevant instances and tasks.
 5.2.2. LLMs-based TIL
-Recent attention has focused on the convergence of large language models (LLMs) with continual learning methodologies, exemplified by significant contributions such as those by Wang et al.
-(Wang et al., 2024a) and Peng et al.
-(PENG et al., 2024).
+Recent attention has focused on the convergence of LLMs with CL methods, exemplified by significant contributions (Wang et al., 2024b; PENG et al., 2024).
 Benefiting from vast corpora and advanced hardware infrastructure, LLMs showcase remarkable capabilities in language comprehension and generation.
@@
 Traditional Methods.
-DynaMind, introduced by Du et al.
-(Du et al., 2023), emerges as a pioneering framework that intricately incorporates memory mechanisms and modular operators, enhancing the precision of LLM outputs.
-Comprising three essential components, DynaMind includes a memory module dedicated to storing and updating acquired knowledge, a modular operator for processing input data, and a continual learning module responsible for dynamically adjusting LLM parameters in response to new knowledge.
-Furthermore, Luo et al.
-(Luo et al., 2023) conduct a thorough investigation into catastrophic forgetting (CF) in Large Language Models (LLMs) during continual fine-tuning.
-Their experiments across various domains, including domain knowledge, reasoning, and reading comprehension, reveal the prevalence of CF in LLMs ranging from 1b to 7b scale, with severity increasing with model size.
-Comparative analysis between decoder-only BLOOMZ and encoder-decoder mT0 indicates that BLOOMZ exhibits less forgetting.
-Additionally, LLMs demonstrate the ability to mitigate language bias during continual fine-tuning.
-Contrasting ALPACA against LLAMA, the study highlights ALPACA’s superiority in preserving knowledge and capacity, suggesting that general instruction tuning aids in mitigating CF during subsequent fine-tuning phases.
-This research provides valuable insights into CF dynamics in LLMs, offering strategies for knowledge retention and bias mitigation.
-Wang et al.
-(Wang et al., 2024a) present TRACE, an innovative benchmark meticulously designed to evaluate continual learning capabilities in LLMs.
-Comprising eight distinct datasets spanning challenging tasks, including domain-specific challenges, multilingual capabilities, code generation, and mathematical reasoning, TRACE serves as a comprehensive evaluation platform.
-The authors rigorously examine the effectiveness of conventional Continual Learning (CL) methods when applied to LLMs within the TRACE framework.
-Peng et al.
-(PENG et al., 2024) propose the Joint Adaptive ReParameterization (JARe) framework, enhanced with Dynamic Task-related Knowledge Retrieval (DTKR), to facilitate adaptive adjustment of language models tailored to specific downstream tasks.
-This innovative approach leverages task distribution within the vector space, aiming to streamline and optimize the continual learning process seamlessly.
+DynaMind (Du et al., 2023) (Figure 6(b)) stands as a pioneering framework that enhances LLM output precision through the integration of memory mechanisms and modular operators.
+The framework consists of three key components: a memory module that stores and updates learned knowledge, a modular operator that processes incoming data, and a continual learning (CL) module that dynamically adjusts LLM parameters as new knowledge is introduced.
+In parallel, JARe (PENG et al., 2024) incorporates Dynamic Task-related Knowledge Retrieval (DTKR) to enable adaptive adjustments of language models for specific downstream tasks.
+By leveraging task distributions within vector space, JARe aims to optimize the continual learning process, streamlining task-specific knowledge adaptation.
 Parameter-Efficient Tuning Methods.
-Large language models (LLMs) encounter several substantial challenges that limit their practical applications.
-These include high computational requirements, significant memory demands, and a tendency toward catastrophic forgetting.
-Such limitations highlight the need for ongoing research into more efficient and robust approaches to training and deploying these models.
-Continual Parameter-Efficient Tuning (ConPET) (Song et al., 2023) is designed for the continuous adaptation of LLMs across diverse tasks, leveraging parameter-efficient tuning (PET) strategies to enhance both efficiency and performance.
-ConPET encompasses two primary modes: Static ConPET and Dynamic ConPET.
-Static ConPET adapts techniques previously utilized in smaller models for LLMs, thus minimizing tuning costs and reducing the risk of overfitting.
-Conversely, Dynamic ConPET enhances scalability by employing distinct PET modules for varying tasks, supplemented by a sophisticated selector mechanism.
-Moreover, the ELM strategy (Jang et al., 2023) involves initially training a compact expert adapter on the LLM for each specific task, followed by deploying a retrieval method to select the most appropriate expert LLM for each new task.
-Furthermore, Wang et al.
-(Wang et al., 2023a) have proposed orthogonal low-rank adaptation (O-LoRA), a straightforward yet efficacious method for facilitating continual learning in language models.
-O-LoRA mitigates catastrophic forgetting during task acquisition by employing distinct low-rank vector subspaces maintained orthogonally to minimize task interference.
-This method highlights the potential of orthogonal subspace techniques in improving the adaptability of language models to new tasks without compromising previously acquired knowledge.
+Large language models (LLMs) face significant challenges that limit their practical applications, such as high computational and memory demands, and vulnerability to catastrophic forgetting.
+These issues highlight the need for more efficient and robust training and deployment strategies.
+ConPET (Song et al., 2023), designed for the continual adaptation of LLMs across diverse tasks, leverages parameter-efficient tuning (PET) strategies to enhance both efficiency and performance.
+ConPET operates in two modes: Static ConPET, optimized for moderate-scale tasks, and Dynamic ConPET, tailored to large-scale tasks.
+Additionally, ELM (Jang et al., 2023) introduces a compact expert adapter trained for each task, with a retrieval mechanism that selects the most suitable expert LLM for new tasks.
+Furthermore, O-LoRA (Wang et al., 2023a) preserves task-specific knowledge by using low-rank approximations of previous task gradients, while the orthogonality constraint ensures task-specific subspaces remain distinct, minimizing interference from new tasks.
+This method demonstrates robustness across various task sequences.
 Instruction Tuning-based Methods.
-Scialom et al.
-(Scialom et al., 2022a) introduced Continual-T0, an innovative framework aimed at exploring the capabilities of large language models (LLMs) through continual learning, incorporating rehearsal techniques.
-A central aspect of this approach is the employment of instruction tuning, a key strategy designed to enhance the adaptability and effectiveness of LLMs when encountering novel tasks.
-Leveraging self-supervised pre-training, Continual-T0 demonstrates exceptional proficiency in mastering new language generation tasks while maintaining high performance across a diverse range of 70 previously encountered datasets.
-Despite the demonstrated proficiency of LLMs in adhering to instructions, their ability to generalize across underrepresented languages remains suboptimal.
-In response, InstructAlign (Cahyawijaya et al., 2023) is proposed to address this challenge by aligning newly introduced languages with those previously learned, which possess abundant linguistic resources, thereby mitigating instances of catastrophic forgetting.
-The core novelty of this approach lies in its advancement of language adaptation methodologies for instruction-tuned LLMs, with particular emphasis on integrating underrepresented languages.
+Continual-T0 (Scialom et al., 2022b) incorporates rehearsal techniques alongside instruction tuning of LLMs, showing strong performance across 70 datasets.
+However, its generalization to underrepresented languages remains limited.
+In response, InstructAlign (Cahyawijaya et al., 2023) proposes aligning new languages with those previously learned, leveraging abundant linguistic resources to mitigate catastrophic forgetting.
+The key innovation lies in enhancing language adaptation for instruction-tuned LLMs, particularly by incorporating underrepresented languages.
+Unlike traditional methods focused on single-task learning, RoboCoder (Li et al., 2024) enables robots to tackle a wide range of increasingly complex tasks by continuously refining and updating action codes based on environmental feedback.
+EUREKA (Ma et al., 2024) operates by leveraging LLMs to generate reward functions from the environment’s source code without requiring pre-defined templates or task-specific prompts.
+Through its evolutionary approach and in-context learning capabilities, it continually refines the rewards to achieve superior performance across diverse RL tasks.
 5.2.3. VLMs-based TIL
@@
 CTP (Zhu et al., 2023) employs topology preservation and momentum contrast to maintain consistent relationships within sample mini-batches across tasks, thereby preserving the distribution of prior embeddings.
-CTP also introduces the P9D dataset, comprising over one million image-text pairs across nine domains, aimed at visual language continuous pre-training (VLCP).
-Zheng et al.
-(Zheng et al., 2023) address the issue of zero-shot transfer degradation in visual language models by introducing the Zero-Shot Continual Learning (ZSCL) method.
-This novel approach utilizes a label-free dataset to facilitate distillation in the feature space, coupled with the application of weight regularization within the parameter space.
-Furthermore, they introduce a new benchmark, the Multi-Domain Task Incremental Learning (MTIL), designed to evaluate incremental learning strategies across various domains, thereby enhancing the assessment of such methods.
-Moreover, ZSCL has also been adapted for use in the CIL setting, further broadening its applicability
-Instruction Tuning-based Methods.
-By decoupling prompts and prompt interaction strategies, TRIPLET (Qian et al., 2023) effectively captures complex interactions between modalities.
-This includes specific designs for visual, textual, and fused prompts, as well as how to interact between different tasks through these prompts and retain crucial information, thereby reducing catastrophic forgetting.
-Decoupled prompts are designed to separate prompts in terms of multi-modality, layer-wise, and complementary, with each type of prompt containing learnable parameters intended to capture modality-specific knowledge from pre-trained visual-language models and training data.
-The prompt interaction strategies consist of three main components: the Query-and-Match Strategy, the Modality-Interaction Strategy, and the Task-Interaction Strategy.
-These components work together to enhance the model’s adaptability to different tasks and its memory for old tasks.
-Moreover, COIN (Chen et al., 2024) introduces a new continuous instruction tuning benchmark designed to evaluate the performance of Multimodal Large Language Models (MLLMs) in the sequential instruction fine-tuning paradigm.
-CoIN includes 10 commonly used data sets covering 8 task categories, ensuring the diversity of instructions and tasks.
-In addition, the trained model is evaluated from two aspects: instruction following and general knowledge, which respectively evaluate the consistency with human intention and the preservation of knowledge for reasoning.
-CoIN converts commonly used visual-linguistic datasets into instruction fine-tuning data formats by using different templates to explore the behavior of MLLMs in continuous instruction fine-tuning.
-This method takes into account the diversity between different tasks and attempts to enhance the model’s adaptability to new and old tasks through diversified instruction templates.
-To alleviate the catastrophic forgetting problem of MLLMs, CoIN introduces the MoELoRA method.
-This method reduces forgetting by using different experts to learn knowledge on different tasks and using gate functions to regulate the output of these experts.
+ZSCL (Zheng et al., 2023) addresses the challenge of zero-shot transfer degradation in VLMs by employing a label-free dataset for distillation in the feature space, alongside weight regularization in the parameter space.
+Additionally, ZSCL has been adapted for the CIL setting, expanding its applicability to a wider range of continual learning scenarios.
+SND (Yu et al., 2024a) tackles catastrophic forgetting and zero-shot performance degradation in VLMs by leveraging a dual-teacher model.
+The pre-trained VLM retains zero-shot generalization, while the fine-tuned VLM preserves task-specific knowledge.
+A teacher selection mechanism quantifies discrepancies between their feature representations.
+Similarly, AwoForget (Zheng et al., [n.
+d.]) employs a graph-based multi-modal proximity distillation framework to mitigate forgetting in VLMs.
+It constructs a cross-modal graph to represent relationships between image-text pairs and adopts a dual-teacher distillation approach for maintaining both zero-shot and task-specific knowledge.
+This method also extends its applicability to CIL.
 Parameter-Efficient Methods.
-MoE-Adapters4CL (Yu et al., 2024) introduces a parameter-efficient continual learning method to mitigate long-term forgetting in incremental learning of visual-language models.
+MoE-Adapters4CL (Yu et al., 2024b) introduces a parameter-efficient continual learning method to mitigate long-term forgetting in incremental learning of VLMs.
 Their approach involves dynamically extending the pre-trained CLIP model to accommodate new tasks by integrating a Mixture-of-Experts (MoE) adapter.
@@
 Furthermore, the MoE-Adapters4CL framework has also been adapted for use in the CIL setting.
+Instruction Tuning-based Methods.
+By decoupling prompts and prompt interaction strategies, TRIPLET (Qian et al., 2023) (Figure 6(c)) effectively captures complex interactions between modalities.
+This includes specific designs for visual, textual, and fused prompts, as well as how to interact between different tasks through these prompts and retain crucial information, thereby reducing catastrophic forgetting.
+Decoupled prompts are designed to separate prompts in terms of multi-modality, layer-wise, and complementary, with each type of prompt containing learnable parameters intended to capture modality-specific knowledge from pre-trained VLMs and training data.
 5.3. Class-Incremental Learning
@@
 Traditional Methods.
-The study presented in (Monaikul et al., 2021) introduces ExtendNER, a novel framework for continual learning in Named Entity Recognition (NER) that obviates the need for extensive re-annotation.
-This framework employs a knowledge distillation (KD) technique, wherein a pre-existing named entity recognition (NER) model, termed the ”teacher”, imparts knowledge to a newly developed model, termed the ”student”.
-The student model, designed to identify new entity types, progressively learns by emulating the teacher model’s responses to a new dataset.
-This method allows the student to acquire the ability to recognize new entities while retaining knowledge of previously identified ones.
-Additionally, Liu et al.
-(Liu et al., 2019b) propose an innovative method for learning distributed representations of sentences.
-This method initiates with the configuration of sentence encoders using features independent of any specific corpus, followed by iterative updates through Boolean operations on conceptor matrices.
-This technique ensures that the encoders maintain their performance on existing datasets while adapting effectively to new data.
-(a) IDBR
+ExtendNER (Monaikul et al., 2021) (Figure 7(a)) introduces a continual learning framework for Named Entity Recognition (NER) that reduces re-annotation efforts.
+It uses knowledge distillation, where a ”teacher” NER model helps a ”student” model learn to recognize new entity types while retaining knowledge of previous ones.
+SRC (Liu et al., 2019b) focuses on sentence representation learning by initializing corpus-independent encoders and refining them through Boolean operations on conceptor matrices.
+IDBR (Huang et al., 2021) improves continual text classification by disentangling task-generic and task-specific representations.
+It uses regularization techniques and auxiliary tasks like next sentence prediction to enhance knowledge retention and generalization, outperforming many existing methods across various task sequences.
+(a) ExtendNER
 (b) PLE
 (c) Adaptation-CLIP
-Figure 7. Frameworks in CIL: IDBR (PLM-based) (Huang et al., 2021), PLE (PLM-based) (Li et al., 2022b), Adaptation-CLIP (VLM-based) (Liu et al., 2023a).
-Huang et al.
-(Huang et al., 2021) introduced an innovative methodology known as Information Disentanglement-based Regularization (IDBR) to address the enduring challenges associated with continual text classification.
-This method effectively disentangles the hidden spaces of text into task-generic and task-specific representations, employing distinct regularization strategies to enhance knowledge retention and facilitate generalization.
-Furthermore, the integration of two auxiliary tasks, namely next sentence prediction and task-id prediction, serves to augment the learning process by reinforcing the separation between generic and specific representational spaces.
+Figure 7. Frameworks in CIL: ExtendNER (PLM-based) (Monaikul et al., 2021), PLE (PLM-based) (Li et al., 2022b), Adaptation-CLIP (VLM-based) (Liu et al., 2023a).
 Instruction Tuning-based Methods.
-Introduced by Varshney et al.
-(Varshney et al., 2022), Prompt Augmented Generative Replay (PAGeR) is a method that enables continual learning in intent detection without retaining past data.
-PAGeR leverages pre-trained language models to generate intent-specific utterances specific to new intents while preserving accuracy on existing ones.
-Unlike exemplar replay, which stores specific examples, PAGeR is structured to selectively maintain relevant contexts for each specified intent, which are then employed as generation prompts.
-This approach combines utterance generation and classification into one model, enhancing knowledge transfer and optimization.
+Instead of retaining examples, PAGeR (Varshney et al., 2022) uses PLMs to generate intent-specific utterances for new tasks while maintaining performance on previous ones.
+It selectively preserves relevant contexts as prompts, ensuring efficient memory usage.
+By utilizing PLMs, PAGeR is effective in real-world scenarios where new intents and relations are continuously introduced, offering a memory-efficient alternative to traditional methods.
 Parameter-Efficient Tuning Methods.
-In the domain of task-oriented dialogue systems, Continual Few-shot Intent Detection (CFID) focuses on recognizing new intents with few examples.
-Li et al.
-(Li et al., 2022b) propose a Prefix-guided Lightweight Encoder (PLE) to address this by using a parameter-efficient tuning method that combines a Continual Adapter module with a frozen Pre-trained Language Model (PLM) and a Prefix-guided Attention mechanism to reduce forgetting.
-To further mitigate forgetting, the Pseudo Samples Replay (PSR) strategy reinforces prior knowledge by replaying crucial samples from previous tasks.
-The Teacher Knowledge Transfer (TKT) strategy uses distillation to transfer task-specific knowledge to maintain performance on new tasks.
-Additionally, the Dynamic Weighting Replay (DWR) strategy dynamically adjusts weights of previous tasks to balance new knowledge acquisition with the revision of old tasks, navigating the variability and potential negative impacts of prior tasks
-The Efficient Parameter Isolation (EPI) method, introduced in Wang et al.
-(Wang et al., 2023c), assigns unique subsets of private parameters to each task alongside a shared pre-trained model.
-This approach ensures precise parameter retrieval and has been shown to outperform non-caching methods in continual language learning benchmarks, while remaining competitive with caching methods.
-Furthermore, EPI employs random static masking to reduce storage requirements, increasing its viability in resource-constrained environments.
-In complex system environments, efficient and adaptable machine learning architectures are crucial, especially for classification tasks with sequentially presented data.
-Wójcik et al.
-(Wójcik et al., 2023) devise a novel architecture, Domain and Expertise Ensemble (DE&E), comprising a feature extractor, classifier, and gating mechanism.
-The feature extractor employs a multi-layer neural network to convert input data into embeddings, while the classifier, a mixture of binary class-specific experts, leverages a gating mechanism to select the appropriate expert for the current input dynamically.
-This Mixture of Experts-based method promotes incremental learning by training experts with class-specific samples and combines their outputs during testing to derive the final classification.
+In task-oriented dialogue systems, continuous few-shot intent detection (CFID) focuses on recognizing new intents with a few examples.
+To address this, PLE (Li et al., 2022b) (Figure 7(b)) employs a parameter-efficient tuning approach that integrates a Continual Adapter module with a frozen PLM and a Prefix-guided Attention mechanism.
+This design effectively reduces forgetting while minimizing the number of trainable and stored parameters, making it an ideal choice for real-world applications where computational efficiency is critical.
+Similarly, EPI (Wang et al., 2023c) allocates task-specific private parameters alongside a shared model, ensuring precise retrieval and superior performance in continual learning.
+It reduces storage needs through random static masking and consistently outperforms rehearsal-free methods, even competing with rehearsal-based ones on far-domain datasets.
+For environments with sequentially presented data, DE&E (Wójcik et al., 2023) uses a mixture of binary class-specific experts, dynamically selecting the best expert for each input, facilitating incremental learning by combining expert outputs for final classification.
 5.3.2. VLMs-based CIL
 Traditional Methods.
-Kim et al.
-(Kim et al., 2024) introduce VLM-PL, a novel approach for class incremental object detection that incorporates new object classes into a detection model without forgetting old ones.
-Utilizing a visual-language model (VLM), this method enhances the pseudo-labeling process to improve the accuracy and performance of object detection in continual learning settings.
-VLM-PL starts with a pre-trained detector to generate initial pseudo-labels, which are then validated through a visual-language evaluation using a specially designed hint template.
-Accurate pseudo-labels are retained and combined with ground truth labels to train the model, ensuring it remains proficient in recognizing both new and previously learned categories.
-Cao et al.
-(Cao et al., 2024) introduce a framework for a Generative Multi-modal Model (GMM) that leverages large language models for class-incremental learning.
+VLM-PL (Kim et al., 2024) leverages a VLM to enhance the pseudo-labeling process, integrating new object classes into a detection model without compromising previously learned categories.
+GMM (Cao et al., 2024) leverages LLMs for class-incremental learning.
 This innovative approach entails the generation of labels for images by employing an adapted generative model.
-Following the production of detailed textual descriptions, a text encoder is utilized to extract salient features from these descriptions.
-These extracted features are subsequently aligned with existing labels to ascertain the most fitting label for classification predictions.
-PROOF (Zhou et al., 2023c) develops a method to enhance model memory retention when adapting to downstream tasks.
-This method involves a projection technique that maps pre-trained features into a new feature space designed to preserve prior knowledge.
-Additionally, to effectively utilize cross-modal information, PROOF introduces a fusion module that employs an attention mechanism.
-This module adjusts both visual and textual features simultaneously, enabling the capture of semantic information with enhanced expressive power.
-Recently, Jha et al.
-(Jha et al., 2024) present a new approach for adapting pre-trained vision-language models like CLIP to new tasks without forgetting previous knowledge.
-It employs a Variational Inference framework to probabilistically model the distribution of visual-guided text features, enhancing fine-tuning reliability by accounting for uncertainties in visual-textual interactions.
-Key to CLAP is the visual-guided attention (VGA) module, which aligns text and image features to prevent catastrophic forgetting.
-Additionally, CLAP includes lightweight, task-specific inference modules that learn unique stochastic factors for each task, allowing continuous adaptation and knowledge retention.
+PROOF (Zhou et al., 2023c) maps pre-trained features into a new feature space designed to preserve prior knowledge.
+To enhance the use of cross-modal information, it incorporates a fusion module with an attention mechanism, allowing the model to retain essential features when adapting to new tasks.
+CLAP (Jha et al., 2024) offers a new method for adapting VLMs to novel tasks without losing previously learned knowledge.
+This approach employs a Variational Inference framework to probabilistically model the distribution of visual-guided text features, enhancing fine-tuning reliability by accounting for uncertainties in visual-textual interactions.
 Parameter-Efficient Tuning Methods.
-Liu et al.
-(Liu et al., 2023a) introduce Adaptation-CLIP, which employs three strategies for CLIP’s continual learning: linear adapter, self-attention adapter, and prompt tuning.
-The first two strategies add a linear layer and a self-attention mechanism, respectively, after the image encoder while freezing the remaining architecture.
-The third, prompt tuning, integrates trained prompts into the text encoder to enhance task comprehension and splices these with prior prompts to maintain continuity.
-To prevent catastrophic forgetting, a parameter retention strategy preserves significantly altered parameters from $M_{t-1}$ to $M_{t}$, ensuring stability and effective continual learning.
+Adaptation-CLIP (Liu et al., 2023a) (Figure 7(c)) employs three strategies for continual learning: a linear adapter, a self-attention adapter, and prompt tuning.
+The adapters add layers to the image encoder while freezing the rest of the model, and prompt tuning enhances the text encoder by integrating new and prior prompts to maintain task comprehension.
+DIKI (Tang et al., 2024) prevents the disruption of pre-trained knowledge by using a residual attention mechanism that injects new knowledge into a frozen backbone, minimizing interference.
+A distribution-aware calibration scheme is employed to maintain the zero-shot capabilities of VLMs.
+RAPF (Huang et al., 2024) addresses class-incremental learning challenges using a pre-trained CLIP model with a linear adapter.
+It minimizes interference by identifying neighboring categories and applying hinge loss, while generating old class features from a Gaussian distribution to aid classification.
+STAR-Prompt (Menabue et al., 2024) introduces a two-level prompting mechanism to balance stability and plasticity in large models like CLIP and ViT, reducing catastrophic forgetting while enabling continuous learning.
 Instruction Tuning-based Methods.
-Khan et al.
-(Khan et al., 2023) introduce two notable advancements: an enhanced prompt pool key query mechanism and category-level language guidance.
-The key query mechanism uses CLS features to improve prompt selection, featuring key replacement with a fixed CLS tag and dynamic mapping to task-level language representations, thereby enhancing accuracy and robustness across different tasks.
-Meanwhile, category-level language guidance is implemented in the vision transformer to better align output features with category-specific language representations, significantly improving task handling and category differentiation, leading to improved model performance.
+LGCL (Khan et al., 2023) introduces two key innovations: a refined prompt pool key query mechanism and category-level language guidance.
+The former leverages CLS features to optimize prompt selection with dynamic task-level mappings, improving accuracy and robustness, while the latter aligns output features with specific language representations in the vision transformer, enhancing task handling and model performance.
 6. Online Continual Learning
@@
 The Hard Task Boundary (HTB) setting has been developed to enable continuous knowledge acquisition by learning models from a dynamically changing stream of textual data, without the need for dataset identifiers.
-For example, Shen et al.
-(Shen et al., 2019) have implemented HTB in slot filling, Michieli et al.
-(Michieli et al., 2023) have utilized it in audio classification, and Vander et al.
-(Vander Eeckt et al., 2023) have explored its use in automatic speech recognition.
+For example, ProgModel (Shen et al., 2019) have implemented HTB in slot filling, TAP-SLDA (Michieli et al., 2023) have utilized it in audio classification, and AOS (Vander Eeckt et al., 2023) have explored its use in automatic speech recognition.
 Traditional Methods.
@@
 These include parameter-isolation-based methods (de Masson D’Autume et al., 2019; Wang et al., 2020), replay-based methods (Holla et al., 2020) and regularization-based methods (Vander Eeckt et al., 2023; Liu et al., 2021b).
-MBPA++ (de Masson D’Autume et al., 2019) introduces a framework for lifelong language learning, enabling a pre-trained model to learn continually from textual examples without requiring labeled datasets.
+MBPA++ (de Masson D’Autume et al., 2019) (Figure 8(a)) introduces a framework for lifelong language learning, enabling a pre-trained model to learn continually from textual examples without requiring labeled datasets.
 It employs an episodic memory system with sparse experience replay and local adaptation techniques to prevent catastrophic forgetting.
 Extending this framework, Meta-MBPA++ (Wang et al., 2020) integrates three core lifelong learning principles, enhancing performance in text classification and question-answering tasks while using only 1% of the typical memory usage.
-Liu et al.
-(Liu et al., 2021b) introduce a regularization-based strategy, CID, for lifelong intent detection.
-This method uses cosine normalization, hierarchical knowledge distillation, and inter-class margin loss to tackle the challenges of data imbalances in the lifelong intent detection (LID) task, aiming to mitigate the negative impacts associated with these imbalances.
+CID (Liu et al., 2021b) is devised for lifelong intent detection.
+This method uses cosine normalization, hierarchical knowledge distillation, and inter-class margin loss to tackle the challenges of data imbalances in the lifelong intent detection task, aiming to mitigate the negative impacts associated with these imbalances.
 (a) MBPA++
-(b) CBA
-Figure 8. Frameworks in Online Continual Learning: MBPA++ (PLM-based HTB/BTB) (de Masson D’Autume et al., 2019), CBA (VLM-base BTB) (Wang et al., 2023d).
+(b) G-NoCL
+Figure 8. Frameworks in Online Continual Learning: MBPA++ (PLM-based HTB/BTB) (de Masson D’Autume et al., 2019), G-NoCL (VLM-base BTB) (Seo et al., 2024).
+6.1.2. VLMs-based HTB
+PEGP (Qiao et al., 2024) utilizes various parameter-efficient tuning methods and an orthogonal gradient projection mechanism to prevent catastrophic forgetting in continual learning.
+Using backbone networks like ViT and CLIP, PEGP ensures that gradient updates during training are orthogonal to the feature subspace of previous tasks, preserving old knowledge while learning new ones.
 6.2. Blurry Task Boundary
 6.2.1. PLMs-based BTB
-MBPA++ (de Masson D’Autume et al., 2019) and Meta-MBPA++ (Wang et al., 2020) can also exemplify models capable of adapting to environments with indistinct task boundaries.
-TPEM (Geng et al., 2021) employs a tripartite approach within an encoder-decoder framework, consisting of pruning, expanding, and masking techniques.
-Pruning helps preserve essential information from previous tasks, expansion increases the model’s capacity to accommodate new tasks, and masking reduces interference from previous tasks’ weights, thus enhancing learning efficiency.
-Originally, online meta-learning (OML) (Javed and White, 2019) and a neuromodulatory meta-learning algorithm (ANML) (Beaulieu et al., 2020) were intended to continuously learn new sequences of tasks during the testing phase.
+MBPA++ (de Masson D’Autume et al., 2019) and Meta-MBPA++ (Wang et al., 2020) exemplify models capable of adapting to environments with indistinct task boundaries.
+TPEM (Geng et al., 2021) adopts a tripartite approach within an encoder-decoder framework, utilizing pruning, expanding, and masking techniques.
+Pruning preserves essential information from previous tasks, expansion increases model capacity for new tasks, and masking mitigates interference from prior task weights, thereby enhancing learning efficiency.
+Online meta-learning (OML) (Javed and White, 2019) and the neuromodulatory meta-learning algorithm (ANML) (Beaulieu et al., 2020) are initially designed to learn sequences of tasks during the testing phase.
 Holla et al.
-(Holla et al., 2020) adapt these algorithms for a conventional continual learning context, where the focus is on evaluating performance on previously encountered tasks.
-The enhanced versions, named OML-ER and ANML-ER, incorporate an episodic memory module designed for experience replay.
+(Holla et al., 2020) adapts these algorithms for conventional continual learning, focusing on performance evaluation on previous tasks.
+Their enhanced versions, OML-ER and ANML-ER, incorporate an episodic memory module for experience replay.
+The S6 framework addresses the challenges of continual relation learning in real-world scenarios, where data arrives in a streaming manner with noisy labels and shifting distributions.
 6.2.2. VLMs-based BTB
-Traditional Methods.
-Cui et al.
-(Cui et al., 2024) propose the Dynamic Knowledge Rectification (DKR) framework, designed to mitigate the propagation of incorrect information in foundation LMs.
-The DKR framework operates by initially leveraging an existing model to identify and exclude obsolete or erroneous knowledge when confronted with new data.
-Subsequently, a rectification process is employed to amend these inaccuracies while ensuring the preservation of valid data associations.
-This process is especially vital when integrating new data, as it prevents the perpetuation of outdated or incorrect information.
-In cases where data is inaccessible through existing models, DKR utilizes paired ground-truth labels to support the continuous evolution of knowledge bases, thereby enhancing the model’s accuracy.
-Parameter-Efficient Tuning Methods.
-Wang et al.
-(Wang et al., 2023d) introduce the Continual Bias Adaptor (CBA) (Figure 8(b)), a novel method designed to enhance the efficacy of online CL by mitigating catastrophic forgetting.
-The CBA method dynamically modifies the classifier’s network to adapt to significant shifts in data distribution observed during training, thereby preserving knowledge from earlier tasks.
-Notably, the CBA module can be deactivated during testing phases, eliminating additional computational burdens or memory demands.
-This feature highlights the practicality and applicability of the CBA in real-world scenarios.
-Instruction Tuning-based Methods.
-In the BTB scenario, the Mask and Visual Prompt tuning (MVP) method, as detailed by Moon et al.
-(Moon et al., 2023), addresses challenges like intra- and inter-task forgetting and class imbalance effectively.
-MVP features instance-wise logit masking to prevent irrelevant information retention, contrastive visual prompt tuning loss to ensure consistent prompt selection, gradient similarity-based focal loss to focus on overlooked samples, and adaptive feature scaling to balance the integration of new knowledge with existing data retention.
-Online Learning for CV
-Recent advancements in class-incremental online learning predominantly address computer vision tasks, encapsulating various methodologies such as regularization-based (Yin et al., 2021; Gu et al., 2022; Guo et al., 2022; Bang et al., 2022), replay-based (Aljundi et al., 2019; Shim et al., 2021; Caccia et al., 2021; Lin et al., 2023; Zhang et al., 2022a), distillation-based (He et al., 2020; Fini et al., 2020; Koh et al., 2022), and gradient-based approaches (Jin et al., 2021b; Guo et al., 2023; Chen et al., 2020; Yoon et al., 2021; Ororbia et al., 2022).
-Among these, Koh et al.
-(Koh et al., 2021) introduced the Class-Incremental Blurry (CLIB) model, which distinguishes itself through its task-free nature, adaptability to class increments, and prompt response to inference queries, showing superior performance over traditional continual learning methods.
-In a related vein, Gunasekara et al.
-(Gunasekara et al., 2023) explore the Online Streaming Continual Learning (OSCL), a hybrid of Stream Learning (SL) and Online Continual Learning (OCL), which integrates aspects of both domains.
-Furthermore, issues of shortcut learning and bias in online continual learning have been tackled by Wei et al.
-(Wei et al., 2023) and Chrysalis et al.
-(Chrysakis and Moens, 2023).
-Additionally, Semola et al.
-(Semola et al., 2022) propose Continual-Learning-as-a-Service (CLaaS), a service model that leverages continual learning to monitor shifts in data distribution and update models efficiently.
-This array of developments highlights the dynamic capabilities of online continual learning frameworks to effectively address complex challenges across a variety of computer vision tasks.
-Table 1. The statistics information of the existing CL datasets. #D/T/C means the number of domains/tasks/classes, respectively.
-Datasets #Train #Val #Test #Total CL Settings NLP Problems Language #D/T/C
+DKR (Cui et al., 2024) is devised to mitigate the propagation of incorrect information in foundation LMs.
+It operates by initially leveraging an existing model to identify and exclude obsolete or erroneous knowledge when confronted with new data.
+Subsequently, a rectification process is employed to amend these inaccuracies while preserving valid data associations.
+SIT (Wang et al., 2024a) addresses the challenge of asymmetry in Online Lifelong Learning (OLL) with VLMs, where the image encoder is tuned on the current batch of data, while the text encoder can access features from previously seen classes.
+SIT reformulates the loss function by restricting the text features in the comparison to only those present in the current batch, ensuring that both image and text feature updates remain balanced.
+OLiVia-Nav (Narasimhan et al., 2024) is an innovative architecture designed for social navigation in human-centered environments like hospitals and offices, integrating VLMs with an online lifelong learning framework.
+The system employs a novel distillation approach, Social Context Contrastive Language Image Pre-training (SC-CLIP), to transfer the social reasoning capabilities of large VLMs into lightweight models, enabling real-time encoding of social and environmental context.
+Generative Name only Continual Learning (G-NoCL) (Seo et al., 2024) (Figure 8(b)) is a framework designed to overcome the challenges of manual annotation and web-scraped data in continual learning.
+The core of the model is the Prompt Refiner Module, which transforms basic concept names into diverse, context-rich prompts.
+Table 1. The statistics information of the existing CL datasets. #D/T/C means the number of domains/tasks/classes for DIL, TIL and CIL, respectively.
+Datasets #Train #Val #Test #Total CL Settings NLP Problems Source Language #D/T/C
 Offline
-Progressive Prompts (Razdaibiedina et al., 2023) - - - - TIL Sentiment analysis, topic classification, boolean QA, QA, paraphrase detection, word sense disambiguation, natural language inference English 15 tasks
-MeLL (Wang et al., 2021a) 1,430,880 173,781 118,240 1,722,901 TIL Intent classification English 1184 tasks
-Continual-T0 (Scialom et al., 2022b) 800,000 - 33,382 833,382 TIL Text Simplification, Headline Generation with Constraint, Haiku Generation, Covid QA, Inquisitive Question Generation, Empathetic Dialogue Generation, Explanation Generation, Twitter Stylometry English 8 tasks
-COPR (Zhang et al., 2023b) - - - - TIL QA tasks, summary task, positive file review generation task English 3 tasks
-Adaptive Compositional Modules (Zhang et al., 2022b) 50,725 - 27,944 78,669 TIL Natural language generation, SQL query generation, Summarization and Task-oriented dialogue English 8 tasks
-CODETASKCL (Yadav et al., 2023) 181,000 9,700 10,000 200,700 TIL Code generation, code translation, code summarization, and code refinement Hybrid 4 tasks
-Lifelong Simple Questions (Wang et al., 2019b) - - - - TIL single-relation questions English 20 tasks
-Lifelong FewRel (Wang et al., 2019b) - - - - TIL few-shot relation detection English 10 tasks
-InstrDialog (Zhang et al., 2023a) 9,500 950 1,900 12,350 TIL Dialogue state tracking, dialogue generation, intent identification English 19 tasks
-InstrDialog++ (Zhang et al., 2023a) 3,800 1,900 3,800 9,500 TIL Dialogue Generation, Intent Identification, Dialogue State Tracking, Style Transfer, Sentence Ordering, Word Semantics, Text Categorization, Pos Tagging, Fill in The Blank, Program Execution, Question Generation, Misc, Coherence Classification, Question Answering, Summarization, Commonsense Classification, Wrong Candidate Generation and Toxic Language Detection English 38 tasks
-ConTinTin (Yin et al., 2022) - - - - TIL Question generation tasks (QG), answer generation tasks (AG), classification tasks (CF), incorrect answer generation tasks (IAG), minimal modification tasks (MM) and verification tasks (VF) English 61 tasks
-Tencent TL (Yuan et al., 2021) - - - - TIL personalized recommendations and profile predictions English 6 tasks
-Movielens (Yuan et al., 2021) - - - - TIL personalized recommendations and profile predictions English 3 tasks
-NAVER Shopping (Kim et al., 2023) - - - - TIL search query prediction tasks English 6 tasks
-TRACE (Wang et al., 2024a) 40,000 - 16,000 56,000 TIL Domain-specific task, Multi-lingual task, Code completion task, Mathematical reasoning task Hybrid 8 tasks
-ABSC (Ke and Xu, 2021) 3,452 150 1,120 4,722 DIL Aspect-based sentiment classification English 19 Domains
-DecaNLP (Sun et al., 2020a) 169,824 - 32,116 201,940 DIL Question answering, semantic parsing, sentiment analysis, semantic role labeling, and goal-oriented dialogue English 5 domains
-Foundational text classification (Sun et al., 2020a) 115,000 - 7,600 122,600 DIL News classification, sentiment analysis, Wikipedia article classification, and question-and-answer categorization English 5 domains
-RVAE_LAMOL (Wang et al., 2022a) 15,870 - 5,668 21,538 DIL Oriented dialogue of the restaurant reservation task, semantic role labeling, sentiment classification English 3 domains
-COPR (Zhang et al., 2023b) - - - - DIL - English 18 domains
-SGD (Zhu et al., 2022) 38,745 5,210 11,349 40,287 DIL Dialogue state tracking English 19 Domains
-CPT (Ke et al., 2022) 3,121,926 - - 3,121,926 DIL Domain-adaptive pre-training task English 4 Domains
-CKL(Jang et al., 2022b) - - - 30,372 DIL Domain-adaptive pre-training task English 3 Domains
-ELLE (Qin et al., 2022) - - - - DIL Domain-adaptive pre-training task English 5 Domains
-Domain-incremental Paper Stream (Jin et al., 2022) - - - - DIL Relation extraction and named entity recognition English 4 domains
-Chronologically-ordered Tweet Stream (Jin et al., 2022) - - - - DIL multi-label hashtag prediction and single-label emoji prediction English 4 domains
-AdapterCL (Madotto et al., 2020) 31,426 4,043 4,818 40,287 DIL Intent classification, Dialogue State Tracking (DST), Natural Language Generation (NLG), end-to-end (E2E) modeling English 37 Domains
-DE&E (Wójcik et al., 2023) 28,982 - 12,089 41,071 CIL Text classification English 3 tasks
-EPI (Wang et al., 2023c) 12,840 3,524 6,917 23,281 CIL Text classification, topic classification English 13 Classes
-PAGeR (Varshney et al., 2022) 59,754 7,115 15,304 82,173 CIL Intent classification English 355 Classes
-PLE (Li et al., 2022b) 4,669 4,650 31,642 40,961 CIL Intent classification English 477 Classes
-CoNLL-03 (Monaikul et al., 2021) 23,326 5,902 5,613 34,841 CIL Named Entity Recognition English 4 Classes
-OntoNotes (Monaikul et al., 2021) 107,169 16,815 10,464 134,448 CIL Named Entity Recognition English 6 Classes
+Progressive Prompts (Razdaibiedina et al., 2023) - - - - TIL Mixed 15 classification tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers, MNLI, QQP, RTE, SST2, WiC, CB, COPA, MultiRC, BoolQ English 15 tasks
+MeLL (Wang et al., 2021) 1,430,880 173,781 118,240 1,722,901 TIL Intent classification Snips, TOP semantic parsing, Facebook Multilingual Task Oriented Dataset, e-commerce dialogue system AliMe English 1184 tasks
+Continual-T0 (Scialom et al., 2022a) 800,000 - 33,382 833,382 TIL Language generation task WikiAuto, Gigaword, Subreddit haiku, ELI5, Empathetic dialogues, e-SNLI, Twitter English 8 tasks
+COPF (Zhang et al., 2023b) - - - - TIL Value alignment IMDB, HH-RLHF, Reddit TL, DR English 3 tasks
+ACM (Zhang et al., 2022) 50,725 - 27,944 78,669 TIL Mixed 4 generation tasks E2E NLG, RNNLG, WikiSQL, CNN/DailyMail, MultiWOZ English 4 tasks
+CODETASKCL (Yadav et al., 2023) 181,000 9,700 10,000 200,700 TIL Mixed 4 code-centric tasks CONCODE, CodeTrans, CodeSearchNet, BFP Hybrid 4 tasks
+Lifelong SimpleQuestions (Wang et al., 2019b) - - - - TIL Relation extraction SimpleQuestions English 20 tasks
+Lifelong FewRel (Wang et al., 2019b) - - - - TIL Relation extraction FewRel English 10 tasks
+InstrDialog (Zhang et al., 2023a) 9,500 950 1,900 12,350 TIL Mixed 3 dialogue tasks SuperNI English 19 tasks
+InstrDialog++ (Zhang et al., 2023a) 3,800 1,900 3,800 9,500 TIL Mixed 22 dialogue tasks SuperNI English 38 tasks
+ConTinTin (Yin et al., 2022) - - - - TIL Mixed classification and generation tasks NATURAL-INSTRUCTIONS English 61 tasks
+Conure (Yuan et al., 2021) - - - - TIL User representations Tencent TL, Movielens English 9 tasks
+NAVER Shopping (Kim et al., 2023) - - - - TIL Search query prediction Tencent TL, Movielens English 6 tasks
+TRACE (Wang et al., 2024b) 40,000 - 16,000 56,000 TIL Mixed diverse tasks ScienceQA, FOMC, MeetingBank, C-STANCE, 20Minuten, CodeXGLUE, NumGLUE Hybrid 8 tasks
+ABSC (Ke and Xu, 2021) 3,452 150 1,120 4,722 DIL Aspect-based sentiment classification L5Domains, Liu3Domains, Ding9Domains, SemEval14 English 19 domains
+LAMOL (Sun et al., 2020a) 284,824 - 39,716 324,540 DIL Mixed classification and labeling tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers, SQuAD, WikiSQL, SST, QA-SRL, WOZ English 5 domains
+RVAE_LAMOL (Wang et al., 2022a) 15,870 - 5,668 21,538 DIL Generation tasks WOZ, QA-SRL, SST English 3 domains
+COPF (Zhang et al., 2023b) - - - - DIL Value alignment Standard Human Preference (SHP) English 18 domains
+C-PT (Zhu et al., 2022) 38,745 5,210 11,349 40,287 DIL Dialogue state tracking Schema-Guided Dialog dataset (SGD) English 19 domains
+CPT (Ke et al., 2022) 3,121,926 - - 3,121,926 DIL Domain-adaptive pre-training task Yelp Restaurant, AI Papers, ACL Papers, AGNews, SemEval-res, ACL-ARC, SCIERC English 4 domains
+CKL(Jang et al., 2022b) - - - 30,372 DIL Domain-adaptive pre-training task INVARIANTLAMA, UPDATEDLAMA, NEWLAMA, NEWLAMA-EASY English 3 domains
+ELLE (Qin et al., 2022) - - - - DIL Domain-adaptive pre-training task BOOKCORPUS(WB), NEWS ARTICLES(NS), AMAZON REVIEWS(REV), BIOMEDICAL PAPERS(BIO), COMPUTER SCIENCE PAPERS(CS) English 5 domains
+Domain-incremental Paper Stream (Jin et al., 2022) - - - - DIL Mixed 2 information extraction tasks S2ORC English 4 domains
+Chronologically-ordered Tweet Stream (Jin et al., 2022) - - - - DIL Mixed 2 classification tasks Tweets English 4 domains
+AdapterCL (Madotto et al., 2020) 31,426 4,043 4,818 40,287 DIL Mixed 4 dialogue tasks TaskMaster2019(TM19), TaskMaster2020(TM20), Schema Guided Dialogue(SGD), MultiWoZ English 37 domains
+DE&E (Wójcik et al., 2023) 28,982 - 12,089 41,071 CIL Text classification BBC News, Newsgroups, Complaints English 3 tasks
+EPI (Wang et al., 2023c) 12,840 3,524 6,917 23,281 CIL Mixed 2 classification tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers, Web of Science (WOS) English 13 classes
+PAGeR (Varshney et al., 2022) 59,754 7,115 15,304 82,173 CIL Intent classification CLINC150, HWU64, BANKING77, Stackoverflow S20, SGD, MWOZ, FewRel English 355 classes
+PLE (Li et al., 2022b) 4,669 4,650 31,642 40,961 CIL Intent classification CLINC150, ATIS, HWU64, BANKING77, MTOP, SNIPS, LEYZER, MSLU, TOP English 477 classes
+ExtendNER (Monaikul et al., 2021) 130,495 22,717 16,077 169,289 CIL Named Entity Recognition CoNLL-03, OntoNotes English 4 classes
 Online
-Foundational text classification (de Masson D’Autume et al., 2019) 115,000 - 7,600 122,600 Hard and Blurry News classification, sentiment analysis, Wikipedia article classification, and question-and-answer categorization English 5 tasks
-MBPA++ (de Masson D’Autume et al., 2019) 881,000 35,000 38,000 954,000 Hard and Blurry News classification, sentiment analysis, Wikipedia article classification, questions and answers categorization, question answering English 9 tasks
-Lifelong FewRel (Holla et al., 2020) - - - - Hard and Blurry few-shot relation detection English 10 tasks
-Firehose (Hu et al., 2022) - - - 110,000,000 Blurry Personalized online language learning English 1 tasks
-TemporalWiki (Jang et al., 2022a) - - - - - - English -
+MBPA++ (de Masson D’Autume et al., 2019) 115,000 - 7,600 122,600 Hard and Blurry Mixed 5 classification tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers English 5 tasks
+MBPA++ (de Masson D’Autume et al., 2019) 306,000 35,000 - 341,000 Hard and Blurry Question answering QuAC, SQuAD, Trivia Wikipedia, Trivia Web English 4 tasks
+Lifelong FewRel (Holla et al., 2020) - - - - Hard and Blurry Few-shot relation detection FewRel English 10 tasks
+Firehose (Hu et al., 2022) - - - 110,000,000 Blurry Personalized online language learning Twitter English 1 tasks
+TemporalWiki (Jang et al., 2022a) - - - - - - Wikipedia English -
 7. Datasets
+In this section, we review the typical datasets for offline and online continual learning for different tasks (Table 1).
 7.1. Offline Datasets for NLP
@@
 Particularly, the AG News dataset has 4 classes for news classification; the Amazon and Yelp dataset has 5 classes for sentiment analysis; the DBpedia dataset has 14 classes for Wikipedia text classification; and the Yahoo dataset has 10 classes for Q&A classification.
-The text classification benchmark includes 115,000 training and 7,600 test examples for each task, holding out 500 samples per class from the training set for validation.
 Building upon this, Razdaibiedina et al.
@@
 This benchmark not only utilizes the foundational text classification benchmark but also integrates additional datasets from the GLUE benchmark (Wang et al., 2018), SuperGLUE benchmark (Wang et al., 2019a), and the IMDB dataset (Maas et al., 2011).
-Specifically, the GLUE benchmark datasets included are MNLI, QQP, RTE, and SST2, focusing on tasks like natural language inference, paraphrase detection, and sentiment analysis.
-Similarly, the SuperGLUE datasets—WiC, CB, COPA, MultiRC, and BoolQ—encompass tasks ranging from word sense disambiguation to question answering.
 DE&E (Wójcik et al., 2023) uses three common text classification data sets with different characteristics-News-groups, BBC News, and Consumer Finance Complaints2.
@@
 In contrast, the near-domain group uses the Web of Science (WOS) (Kowsari et al., 2017) and 20 Newsgroups (Lang, 1995), which are restructured according to their high inter-task relevance.
-The WOS dataset comprises seven parent classes, each with five closely related sub-classes, while the 20 Newsgroups dataset, containing six news topics, is reorganized into four tasks to maximize inter-task correlation.
 Intent Classification.
 Some studies focus on intent classification tasks, where the classes are quite different in various domains or scenarios.
-In the realm of intent classification and detection, several datasets have been specifically designed to advance the field by addressing different challenges and providing diverse environments for model training and evaluation.
 The dataset, as introduced in PAGeR (Varshney et al., 2022), aims to tackle the lifelong intent detection problem by combining three public intent classification datasets (CLINC150 (Larson et al., 2019), HWU64 (Liu et al., 2021a), BANKING77 (Casanueva et al., 2020)), one text classification dataset (Stackoverflow S20 (Xu et al., 2017)), and two public multidomain dialog intent detection datasets (SGD (Rastogi et al., 2020), MWOZ (Budzianowski et al., 2018)).
@@
 Conversely, the dataset compiled in PLE (Li et al., 2022b) consolidates nine well-regarded intent detection datasets, including CLINC150 (Larson et al., 2019) and HWU64 (Liu et al., 2021a), among others, arranged in a fixed random sequence to form a standardized benchmark.
-This dataset emphasizes the importance of consistency and comparability in performance evaluations across different intent detection models, providing a platform for assessing and enhancing various methodologies.
-The dataset described by MeLL (Wang et al., 2021a) specifically addresses intent detection within two distinct contexts: task-oriented dialogues (TaskDialog-EUIC) and real-world e-commerce interactions (Hotline-EUIC).
+The dataset described by MeLL (Wang et al., 2021) specifically addresses intent detection within two distinct contexts: task-oriented dialogues (TaskDialog-EUIC) and real-world e-commerce interactions (Hotline-EUIC).
 TaskDialog-EUIC integrates data from Snips (Coucke et al., 2018), TOP semantic parsing (Gupta et al., 2018), and Facebook’s Multilingual Task Oriented Dataset (Schuster et al., 2018) into 90 tasks with overlapping label sets, amounting to over ten thousand samples.
-Hotline-EUIC is derived from an e-commerce dialogue system (Li et al., 2017) and the hotline audios are transcribed to text by a high-accuracy industrial Automatic Speech Recognition (ASR) system.
 Fine-grained Sentiment Analysis.
 Ke et al.
-(Ke and Xu, 2021) developed a task incremental learning dataset for aspect-based sentiment classification (ABSC).
+(Ke and Xu, 2021) develop a task incremental learning dataset for aspect-based sentiment classification (ABSC).
 This dataset aggregates reviews from four distinct sources, thereby enhancing its diversity and applicability across multiple domains.
 The sources include the L5Domains dataset by Hu et al.
-(Hu and Liu, 2004), which features consumer reviews for five different products; the Liu3Domains dataset by Liu (Liu et al., 2015), comprising reviews pertaining to three products; the Ding9Domains dataset by Ding et al.
-(Ding et al., 2008), which includes reviews of nine varied products; and the SemEval14 dataset, which is focused on reviews of two specific products—laptops and restaurants.
+(Hu and Liu, 2004), the Liu3Domains dataset (Liu et al., 2015), the Ding9Domains dataset (Ding et al., 2008), and the SemEval14 dataset.
 7.1.2. Datasets for Generation.
-In the rapidly advancing field of machine learning, diverse datasets function as crucial benchmarks for exploring various dimensions of language and code generation.
-These datasets address both universal and task-specific challenges, enabling a comprehensive evaluation of model capabilities.
-A particularly significant dataset highlighted in the work by Continual-T0 (Scialom et al., 2022b) focuses on English language generation tasks, including text simplification and empathetic dialogue generation, among others (Camburu et al., 2018; Bin Tareaf, 2017).
-The design of this dataset maintains uniformity in size, facilitating effective comparative analyses of performance across distinct tasks by ensuring a consistent volume of data for training.
+Diverse datasets function as crucial benchmarks for exploring various dimensions of language and code generation.
+A particularly significant dataset highlighted in the work by Continual-T0 (Scialom et al., 2022a) focuses on English language generation tasks, including text simplification and empathetic dialogue generation, among others (Camburu et al., 2018; Bin Tareaf, 2017).
 In a subsequent study, Luo et al.
@@
 d.]) and the foundational text classification benchmark (Zhang et al., 2015).
-This dataset encompasses five distinct NLP tasks originally sourced from DecaNLP: question answering, semantic parsing, sentiment analysis, semantic role labeling, and goal-oriented dialogue.
-For the purposes of this dataset, all tasks, whether derived from DecaNLP or the foundational text classification benchmark, are restructured into a uniform format, conceptualized under the framework of a question answering task.
-Moreover, the dataset devised in RVAE_LAMOL (Wang et al., 2022a), employs three tasks from DecaNLP: the English Wizard of Oz (WOZ) for goal-oriented dialogue, QA-SRL for semantic role labeling in a SQuAD-style format, and SST, which is a binary version of the Stanford Sentiment Treebank categorizing sentiments as positive or negative.
-These tasks are specifically treated as sequence generation tasks.
-The dataset introduced in COPR (Zhang et al., 2023b) represents a pioneering effort in applying both Task Incremental Learning (TIL) and Domain Incremental Learning (DIL) within the context of benchmarks that utilize existing human preferences.
+Moreover, the dataset devised in RVAE_LAMOL (Wang et al., 2022a), employs three tasks from DecaNLP: the English Wizard of Oz (WOZ) for goal-oriented dialogue, QA-SRL for semantic role labeling in a SQuAD-style format, and SST, which is a binary version of the Stanford Sentiment Treebank.
+The dataset introduced in COPF (Zhang et al., 2023b) represents a pioneering effort in applying both TIL and DIL within the context of benchmarks that utilize existing human preferences.
 Specifically, the TIL framework in this dataset mandates that the model sequentially acquires knowledge from three distinct tasks.
@@
 (Ethayarajh et al., 2022).
-The dataset described in Adaptive Compositional Modules (Zhang et al., 2022b) explores sequence generation and categorizes tasks into ”similar” and ”dissimilar” groups based on their characteristics.
+The dataset described in ACM (Zhang et al., 2022) explores sequence generation and categorizes tasks into “similar” and “dissimilar” groups based on their characteristics.
 Tasks classified as similar, including E2ENLG (Novikova et al., 2017) and four domains (restaurant, hotel, TV, laptop) from RNNLG (Wen et al., 2015), demonstrate shared patterns and are tested across four sequence orders, comprising a total of five tasks.
-In contrast, dissimilar tasks such as WikiSQL (SQL query generation) (Zhong et al., 2017), CNN/DailyMail (news article summarization) (See et al., 2017), and MultiWOZ (semantic state sequence generation) (Budzianowski et al., 2018) exhibit significant distributional shifts from previously encountered tasks.
+In contrast, dissimilar tasks such as WikiSQL (Zhong et al., 2017), CNN/DailyMail (See et al., 2017), and MultiWOZ (Budzianowski et al., 2018) exhibit significant distributional shifts from previously encountered tasks.
 The CODETASKCL dataset, explored by Yadav et al.
@@
 7.1.3. Datasets for Information Extraction.
-In the realm of natural language processing (NLP), various datasets are tailored to specific aspects of the task under continual learning paradigms.
 The dataset introduced in ExtendNER (Monaikul et al., 2021), exemplifies a continual learning approach to Named Entity Recognition (NER).
@@
 7.1.4. Datasets for Continual Pre-training.
-In the realm of continual pre-training for large language models (LMs), the development and utilization of specialized benchmarks play a pivotal role in evaluating and enhancing the effectiveness of continual learning systems.
+In the realm of continual pre-training for LLMs, the development and utilization of specialized benchmarks play a pivotal role in evaluating and enhancing the effectiveness of continual learning systems.
 The dataset, introduced in CPT (Ke et al., 2022), primarily focuses on the continual post-training of LMs across a series of domain-specific, unlabeled datasets.
@@
 Whereas the aforementioned two datasets assess more controlled dimensions of knowledge integration and retention, the dataset introduced in ELLE (Qin et al., 2022) focuses on the dynamic scenario of accumulating streaming data from diverse sources in a lifelong learning context.
-This dataset mirrors the real-world challenge of a language model (LM) that must continuously adapt to new data inflows from multiple domains, including BOOKCORPUS (WB) (Zhu et al., 2015), NEWS ARTICLES (NS) (Zellers et al., 2019), AMAZON REVIEWS (REV) (He and McAuley, 2016), BIOMEDICAL PAPERS (BIO) (Lo et al., 2020) and COMPUTER SCIENCE PAPERS (CS) (Lo et al., 2020).
+This dataset mirrors the real-world challenge of a LM that must continuously adapt to new data inflows from multiple domains, including BOOKCORPUS (Zhu et al., 2015), NEWS ARTICLES (Zellers et al., 2019), AMAZON REVIEWS (He and McAuley, 2016), BIOMEDICAL PAPERS (Lo et al., 2020) and COMPUTER SCIENCE PAPERS (Lo et al., 2020).
 The benchmark evaluates the LM’s capacity to effectively integrate new information from these varied sources over time, highlighting the essential need for LMs to evolve in response to continual data growth and shifts in data distribution.
@@
 7.1.5. Datasets for Hybrid Tasks.
-An increasing number of datasets are adopting a hybrid task approach that integrates multiple learning paradigms and task types, aimed at testing and enhancing the adaptability of models.
-A notable example is the dataset introduced in AdapterCL (Madotto et al., 2020), which is tailored for task-oriented dialogue systems.
-This dataset incorporates four task-oriented datasets: TaskMaster 2019 (TM19) (Byrne et al., 2019), TaskMaster 2020 (TM20) (Byrne et al., 2019), Schema Guided Dialogue (SGD) (Rastogi et al., 2020), and MultiWoZ (Budzianowski et al., 2018).
-These datasets have been pre-processed to form a curriculum encompassing 37 domains, structured under four continual learning settings: INTENT classification, Dialogue State Tracking (DST), Natural Language Generation (NLG), and end-to-end (E2E) modeling.
+An increasing number of datasets are adopting a hybrid task approach that integrates multiple learning paradigms and task types, aimed at enhancing the adaptability of models.
+AdapterCL (Madotto et al., 2020) introduces a dataset for task-oriented dialogue systems, which incorporates four task-oriented datasets: TaskMaster 2019 (Byrne et al., 2019), TaskMaster 2020 (Byrne et al., 2019), Schema Guided Dialogue (Rastogi et al., 2020), and MultiWoZ (Budzianowski et al., 2018).
 Continual Instruction Tuning Benchmark (CITB) (Zhang et al., 2023a) extends the concept of continual learning by focusing on instruction-based NLP tasks.
 Built on the comprehensive SuperNI (Wang et al., 2022c) dataset, it includes over 1,600 tasks across diverse NLP categories.
-CITB differentiates itself by formulating two distinct streams—InstrDialog and InstrDialog++—to examine how models integrate and retain new dialogue-oriented and varied NLP tasks under continual learning settings.
-This benchmark suite not only tests task retention and adaptability but also explores how instruction tuning can be optimized for a continual learning framework.
-The ConTinTin (Yin et al., 2022) is an adaptation of the NATURAL-INSTRUCTIONS dataset, specifically restructured to facilitate a continual learning framework.
-This adaptation involves decomposing the original crowdsourcing instructions into smaller, distinct sub-tasks to create a new dataset.
-Additionally, the new dataset incorporates a novel experimental design where tasks are selected randomly to create diverse sequences, enabling the evaluation of a model’s adaptability to novel instructions without prior exposure.
+The ConTinTin (Yin et al., 2022) is an adaptation of the NATURAL-INSTRUCTIONS dataset, specifically restructured to facilitate a CL framework.
 The dataset, used in Conure (Yuan et al., 2021), consists of Tencent TL (TTL) (Yuan et al., 2020) and Movielens (ML).
 The TTL dataset is designed to address three item recommendation tasks and three user profiling tasks, whereas the ML dataset exclusively focuses on three item recommendation tasks.
-Both datasets have been pre-processed to facilitate a continual learning framework, simulating environments where models must adapt to evolving data streams.
 Furthermore, Kim et al.
@@
 Finally, the TRACE dataset, introduced by Wang et al.
-(Wang et al., 2024a), is specifically designed to bridge the existing gap in the evaluation of large language models (LLMs) within the continual learning framework, encompassing a wide range of complex and specialized tasks.
-Distinguished from other datasets, TRACE targets domain-specific tasks that are multilingual and technical, including code completion and mathematical reasoning.
-This diversity presents a unique set of challenges that span both specialized and broad dimensions.
-Moreover, TRACE rigorously assesses the models’ ability to sustain performance across tasks that demand different knowledge bases and cognitive skills.
-This evaluation highlights the essential need for adaptability in LLMs trained to operate under continual learning conditions, underscoring their potential in dynamic real-world applications.
+(Wang et al., 2024b), is specifically designed to bridge the existing gap in the evaluation of LLMs within the CL framework, encompassing a wide range of complex and specialized tasks.
 7.2. Online Datasets for NLP
@@
 The foundational text classification benchmark, as introduced by Zhang et al.
-(Zhang et al., 2015) has traditionally been applied in offline continual learning settings.
-Recent advancements have adapted this benchmark for online continual learning, notably in studies such as MBPA++ (de Masson D’Autume et al., 2019) and OML-ER (Holla et al., 2020)
+(Zhang et al., 2015) has traditionally been applied in offline CL settings.
+Recent advancements have adapted this benchmark for online CL, notably in studies such as MBPA++ (de Masson D’Autume et al., 2019) and OML-ER (Holla et al., 2020)
 7.2.2. Datasets for Generation.
@@
 The model’s performance is subsequently tested on the Flickr30K, MS-COCO, and EC datasets.
-6.4. Offline and Online Datasets for Other Tasks
-In this paper, we provide a comprehensive review of recent studies in natural language processing (NLP) and multi-modal tasks, with a particular focus on the use of PLMs, LLMs and VLMs.
-Additionally, we introduce a range of offline continual learning (CL) datasets employed across various applications, including automatic speech recognition (Vander Eeckt et al., 2023; Della Libera et al., 2023), autonomous driving (Verwimp et al., 2023), disease classification (Derakhshani et al., 2022), reinforcement learning (Wołczyk et al., 2021), graph data (Ko et al., 2022), and computer vision (Faber et al., 2023; Lin et al., 2021).
-Furthermore, some online CL benchmarks are also proposed for computer version tasks, such as continual visual learning (Cai et al., 2021), continual object detection (Wang et al., 2021b).
+LAION-400M (Schuhmann et al., 2021) and LAION-5B (Schuhmann et al., 2022) are both CLIP-filtered, large-scale, open-access datasets designed to support multimodal models like CLIP, but they differ significantly in scale and scope.
+LAION-400M contains 400 million image-text pairs primarily in English, while LAION-5B is a much larger and more diverse dataset with 5.85 billion pairs, including multilingual and language-agnostic examples.
+CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021) are large-scale datasets for vision-and-language pre-training, differing in scale, filtering, and diversity.
+CC3M has 3.3 million image-text pairs with strict filtering for high-precision captions, while CC12M, with 12.4 million pairs, relaxes these filters to capture a broader range of concepts, especially long-tail categories.
+COYO-700M (Byeon et al., 2022) contains 747 million image-text pairs, designed to train various models with additional meta-attributes.
+Datacomp (Gadre et al., 2024) is a benchmark designed to advance the creation of multimodal datasets, particularly for image-text models such as CLIP.
 Figure 9. Illustration of calculating metrics.
@@
 8.2. Memory Stability.
-Memory stability is typically evaluated by backward transfer (BWT) (Lopez-Paz and Ranzato, 2017) and forgetting measure (FM) (Chaudhry et al., 2018).
+Memory stability is typically evaluated by backward transfer (Lopez-Paz and Ranzato, 2017) and forgetting measure (Chaudhry et al., 2018).
 Backward Transfer (BWT) emerges as a pivotal concept extensively documented in the literature, notably by Lopez et al.
@@
 Davari et al.
-(Davari et al., 2022) propose a method named linear probes (LP) to assess representation forgetting.
-This approach measures the effectiveness of learned representations via an optimal linear classifier trained on the frozen activations of a base network.
+(Davari et al., 2022) propose linear probes (LP) to assess representation forgetting via an optimal linear classifier trained on the frozen activations of a base network.
 Representation forgetting is quantified by evaluating the change in Language Processing (LP) performance before and after the introduction of a new task.
@@
 (6) $FWT=\frac{1}{N-1}\sum_{i=2}^{N}R_{i-1,i}-R_{0,i}$
-where $R_{0,i}$ denotes the performance metric associated with training on task $i$ independently.
-Higher values of FWT indicate superior model performance.
-It is important to note that discussing backward transfer for the initial task is not applicable, as there are no preceding tasks to influence its performance.
+where $R_{0,i}$ denotes the performance metric associated with training on task $i$ independently. Higher values of FWT indicate superior model performance.
 Intransigence measure (IM), as defined by Chaudhry et al.
-(Chaudhry et al., 2018), quantifies a model’s inability to learn new tasks.
-This measure is calculated by comparing the performance difference of a task when trained jointly with other tasks versus when trained in a continual learning setting.
+(Chaudhry et al., 2018), quantifies a model’s inability to learn new tasks by comparing the performance difference of a task when trained jointly with other tasks versus.
 Moreover, Koh et al.
-(Koh et al., 2022) introduce novel metrics, known as Knowledge Gain Ratio (KGR), which quantifies the capacity to acquire new knowledge through the calculation of knowledge gain.
+(Koh et al., 2022) introduce novel metrics, known as Knowledge Gain Ratio (KGR), which quantifies the capacity to acquire new knowledge by calculating knowledge gain.
+Table 2. Comparison of typical continual learning methods, where Reg. and Para. mean Regularization and Parameter-isolation.
+CL settings Method FM Arch. Continual Learning Tech.
+Traditional Continual Pre-training Parameter-Efficient Instruction Tuning
+DIL LFPT5 (Qin and Joty, 2022b) T5 Replay — Reg. — Para.
+B-CL (Ke and Xu, 2021) BERT Para.
+ELLE (Qin et al., 2022) BERT — GPT Reply — Para.
+AdapterCL (Madotto et al., 2020) GPT2 Para.
+DEMIX (Gururangan et al., 2022) GPT2 — GPT3 Para.
+CLASSIC (Ke et al., 2021b) BERT Para.
+LLM-CL (Ding et al., 2024) LLaMA Replay — Reg. — Para.
+CPT (Ke et al., 2022) RoBERTa Para.
+C-PT (Zhu et al., 2022) T5 Para.
+CL-KD (Castellucci et al., 2021) BERT Reg.
+PlugLM (Cheng et al., 2023b) BERT Para.
+Pretr (Cossu et al., 2022) BERT — RoBERTa Para.
+AEWC (Lee, 2017) LSTM-RNNs Reg.
+Continual DAP-training (Ke et al., 2023) RoBERTa Reg.
+COPF (Zhang et al., 2023b) LLaMA Replay — Reg. — Para.
+LAMOL (Sun et al., 2020a) GPT2 Replay
+RVAE_LAMOL (Wang et al., 2022a) GPT2 Replay
+Adapt-Retrieve-Revise (Zhang et al., 2023c) Baichuan RAG
+Lifelong-MoE (Chen et al., 2023) GLaM Para.
+CPPO (Zhang et al., 2024b) GPT2 Reg.
+EcomGPT-CT (Ma et al., 2023) BLOOM Replay
+S-Prompt (Wang et al., 2022b) CLIP Para.
+VQACL (Zhang et al., 2023d) VL-T5 Replay — Reg.
+TIL PP (Razdaibiedina et al., 2023) BERT — T5 Para.
+CTR (Ke et al., 2021a) BERT Para.
+MeLL (Wang et al., 2021) BERT — RoBERTa Replay
+ERDA (Qin and Joty, 2022a) Bi-LSTM — BERT Replay — Reg.
+PCLL (Zhao et al., 2022) GPT2 Replay
+BiHNet-Reg (Jin et al., 2021) BART — BART-Adapter Para.
+ConTinTin (Yin et al., 2022) BART Replay — InstructionSpeak
+HMI (Maekawa et al., 2023) BERT, GPT2 Replay
+ACM (Zhang et al., 2022) GPT2 Replay — Para.
+DYNAINST (Mok et al., 2023) BART Replay — Reg.
+Conure (Yuan et al., 2021) TCN Para.
+TERACON (Kim et al., 2023) NextitNet Para.
+ERNIE 2.0 (Sun et al., 2020b) ERNIE Replay
+RecyclableTuning (Qin et al., 2023) RoBERTa Replay — Para.
+Conpet (Song et al., 2023) LLaMA Replay — Para.
+InstructAlign (Cahyawijaya et al., 2023) BLOOMZ Replay
+Continual-T0 (Scialom et al., 2022b) T0 Replay
+DynaMind (Du et al., 2023) LLaMA — GPT — Falcon Replay — Para.
+ELM (Jang et al., 2023) T5 Para.
+O-LoRA (Wang et al., 2023a) LLaMA — Alpaca Para.
+COPF (Zhang et al., 2023b) LLaMA Replay — Reg. — Para.
+SLM (PENG et al., 2024) LLaMA, T5, BERT Para.
+CTP (Zhu et al., 2023) BERT — ViT Para. -
+ZSCL (Zheng et al., 2023) CLIP Replay — Reg.
+MoE-Adapters4CL (Yu et al., 2024b) CLIP Para.
+TRIPLET (Qian et al., 2023) ALBEF — FLAVA Para.
+RecAdam (Chen et al., 2020) BERT — ALBERT Reg.
+CIL EPI (Wang et al., 2023c) BERT Para.
+IDBR (Huang et al., 2021) BERT Replay
+PAGeR (Varshney et al., 2022) GPT2 Replay
+ENTAILMENT (Xia et al., 2021) RoBERTa prompt-based
+ExtendNER (Monaikul et al., 2021) BERT Reg.
+PLE (Li et al., 2022b) RoBERTa Replay — Para.
+DE&E (Wójcik et al., 2023) Distilbert Para.
+SRC (Liu et al., 2019b) SIF Reg.
+MoE-Adapters4CL (Yu et al., 2024b) TIL RandSel — KMeansSel
+VLM-PL (Kim et al., 2024) CLIP — Ferret Replay
+Adaptation-CLIP (Liu et al., 2023a) CLIP Para.
+PROOF (Zhou et al., 2023c) CLIP Para.
+LGCL (Khan et al., 2023) CLIP — ViT Para.
+ZSCL (Zheng et al., 2023) CLIP Replay — Reg.
+CLAP (Jha et al., 2024) CLIP Replay — Para.
+GMM (Cao et al., 2024) EVA-CLIP — BLIP2 Reg.
+Online CL MBPA++ (de Masson D’Autume et al., 2019) BERT Replay
+Meta-MBPA++ (Wang et al., 2020) BERT Replay
+OML-ER (Holla et al., 2020) BERT Replay
+TPEM (Geng et al., 2021) GRU Para.
+MSR (Liu et al., 2021b) BERT Replay — Reg.
+ProgModel (Shen et al., 2019) RNN Para.
+DKR (Cui et al., 2024) CLIP Reg.
+GMM (Cao et al., 2024) EVA-CLIP — BLIP2 — MiniGPT-4 Para.
 8.4. Metrics for Continual Pre-training.
-CKL (Jang et al., 2022b) introduces a novel metric, named FUAR (FORGOTTEN / (UPDATED + ACQUIRED) RATIO), which quantitatively measures the efficiency of each CKL method.
-It calculates the number of instances of time-invariant knowledge that a model forgets in order to learn or update one instance of new knowledge.
-When FUAR is equal to 1.0, it signifies an equilibrium where one time-invariant knowledge instance is forgotten on average to obtain a new or updated knowledge instance.
+CKL (Jang et al., 2022b) introduces FUAR, which quantitatively measures the efficiency of each CKL method.
+It calculates the instances of time-invariant knowledge that a model forgets to learn or update one instance of new knowledge.
+Table 3.
+A summary of the performance of various Foundation LMs.
+A-RG: Average of ROUGE-1, -2 and -L scores; F1: F1 score; Acc.: Accuracy; MF1: Macro-F1; AMF1: averaged Macro-F1; JGA: Joint Goal Accuracy; SER: Slot Error Rate; AA: average accuracy; AIA: average incremental accuracy; FM: forgetting measure.
+Methods Dataset F1 Acc A-RG MF1 JGA SER BLEU AA AIA FM FWT BWT AMF1
+LFPT5 (Qin and Joty, 2022b) AGNews, Amazon, DBPedia, Yahoo - 52.71 - - - - - - - - - - -
+SLM(T5) (PENG et al., 2024) AGNews, Amazon, DBpedia, Yahoo - - - - - - - 73.10 - - - - -
+PP(T5) (Razdaibiedina et al., 2023) AGNews, Amazon, DBpedia, Yahoo - - - - - - - 75.10 - - - - -
+MBPA++ (de Masson D’Autume et al., 2019) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 70.60 - - - - -
+IDBR (Huang et al., 2021) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 73.19 - - - - -
+EPI (Wang et al., 2023c) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 74.43 - - - - -
+OML-ER (Holla et al., 2020) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 75.70 - - - - -
+O-LoRA(Wang et al., 2023a) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 75.80 - - - - -
+LAMOL (Sun et al., 2020a) AGNews, Amazon, DBPedia, Yahoo, Yelp - - - - - - - 76.50 - - - - -
+PP(Bert) (Razdaibiedina et al., 2023) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 77.90 - - - - -
+Meta-MBPA++ (Wang et al., 2020) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 77.30 - - - - -
+SLM(Bert) (PENG et al., 2024) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 79.10 - - - - -
+B-CL (Ke and Xu, 2021) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 88.29 - 81.40 - - - - - - - - -
+CLASSIC (Ke et al., 2021b) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 90.22 - 85.12 - - - - - - - - -
+CTR (Ke et al., 2021a) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 89.47 - 83.62 - - - - - - - - -
+LLM-CT (Ding et al., 2024) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 94.91 - 91.43 - - - - - - - - -
+PP(Bert) (Razdaibiedina et al., 2023) MNLI, CB, WiC, COPA, QQP, BoolQ, RTE, IMDB, Yelp, Amazon, SST2, DBpedia, AGNews, Yelp, Amazon, Yahoo - - - - - - - 69.3 - - - - -
+O-LoRA(Wang et al., 2023a) MNLI, CB, WiC, COPA, QQP, BoolQ, RTE, IMDB, Yelp, Amazon, SST2, DBpedia, AGNews, Yelp, Amazon, Yahoo - - - - - - - 69.6 - - - - -
+PP(T5) (Razdaibiedina et al., 2023) MNLI, CB, WiC, COPA, QQP, BoolQ, RTE, IMDB, Yelp, Amazon, SST2, DBpedia, AGNews, Yelp, Amazon, Yahoo - - - - - - - 79.5 - - - - -
+OML-ER (Holla et al., 2020) FewRel - 69.5 - - - - - - - - - - -
+DYNAMIC CONPET(Song et al., 2023) FewRel - 88.62 - - - - - 88.62 - - - - -
+PAGeR (Varshney et al., 2022) FewRel - - - - - - - 91.30 - - - - -
+CPPO (Zhang et al., 2024b) IMDB, HH-RLHF, Reddit TL, DR - - - - - - - 77.40 84.20 2.9 - -2.9 -
+COPF (Zhang et al., 2023b) IMDB, HH-RLHF, Reddit TL, DR - - - - - - - 86.30 85.10 -0.60 - 0.60 -
+LAMOL (Sun et al., 2020a) SST, QA-SRL, WOZ - - - - - - - 81.00 - - - - -
+RVAE_LAMOL (Wang et al., 2022a) SST, QA-SRL, WOZ - - - - - - - 81.60 - - - - -
+PLE(5-shot regime) (Li et al., 2022b) CLINC150, ATIS, HWU64, BANKING77, MTOP, SNIPS, LEYZER, MSLU, TOP - - - - - - - 85.61 - - - - -
+PLE(10-shot regime) (Li et al., 2022b) CLINC150, ATIS, HWU64, BANKING77, MTOP, SNIPS, LEYZER, MSLU, TOP - - - - - - - 88.16 - - - - -
+MBPA++ (de Masson D’Autume et al., 2019) QuAC, TrWeb, TrWik, SQuAD - - - - - - - - - - - - 62.40
+Meta-MBPA++ (Wang et al., 2020) QuAC, TrWeb, TrWik, SQuAD - - - - - - - - - - - - 64.90
+CPT (Ke et al., 2022) Yelp Restaurant, AI Papers, ACL Papers, AGNews - 52.59 - 46.41 - - - - - - - - -
+Continual DAP-training (Ke et al., 2023) Yelp Restaurant, Amazon Phone, Amazon Camera, ACL Papers, AI Papers, PubMed Papers - 81.91 - 77.93 - - - - - - - - -
+DYNAMIC CONPET(Song et al., 2023) OntoNotes - 84.47 - - - - - 85.83 - - - - -
+LFPT5 (Qin and Joty, 2022b) CoNLL03, OntoNotes 47.59 - - - - - - - - - - - -
+AdapterCL (Madotto et al., 2020) TM19, TM20, SGD, MultiWoZ - 90.50 - - 35.10 31.78 16.76 - - - - - -
+RMR_ DSE (Li et al., 2022a) MultiWoZ-2.0 - - - - - 48.79 39.86 - - - - - -
+DEMIX (Gururangan et al., 2022) In-Car Assistant, Multi-WOZ 2.1, CamRest 45.84 - - - - - 11.96 - - - - - -
+LFPT5 (Qin and Joty, 2022b) CNNDM, WikiHow, XSum - - 17.05 - - - - - - - - - -
+C-PT (Zhu et al., 2022) Schema-Guided Dialog dataset (SGD) - - - - 61.20 - - - - - 13.70 0.50 -
+Lifelong-MoE (Chen et al., 2023) TriviaQA 20.22 - - - - - - - - - - - -
+Lifelong-MoE (Chen et al., 2023) WMT16 - - - - - - 19.16 - - - - - -
+CTR (Ke et al., 2021a) 10 DSC datasets - 89.31 - 88.75 - - - - - - - - -
+CTR (Ke et al., 2021a) 20News - 95.25 - 95.23 - - - - - - - - -
+MeLL (Wang et al., 2021) TaskDialog-EUIC - - - - - - - 93.79 - - - - 93.42
+MeLL (Wang et al., 2021) Hotline-EUIC - - - - - - - 96.73 - - - - 93.41
+PCLL (Zhao et al., 2022) HWU, BANKING, CLINC, SNIPS, TOP - - - - - - - 90.25 - - - - -
+PCLL (Zhao et al., 2022) SNIPS, AITS, DSTC, MIT-MOVIE, MIT-RESTAURANT - - - - - - - - - - - - 74.48
+SLM(LLaMA) (PENG et al., 2024) Medical, MMLU, Finance - - - - - - - 82.3 - - - - -
+EPI (Wang et al., 2023c) Web of Science (WOS) - - - - - - - 77.83 - - - - -
+IDBR (Huang et al., 2021) Yahoo, AGNews, Yelp - - - - - - - 72.53 - - - - -
 8.5. Online CL-Specific Metrics.
-Near-future accuracy (NFA) (Al Kader Hammoud et al., 2023) is introduced as a novel evaluation metric for OCL problem.
+Near-future accuracy (NFA) (Al Kader Hammoud et al., 2023) is introduced as a novel evaluation metric for OCL problems.
 Unlike traditional evaluation methods that assess models on immediately subsequent samples, NFA evaluates models on samples slightly further into the future, using a minimal shift $S$.
 Such operation can mitigate label correlation effects, which can adversely impact the accuracy of model adaptability assessments.
-The smallest shift $S$ is selected to ensure that the test sample aligns closely with the distribution of recently observed training data.
 Yogatama et al.
 (Yogatama et al., 2019) proposed a novel online codelength, inspired by prequential encoding (Blier and Ollivier, 2018), to quantify how quickly an existing model can adapt to a new task.
+9. Analysis and Discussion
+In this section, we compare the performance and characteristics of typical foundation model-based continual learning methods (Table 2 and Table 3).
+The effectiveness of each continual learning method is shaped by the specific task environment, resource constraints, and application goals.
+Replay-based methods like LAMOL, Meta-MBPA++, PAGeR, and COPF excel in scenarios where memory efficiency and maintaining a balance between old and new knowledge are essential.
+These methods strike a delicate balance between retaining previously acquired knowledge and adapting to new tasks.
+Parameter isolation is key to minimizing interference between highly distinct tasks.
+By allocating dedicated parameters to each task, these approaches prevent overlap and ensure that task-specific learning is preserved.
+Techniques such as AdapterCL, B-CL, and EPI are particularly effective in environments with highly variable tasks, where reducing interference is critical.
+Regularization techniques, on the other hand, enhance stability in structured task settings.
+These methods constrain model updates to protect knowledge from previous tasks, making them ideal in situations with well-defined task boundaries.
+Regularization-based approaches like RMR_DSE, CLASSIC, and EWC mitigate catastrophic forgetting by limiting model changes, thereby preserving prior learning.
+These methods are particularly effective in domain-incremental learning, where tasks evolve gradually, and long-term stability is vital.
+First, most continual learning (CL) methods are built on PLMs like BERT, RoBERTa, and BART, or smaller-scale LLMs such as LLaMA.
+Models leveraging LLMs often outperform PLM-based ones, thanks to their greater capacity for handling complex tasks and enhanced adaptability.
+Second, the majority of CL methods are tailored for offline settings, where the full target dataset is assumed to be available for training.
+However, in real-world scenarios, data typically arrives in a streaming fashion, making it challenging to anticipate distribution shifts and effectively manage dynamic data.
+Third, a significant hurdle in the field is the lack of unified benchmarks and standardized metrics for evaluating CL algorithms.
+This fragmentation complicates the comparison of methods across tasks and domains.
+To propel the field forward, there is a critical need for benchmark datasets and robust metrics that can consistently gauge the performance of CL models, particularly in dynamic, data-streaming environments.
+These benchmarks should span diverse tasks and modalities to ensure comprehensive evaluation across various domains.
 10. Challenges and Further Work
@@
 These assumptions do not hold in real-world applications, where environments continually evolve and introduce novel stimuli.
-A key challenge is to develop continual learning models that operate effectively in complex, noisy environments where clear labels are not always available, and task domains frequently change.
+A key challenge is developing continual learning models that can autonomously detect and adapt to data distribution shifts, thereby improving the applicability of AI in dynamic real-world scenarios.
 Liu et al.
 (Liu et al., 2023b) recently proposed the SOLA framework to address these limitations by facilitating autonomous adaptation in AI systems.
-Despite this progress, significant challenges remain in enabling these systems to independently adjust to new, dynamic environments without ongoing human oversight.
-Future research should focus on developing algorithms capable of autonomously detecting and adapting to shifts in data distribution, thereby improving the applicability of AI in dynamic real-world scenarios.
+Despite this progress, significant challenges remain in enabling these systems to adjust to new, dynamic environments without ongoing human oversight.
+Bridging Machine Learning and Cognitive Science in Continual Learning.
+While continual learning aims to mimic human-like learning, a gap persists in how well these methods align with cognitive science findings on human continuous learning.
+In human cognition, processes such as rehearsal, memory consolidation, and adaptive forgetting are key to learning in dynamic environments.
+In contrast, machine learning models often face challenges like catastrophic forgetting, a phenomenon without a clear biological equivalent.
+Bridging these fields could enhance the design of continual learning algorithms by drawing on cognitive science insights, potentially leading to more robust, adaptive, and human-like learning systems.
 Learning Knowledge from Conversation.
@@
 Techniques such as differential privacy (Dwork, 2006), federated learning (Zhang et al., 2021), and secure multi-party computation (Goldreich, 1998) offer promising solutions by allowing models to learn from decentralized data sources without needing to access the actual data directly.
-Future research in continual learning should not only focus on enhancing learning efficiency and adaptability but also prioritize the development of robust frameworks that safeguard user privacy across all phases of data handling and model updating.
+Continual Alignment.
+In dynamic real-world environments, human preferences and requirements are ever-evolving, making continual alignment crucial for maintaining AI system effectiveness.
+Traditional alignment methods often assume static preferences, which is insufficient in changing contexts.
+Integrating Reinforcement Learning from Human Feedback (RLHF) into continual learning presents a unique challenge, as models must adapt to new preferences while retaining previously learned ones.
+Future research should prioritize developing methods like CPPO and COPF that balance integrating new feedback with preserving past alignment, ensuring long-term adaptability to human needs.
 Robust Continual Learning.
@@
 This work hopes to inspire further research that will ultimately lead to more robust, efficient, and intelligent systems capable of lifelong learning.
-  • (1)
-  • Achiam et al.
-(2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al.
-2023.
-Gpt-4 technical report.
-arXiv preprint arXiv:2303.08774 (2023).
-  • Al Kader Hammoud et al.
-(2023) Hasan Abed Al Kader Hammoud, Ameya Prabhu, Ser-Nam Lim, Philip HS Torr, Adel Bibi, and Bernard Ghanem.
-2023.
-Rapid Adaptation in Online Continual Learning: Are We Evaluating It Right?.
-In ICCV.
-18852–18861.
-  • Aljundi et al.
-(2019) Rahaf Aljundi, Eugene Belilovsky, Tinne Tuytelaars, Laurent Charlin, Massimo Caccia, Min Lin, and Lucas Page-Caccia.
-2019.
-Online Continual Learning with Maximal Interfered Retrieval.
-In NeurIPS, Vol.
-32.
-  • Bai et al.
-(2022) Yuntao Bai, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, Nova DasSarma, Dawn Drain, Stanislav Fort, Deep Ganguli, Tom Henighan, et al.
-2022.
-Training a helpful and harmless assistant with reinforcement learning from human feedback.
-arXiv preprint arXiv:2204.05862 (2022).
-  • Bang et al.
-(2022) Jihwan Bang, Hyunseo Koh, Seulki Park, Hwanjun Song, Jung-Woo Ha, and Jonghyun Choi.
-2022.
-Online continual learning on a contaminated data stream with blurry task boundaries.
-In CVPR.
-9275–9284.
-  • Beaulieu et al.
-(2020) Shawn Beaulieu, Lapo Frati, Thomas Miconi, Joel Lehman, Kenneth O Stanley, Jeff Clune, and Nick Cheney.
-2020.
-Learning to continually learn.
-arXiv preprint arXiv:2002.09571 (2020).
-  • Belouadah et al.
-(2021) Eden Belouadah, Adrian Popescu, and Ioannis Kanellos.
-2021.
-A comprehensive study of class incremental learning algorithms for visual tasks.
-Neural Networks 135 (2021), 38–54.
-  • Biesialska et al. (2020) Magdalena Biesialska, Katarzyna Biesialska, and Marta R Costa-jussà. 2020. Continual Lifelong Learning in Natural Language Processing: A Survey. In COLING. 6523–6541.
-  • Bin Tareaf (2017) Raad Bin Tareaf. 2017. Tweets Dataset - Top 20 most followed users in Twitter social platform.
-  • Blier and Ollivier (2018) Léonard Blier and Yann Ollivier. 2018. The description length of deep learning models. NeurIPS 31 (2018).
-  • Bordes et al. (2015) Antoine Bordes, Nicolas Usunier, Sumit Chopra, and Jason Weston. 2015. Large-scale simple question answering with memory networks. arXiv preprint arXiv:1506.02075 (2015).
-  • Budzianowski et al.
-(2018) Paweł Budzianowski, Tsung-Hsien Wen, Bo-Hsiang Tseng, Inigo Casanueva, Stefan Ultes, Osman Ramadan, and Milica Gašić.
-2018.
-MultiWOZ–a large-scale multi-domain wizard-of-oz dataset for task-oriented dialogue modelling.
-arXiv preprint arXiv:1810.00278 (2018).
-  • Byrne et al.
-(2019) Bill Byrne, Karthik Krishnamoorthi, Chinnadhurai Sankar, Arvind Neelakantan, Daniel Duckworth, Semih Yavuz, Ben Goodrich, Amit Dubey, Andy Cedilnik, and Kyu-Young Kim.
-2019.
-Taskmaster-1: Toward a realistic and diverse dialog dataset.
-arXiv preprint arXiv:1909.05358 (2019).
-  • Caccia et al.
-(2021) Lucas Caccia, Rahaf Aljundi, Nader Asadi, Tinne Tuytelaars, Joelle Pineau, and Eugene Belilovsky.
-2021.
-New insights on reducing abrupt representation change in online continual learning.
-arXiv preprint arXiv:2104.05025 (2021).
-  • Cahyawijaya et al.
-(2023) Samuel Cahyawijaya, Holy Lovenia, Tiezheng Yu, Willy Chung, and Pascale Fung.
-2023.
-InstructAlign: High-and-Low Resource Language Alignment via Continual Crosslingual Instruction Tuning.
-In the First Workshop in South East Asian Language Processing.
-55–78.
-  • Cai et al. (2021) Zhipeng Cai, Ozan Sener, and Vladlen Koltun. 2021. Online continual learning with natural distribution shifts: An empirical study with visual data. In ICCV. 8281–8290.
-  • Camburu et al. (2018) Oana-Maria Camburu, Tim Rocktäschel, Thomas Lukasiewicz, and Phil Blunsom. 2018. e-snli: Natural language inference with natural language explanations. NeurIPS 31 (2018).
-  • Cao et al.
-(2024) Xusheng Cao, Haori Lu, Linlan Huang, Xialei Liu, and Ming-Ming Cheng.
-2024.
-Generative Multi-modal Models are Good Class-Incremental Learners.
-arXiv preprint arXiv:2403.18383 (2024).
-  • Casanueva et al.
-(2020) Iñigo Casanueva, Tadas Temčinas, Daniela Gerz, Matthew Henderson, and Ivan Vulić.
-2020.
-Efficient intent detection with dual sentence encoders.
-arXiv preprint arXiv:2003.04807 (2020).
-  • Castellucci et al. (2021) Giuseppe Castellucci, Simone Filice, Danilo Croce, and Roberto Basili. 2021. Learning to Solve NLP Tasks in an Incremental Number of Languages. In ACL. 837–847.
-  • Chaudhry et al.
-(2018) Arslan Chaudhry, Puneet K Dokania, Thalaiyasingam Ajanthan, and Philip HS Torr.
-2018.
-Riemannian walk for incremental learning: Understanding forgetting and intransigence.
-In ECCV.
-532–547.
-  • Chaudhry et al. (2019) Arslan Chaudhry, Marc’Aurelio Ranzato, Marcus Rohrbach, and Mohamed Elhoseiny. 2019. Efficient Lifelong Learning with A-GEM. In ICLR.
-  • Chen et al.
-(2024) Cheng Chen, Junchen Zhu, Xu Luo, Hengtao Shen, Lianli Gao, and Jingkuan Song.
-2024.
-CoIN: A Benchmark of Continual Instruction tuNing for Multimodel Large Language Model.
-arXiv preprint arXiv:2403.08350 (2024).
-  • Chen et al.
-(2020) Hung-Jen Chen, An-Chieh Cheng, Da-Cheng Juan, Wei Wei, and Min Sun.
-2020.
-Mitigating forgetting in online continual learning via instance-aware parameterization.
-NeurIPS 33 (2020), 17466–17477.
-  • Chen et al.
-(2023) Wuyang Chen, Yanqi Zhou, Nan Du, Yanping Huang, James Laudon, Zhifeng Chen, and Claire Cui.
-2023.
-Lifelong language pretraining with distribution-specialized experts.
-In ICML.
-PMLR, 5383–5395.
-  • Cheng et al. (2023a) Daixuan Cheng, Shaohan Huang, and Furu Wei. 2023a. Adapting large language models via reading comprehension. arXiv preprint arXiv:2309.09530 (2023).
-  • Cheng et al. (2023b) Xin Cheng, Yankai Lin, Xiuying Chen, Dongyan Zhao, and Rui Yan. 2023b. Decouple knowledge from paramters for plug-and-play language modeling. In ACL. 14288–14308.
-  • Choi et al. (2018) Eunsol Choi, He He, Mohit Iyyer, Mark Yatskar, Wen-tau Yih, Yejin Choi, Percy Liang, and Luke Zettlemoyer. 2018. QuAC: Question Answering in Context. In EMNLP. 2174–2184.
-  • Chrysakis and Moens (2023) Aristotelis Chrysakis and Marie-Francine Moens. 2023. Online bias correction for task-free continual learning. ICLR (2023).
-  • Cossu et al.
-(2022) Andrea Cossu, Tinne Tuytelaars, Antonio Carta, Lucia Passaro, Vincenzo Lomonaco, and Davide Bacciu.
-2022.
-Continual pre-training mitigates forgetting in language and vision.
-arXiv preprint arXiv:2205.09357 (2022).
-  • Coucke et al.
-(2018) Alice Coucke, Alaa Saade, Adrien Ball, Théodore Bluche, Alexandre Caulier, David Leroy, Clément Doumouro, Thibault Gisselbrecht, Francesco Caltagirone, Thibaut Lavril, et al.
-2018.
-Snips voice platform: an embedded spoken language understanding system for private-by-design voice interfaces.
-arXiv preprint arXiv:1805.10190 (2018).
-  • Cui et al. (2024) Zhenyu Cui, Yuxin Peng, Xun Wang, Manyu Zhu, and Jiahuan Zhou. 2024. Continual Vision-Language Retrieval via Dynamic Knowledge Rectification. In AAAI, Vol. 38. 11704–11712.
-  • Davari et al.
-(2022) MohammadReza Davari, Nader Asadi, Sudhir Mudur, Rahaf Aljundi, and Eugene Belilovsky.
-2022.
-Probing representation forgetting in supervised and unsupervised continual learning.
-In CVPR.
-16712–16721.
-  • De Lange et al.
-(2021) Matthias De Lange, Rahaf Aljundi, Marc Masana, Sarah Parisot, Xu Jia, Aleš Leonardis, Gregory Slabaugh, and Tinne Tuytelaars.
-2021.
-A continual learning survey: Defying forgetting in classification tasks.
-TPAMI 44, 7 (2021), 3366–3385.
-  • de Masson D’Autume et al. (2019) Cyprien de Masson D’Autume, Sebastian Ruder, Lingpeng Kong, and Dani Yogatama. 2019. Episodic memory in lifelong language learning. NeurIPS 32 (2019).
-  • Della Libera et al.
-(2023) Luca Della Libera, Pooneh Mousavi, Salah Zaiem, Cem Subakan, and Mirco Ravanelli.
-2023.
-CL-MASR: A Continual Learning Benchmark for Multilingual ASR.
-arXiv preprint arXiv:2310.16931 (2023).
-  • Derakhshani et al.
-(2022) Mohammad Mahdi Derakhshani, Ivona Najdenkoska, Tom van Sonsbeek, Xiantong Zhen, Dwarikanath Mahapatra, Marcel Worring, and Cees GM Snoek.
-2022.
-Lifelonger: A benchmark for continual disease classification.
-In International Conference on Medical Image Computing and Computer-Assisted Intervention.
-314–324.
-  • Ding et al. (2008) Xiaowen Ding, Bing Liu, and Philip S Yu. 2008. A holistic lexicon-based approach to opinion mining. In WSDM. 231–240.
-  • Dong et al.
-(2022) Qingxiu Dong, Lei Li, Damai Dai, Ce Zheng, Zhiyong Wu, Baobao Chang, Xu Sun, Jingjing Xu, and Zhifang Sui.
-2022.
-A survey on in-context learning.
-arXiv preprint arXiv:2301.00234 (2022).
-  • Douillard et al.
-(2020) Arthur Douillard, Matthieu Cord, Charles Ollion, Thomas Robert, and Eduardo Valle.
-2020.
-Podnet: Pooled outputs distillation for small-tasks incremental learning.
-In ECCV.
-86–102.
-  • Du et al. (2023) Mingzhe Du, Anh Tuan Luu, Bin Ji, and See-kiong Ng. 2023. From Static to Dynamic: A Continual Learning Framework for Large Language Models. arXiv preprint arXiv:2310.14248 (2023).
-  • Du et al. (2022) Yifan Du, Zikang Liu, Junyi Li, and Wayne Xin Zhao. 2022. A Survey of Vision-Language Pre-Trained Models. In IJCAI. 5436–5443.
-  • Dwork (2006) Cynthia Dwork. 2006. Differential privacy. In International colloquium on automata, languages, and programming. 1–12.
-  • Ethayarajh et al. (2022) Kawin Ethayarajh, Yejin Choi, and Swabha Swayamdipta. 2022. Understanding dataset difficulty with V-usable information. In ICML. 5988–6008.
-  • Faber et al.
-(2023) Kamil Faber, Dominik Zurek, Marcin Pietron, Nathalie Japkowicz, Antonio Vergari, and Roberto Corizzo.
-2023.
-From MNIST to ImageNet and Back: Benchmarking Continual Curriculum Learning.
-arXiv preprint arXiv:2303.11076 (2023).
-  • Fini et al. (2020) Enrico Fini, Stéphane Lathuiliere, Enver Sangineto, Moin Nabi, and Elisa Ricci. 2020. Online continual learning under extreme memory constraints. In ECCV 2020. 720–735.
-  • Geng et al.
-(2021) Binzong Geng, Fajie Yuan, Qiancheng Xu, Ying Shen, Ruifeng Xu, and Min Yang.
-2021.
-Continual Learning for Task-oriented Dialogue System with Iterative Network Pruning, Expanding and Masking.
-In ACL.
-517–523.
-  • Gogoulou et al. (2023) Evangelia Gogoulou, Timothée Lesort, Magnus Boman, and Joakim Nivre. 2023. A study of continual learning under language shift. arXiv preprint arXiv:2311.01200 (2023).
-  • Goldreich (1998) Oded Goldreich. 1998. Secure multi-party computation. Manuscript. Preliminary version 78, 110 (1998), 1–108.
-  • Goyal et al.
-(2017) Yash Goyal, Tejas Khot, Douglas Summers-Stay, Dhruv Batra, and Devi Parikh.
-2017.
-Making the v in vqa matter: Elevating the role of image understanding in visual question answering.
-In CVPR.
-6904–6913.
-  • Grubinger et al.
-(2006) Michael Grubinger, Paul Clough, Henning Müller, and Thomas Deselaers.
-2006.
-The iapr tc-12 benchmark: A new evaluation resource for visual information systems.
-In International workshop ontoImage, Vol.
-2.
-  • Gu et al. (2022) Yanan Gu, Xu Yang, Kun Wei, and Cheng Deng. 2022. Not just selection, but exploration: Online class-incremental continual learning via dual view consistency. In CVPR. 7442–7451.
-  • Gunasekara et al. (2023) Nuwan Gunasekara, Bernhard Pfahringer, Heitor Murilo Gomes, and Albert Bifet. 2023. Survey on Online Streaming Continual Learning. In IJCAI.
-  • Guo et al. (2022) Yiduo Guo, Bing Liu, and Dongyan Zhao. 2022. Online continual learning through mutual information maximization. In ICML. 8109–8126.
-  • Guo et al. (2023) Yiduo Guo, Bing Liu, and Dongyan Zhao. 2023. Dealing with Cross-Task Class Discrimination in Online Continual Learning. In CVPR. 11878–11887.
-  • Gupta et al.
-(2018) Sonal Gupta, Rushin Shah, Mrinal Mohit, Anuj Kumar, and Mike Lewis.
-2018.
-Semantic parsing for task oriented dialog using hierarchical representations.
-arXiv preprint arXiv:1810.07942 (2018).
-  • Gururangan et al.
-(2022) Suchin Gururangan, Mike Lewis, Ari Holtzman, Noah A Smith, and Luke Zettlemoyer.
-2022.
-DEMix Layers: Disentangling Domains for Modular Language Modeling.
-In NAACL.
-5557–5576.
-  • Han et al.
-(2018) Xu Han, Hao Zhu, Pengfei Yu, Ziyun Wang, Yuan Yao, Zhiyuan Liu, and Maosong Sun.
-2018.
-FewRel: A large-scale supervised few-shot relation classification dataset with state-of-the-art evaluation.
-arXiv preprint arXiv:1810.10147 (2018).
-  • Han et al. (2024) Zeyu Han, Chao Gao, Jinyang Liu, Sai Qian Zhang, et al. 2024. Parameter-Efficient Fine-Tuning for Large Models: A Comprehensive Survey. arXiv preprint arXiv:2403.14608 (2024).
-  • Hayes et al.
-(2021) Tyler L Hayes, Giri P Krishnan, Maxim Bazhenov, Hava T Siegelmann, Terrence J Sejnowski, and Christopher Kanan.
-2021.
-Replay in deep learning: Current approaches and missing biological elements.
-Neural computation 33, 11 (2021), 2908–2950.
-  • He et al. (2020) Jiangpeng He, Runyu Mao, Zeman Shao, and Fengqing Zhu. 2020. Incremental learning in online scenario. In CVPR. 13926–13935.
-  • He and McAuley (2016) Ruining He and Julian McAuley. 2016. Ups and downs: Modeling the visual evolution of fashion trends with one-class collaborative filtering. In WWW. 507–517.
-  • Holla et al.
-(2020) Nithin Holla, Pushkar Mishra, Helen Yannakoudakis, and Ekaterina Shutova.
-2020.
-Meta-learning with sparse experience replay for lifelong language learning.
-arXiv preprint arXiv:2009.04891 (2020).
-  • Houlsby et al.
-(2019) Neil Houlsby, Andrei Giurgiu, Stanislaw Jastrzebski, Bruna Morrone, Quentin De Laroussilhe, Andrea Gesmundo, Mona Attariyan, and Sylvain Gelly.
-2019.
-Parameter-efficient transfer learning for NLP.
-In ICML.
-2790–2799.
-  • Hovy et al. (2006) Eduard Hovy, Mitchell Marcus, Martha Palmer, Lance Ramshaw, and Ralph Weischedel. 2006. OntoNotes: The 90% Solution. In NAACL. 57–60.
-  • Hu et al. (2022) Hexiang Hu, Ozan Sener, Fei Sha, and Vladlen Koltun. 2022. Drinking from a firehose: Continual learning with web-scale natural language. TPAMI 45, 5 (2022), 5684–5696.
-  • Hu and Liu (2004) Minqing Hu and Bing Liu. 2004. Mining and summarizing customer reviews. In SIGKDD. 168–177.
-  • Huang et al.
-(2021) Yufan Huang, Yanzhe Zhang, Jiaao Chen, Xuezhi Wang, and Diyi Yang.
-2021.
-Continual Learning for Text Classification with Information Disentanglement Based Regularization.
-In ACL.
-2736–2746.
-  • Husain et al.
-(2019) Hamel Husain, Ho-Hsiang Wu, Tiferet Gazit, Miltiadis Allamanis, and Marc Brockschmidt.
-2019.
-Codesearchnet challenge: Evaluating the state of semantic code search.
-arXiv preprint arXiv:1909.09436 (2019).
-  • Iyer et al. (2018) Srinivasan Iyer, Ioannis Konstas, Alvin Cheung, and Luke Zettlemoyer. 2018. Mapping language to code in programmatic context. arXiv preprint arXiv:1808.09588 (2018).
-  • Jang et al.
-(2023) Joel Jang, Seungone Kim, Seonghyeon Ye, Doyoung Kim, Lajanugen Logeswaran, Moontae Lee, Kyungjae Lee, and Minjoon Seo.
-2023.
-Exploring the benefits of training expert language models over instruction tuning.
-In ICML.
-14702–14729.
-  • Jang et al.
-(2022a) Joel Jang, Seonghyeon Ye, Changho Lee, Sohee Yang, Joongbo Shin, Janghoon Han, Gyeonghun Kim, and Minjoon Seo.
-2022a.
-TemporalWiki: A Lifelong Benchmark for Training and Evaluating Ever-Evolving Language Models.
-In EMNLP.
-6237–6250.
-  • Jang et al.
-(2022b) Joel Jang, Seonghyeon Ye, Sohee Yang, Joongbo Shin, Janghoon Han, Gyeonghun Kim, Stanley Jungkyu Choi, and Minjoon Seo.
-2022b.
-Towards Continual Knowledge Learning of Language Models.
-In ICLR.
-  • Javed and White (2019) Khurram Javed and Martha White. 2019. Meta-learning representations for continual learning. NeurIPS 32 (2019).
-  • Jha et al. (2024) Saurav Jha, Dong Gong, and Lina Yao. 2024. CLAP4CLIP: Continual Learning with Probabilistic Finetuning for Vision-Language Models. arXiv preprint arXiv:2403.19137 (2024).
-  • Jin et al.
-(2021a) Xisen Jin, Bill Yuchen Lin, Mohammad Rostami, and Xiang Ren.
-2021a.
-Learn Continually, Generalize Rapidly: Lifelong Knowledge Accumulation for Few-shot Learning.
-In EMNLP.
-714–729.
-  • Jin et al. (2021b) Xisen Jin, Arka Sadhu, Junyi Du, and Xiang Ren. 2021b. Gradient-based editing of memory examples for online task-free continual learning. NeurIPS 34 (2021), 29193–29205.
-  • Jin et al.
-(2022) Xisen Jin, Dejiao Zhang, Henghui Zhu, Wei Xiao, Shang-Wen Li, Xiaokai Wei, Andrew Arnold, and Xiang Ren.
-2022.
-Lifelong Pretraining: Continually Adapting Language Models to Emerging Corpora.
-In NAACL.
-4764–4780.
-  • Joshi et al. (2017) Mandar Joshi, Eunsol Choi, Daniel Weld, and Luke Zettlemoyer. 2017. TriviaQA: A Large Scale Distantly Supervised Challenge Dataset for Reading Comprehension. In ACL. 1601–1611.
-  • Kafle and Kanan (2017) Kushal Kafle and Christopher Kanan. 2017. An analysis of visual question answering algorithms. In ICCV. 1965–1973.
-  • Ke et al. (2022) Zixuan Ke, Haowei Lin, Yijia Shao, Hu Xu, Lei Shu, and Bing Liu. 2022. Continual Training of Language Models for Few-Shot Learning. In EMNLP. 10205–10216.
-  • Ke and Liu (2022) Zixuan Ke and Bing Liu. 2022. Continual learning of natural language processing tasks: A survey. arXiv preprint arXiv:2211.12701 (2022).
-  • Ke et al. (2021a) Zixuan Ke, Bing Liu, Nianzu Ma, Hu Xu, and Lei Shu. 2021a. Achieving Forgetting Prevention and Knowledge Transfer in Continual Learning. In NeurIPS, Vol. 34. 22443–22456.
-  • Ke et al. (2021b) Zixuan Ke, Bing Liu, Hu Xu, and Lei Shu. 2021b. CLASSIC: Continual and Contrastive Learning of Aspect Sentiment Classification Tasks. In EMNLP.
-  • Ke et al. (2023) Zixuan Ke, Yijia Shao, Haowei Lin, Tatsuya Konishi, Gyuhak Kim, and Bing Liu. 2023. Continual Pre-training of Language Models. In ICLR.
-  • Ke and Xu (2021) Zixuan Ke and Hu Xu. 2021. Adapting BERT for Continual Learning of a Sequence of Aspect Sentiment Classification Tasks. In NAACL.
-  • Kemker et al. (2018) Ronald Kemker, Marc McClure, Angelina Abitino, Tyler Hayes, and Christopher Kanan. 2018. Measuring catastrophic forgetting in neural networks. In AAAI, Vol. 32.
-  • Kenton and Toutanova (2019) Jacob Devlin Ming-Wei Chang Kenton and Lee Kristina Toutanova.
-2019.
-BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding.
-In NAACL.
-4171–4186.
-  • Khan et al.
-(2023) Muhammad Gul Zain Ali Khan, Muhammad Ferjad Naeem, Luc Van Gool, Didier Stricker, Federico Tombari, and Muhammad Zeshan Afzal.
-2023.
-Introducing language guidance in prompt-based continual learning.
-In ICCV.
-11463–11473.
-  • Kim et al.
-(2024) Junsu Kim, Yunhoe Ku, Jihyeon Kim, Junuk Cha, and Seungryul Baek.
-2024.
-VLM-PL: Advanced Pseudo Labeling approach Class Incremental Object Detection with Vision-Language Model.
-arXiv preprint arXiv:2403.05346 (2024).
-  • Kim et al. (2023) Sein Kim, Namkyeong Lee, Donghyun Kim, Minchul Yang, and Chanyoung Park. 2023. Task Relation-aware Continual User Representation Learning. In SIGKDD. 1107–1119.
-  • Kirkpatrick et al.
-(2017) James Kirkpatrick, Razvan Pascanu, Neil Rabinowitz, Joel Veness, Guillaume Desjardins, Andrei A Rusu, Kieran Milan, John Quan, Tiago Ramalho, Agnieszka Grabska-Barwinska, et al.
-2017.
-Overcoming catastrophic forgetting in neural networks.
-the national academy of sciences 114, 13 (2017), 3521–3526.
-  • Ko et al.
-(2022) Jihoon Ko, Shinhwan Kang, and Kijung Shin.
-2022.
-BeGin: Extensive Benchmark Scenarios and An Easy-to-use Framework for Graph Continual Learning.
-arXiv preprint arXiv:2211.14568 (2022).
-  • Koh et al.
-(2021) Hyunseo Koh, Dahyun Kim, Jung-Woo Ha, and Jonghyun Choi.
-2021.
-Online continual learning on class incremental blurry task configuration with anytime inference.
-arXiv preprint arXiv:2110.10031 (2021).
-  • Koh et al.
-(2022) Hyunseo Koh, Minhyuk Seo, Jihwan Bang, Hwanjun Song, Deokki Hong, Seulki Park, Jung-Woo Ha, and Jonghyun Choi.
-2022.
-Online Boundary-Free Continual Learning by Scheduled Data Prior.
-In ICLR.
-  • Kowsari et al.
-(2017) Kamran Kowsari, Donald E Brown, Mojtaba Heidarysafa, Kiana Jafari Meimandi, Matthew S Gerber, and Laura E Barnes.
-2017.
-Hdltex: Hierarchical deep learning for text classification.
-In ICMLA.
-364–371.
-  • Lang (1995) Ken Lang. 1995. Newsweeder: Learning to filter netnews. In Machine learning proceedings. 331–339.
-  • Larson et al.
-(2019) Stefan Larson, Anish Mahendran, Joseph J Peper, Christopher Clarke, Andrew Lee, Parker Hill, Jonathan K Kummerfeld, Kevin Leach, Michael A Laurenzano, Lingjia Tang, et al.
-2019.
-An evaluation dataset for intent classification and out-of-scope prediction.
-arXiv preprint arXiv:1909.02027 (2019).
-  • Lee et al. (2023) Kuan-Ying Lee, Yuanyi Zhong, and Yu-Xiong Wang. 2023. Do pre-trained models benefit equally in continual learning?. In WCACV. 6485–6493.
-  • Lee (2017) Sungjin Lee. 2017. Toward continual learning for conversational agents. arXiv preprint arXiv:1712.09943 (2017).
-  • Lesort et al.
-(2020) Timothée Lesort, Vincenzo Lomonaco, Andrei Stoian, Davide Maltoni, David Filliat, and Natalia Díaz-Rodríguez.
-2020.
-Continual learning for robotics: Definition, framework, learning strategies, opportunities and challenges.
-Information fusion 58 (2020), 52–68.
-  • Lewis et al.
-(2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer.
-2020.
-BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension.
-In ACL.
-  • Li et al.
-(2022a) Dingcheng Li, Zheng Chen, Eunah Cho, Jie Hao, Xiaohu Liu, Fan Xing, Chenlei Guo, and Yang Liu.
-2022a.
-Overcoming catastrophic forgetting during domain adaptation of seq2seq language generation.
-In NAACL.
-5441–5454.
-  • Li et al.
-(2017) Feng-Lin Li, Minghui Qiu, Haiqing Chen, Xiongwei Wang, Xing Gao, Jun Huang, Juwei Ren, Zhongzhou Zhao, Weipeng Zhao, Lei Wang, et al.
-2017.
-Alime assist: An intelligent assistant for creating an innovative e-commerce experience.
-In CIKM.
-2495–2498.
-  • Li et al. (2022b) Guodun Li, Yuchen Zhai, Qianglong Chen, Xing Gao, Ji Zhang, and Yin Zhang. 2022b. Continual few-shot intent detection. In COLING. 333–343.
-  • Li et al.
-(2019) Liunian Harold Li, Mark Yatskar, Da Yin, Cho-Jui Hsieh, and Kai-Wei Chang.
-2019.
-Visualbert: A simple and performant baseline for vision and language.
-arXiv preprint arXiv:1908.03557 (2019).
-  • Lin et al.
-(2023) Huiwei Lin, Baoquan Zhang, Shanshan Feng, Xutao Li, and Yunming Ye.
-2023.
-PCR: Proxy-based Contrastive Replay for Online Class-Incremental Continual Learning.
-In CVPR.
-24246–24255.
-  • Lin et al.
-(2014) Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C Lawrence Zitnick.
-2014.
-Microsoft coco: Common objects in context.
-In ECCV.
-740–755.
-  • Lin et al. (2021) Zhiqiu Lin, Jia Shi, Deepak Pathak, and Deva Ramanan. 2021. The clear benchmark: Continual learning on real-world imagery. In NeurIPS.
-  • Liu (2020) Bing Liu. 2020. Learning on the job: Online lifelong and continual learning. In AAAI, Vol. 34. 13544–13549.
-  • Liu and Mazumder (2021) Bing Liu and Sahisnu Mazumder. 2021. Lifelong and continual learning dialogue systems: learning during conversation. In AAAI, Vol. 35. 15058–15063.
-  • Liu et al. (2023b) Bing Liu, Sahisnu Mazumder, Eric Robertson, and Scott Grigsby. 2023b. AI Autonomy: Self-initiated Open-world Continual Learning and Adaptation. AI Magazine (2023).
-  • Liu et al. (2024) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2024. Visual instruction tuning. NeurIPS 36 (2024).
-  • Liu et al. (2015) Qian Liu, Zhiqiang Gao, Bing Liu, and Yuanlin Zhang. 2015. Automated rule selection for aspect extraction in opinion mining. In IJCAI.
-  • Liu et al. (2021b) Qingbin Liu, Xiaoyan Yu, Shizhu He, Kang Liu, and Jun Zhao. 2021b. Lifelong intent detection via multi-strategy rebalancing. arXiv preprint arXiv:2108.04445 (2021).
-  • Liu et al. (2019b) Tianlin Liu, Lyle Ungar, and João Sedoc. 2019b. Continual Learning for Sentence Representations Using Conceptors. In NAACL. 3274–3279.
-  • Liu et al.
-(2023a) Xialei Liu, Xusheng Cao, Haori Lu, Jia-wen Xiao, Andrew D Bagdanov, and Ming-Ming Cheng.
-2023a.
-Class Incremental Learning with Pre-trained Vision-Language Models.
-arXiv preprint arXiv:2310.20348 (2023).
-  • Liu et al.
-(2021a) Xingkun Liu, Arash Eshghi, Pawel Swietojanski, and Verena Rieser.
-2021a.
-Benchmarking natural language understanding services for building conversational agents.
-In International Workshop on Spoken Dialogue Systems.
-165–183.
-  • Liu et al.
-(2022) Xiao Liu, Kaixuan Ji, Yicheng Fu, Weng Tam, Zhengxiao Du, Zhilin Yang, and Jie Tang.
-2022.
-P-Tuning: Prompt Tuning Can Be Comparable to Fine-tuning Across Scales and Tasks.
-In ACL.
-61–68.
-  • Liu et al.
-(2019a) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov.
-2019a.
-Roberta: A robustly optimized bert pretraining approach.
-arXiv preprint arXiv:1907.11692 (2019).
-  • Lo et al. (2020) Kyle Lo, Lucy Lu Wang, Mark Neumann, Rodney Kinney, and Daniel Weld. 2020. S2ORC: The Semantic Scholar Open Research Corpus. In ACL. 4969–4983.
-  • Lopez-Paz and Ranzato (2017) David Lopez-Paz and Marc’Aurelio Ranzato. 2017. Gradient episodic memory for continual learning. NeurIPS 30 (2017).
-  • Lu et al.
-(2021) Shuai Lu, Daya Guo, Shuo Ren, Junjie Huang, Alexey Svyatkovskiy, Ambrosio Blanco, Colin Clement, Dawn Drain, Daxin Jiang, Duyu Tang, et al.
-2021.
-Codexglue: A machine learning benchmark dataset for code understanding and generation.
-arXiv preprint arXiv:2102.04664 (2021).
-  • Lu et al.
-(2017) Xiaoqiang Lu, Binqiang Wang, Xiangtao Zheng, and Xuelong Li.
-2017.
-Exploring models and data for remote sensing image caption generation.
-IEEE Transactions on Geoscience and Remote Sensing 56, 4 (2017), 2183–2195.
-  • Luo et al.
-(2023) Yun Luo, Zhen Yang, Fandong Meng, Yafu Li, Jie Zhou, and Yue Zhang.
-2023.
-An Empirical Study of Catastrophic Forgetting in Large Language Models During Continual Fine-tuning.
-arXiv e-prints (2023).
-  • Ma et al.
-(2023) Shirong Ma, Shen Huang, Shulin Huang, Xiaobin Wang, Yangning Li, Hai-Tao Zheng, Pengjun Xie, Fei Huang, and Yong Jiang.
-2023.
-EcomGPT-CT: Continual pre-training of e-commerce large language models with semi-structured data.
-arXiv preprint arXiv:2312.15696 (2023).
-  • Maas et al. (2011) Andrew L. Maas, Raymond E. Daly, Peter T. Pham, Dan Huang, Andrew Y. Ng, and Christopher Potts. 2011. Learning Word Vectors for Sentiment Analysis. In ACL. 142–150.
-  • Madotto et al.
-(2020) Andrea Madotto, Zhaojiang Lin, Zhenpeng Zhou, Seungwhan Moon, Paul Crook, Bing Liu, Zhou Yu, Eunjoon Cho, and Zhiguang Wang.
-2020.
-Continual learning in task-oriented dialogue systems.
-arXiv preprint arXiv:2012.15504 (2020).
-  • Maekawa et al.
-(2023) Aru Maekawa, Hidetaka Kamigaito, Kotaro Funakoshi, and Manabu Okumura.
-2023.
-Generative Replay Inspired by Hippocampal Memory Indexing for Continual Language Learning.
-In EACL.
-930–942.
-  • Mai et al.
-(2022) Zheda Mai, Ruiwen Li, Jihwan Jeong, David Quispe, Hyunwoo Kim, and Scott Sanner.
-2022.
-Online continual learning in image classification: An empirical survey.
-Neurocomputing 469 (2022), 28–51.
-  • Masana et al.
-(2022) Marc Masana, Xialei Liu, Bartłomiej Twardowski, Mikel Menta, Andrew D Bagdanov, and Joost Van De Weijer.
-2022.
-Class-incremental learning: survey and performance evaluation on image classification.
-TPAMI 45, 5 (2022), 5513–5533.
-  • Mazumder and Liu (2024) Sahisnu Mazumder and Bing Liu. 2024. Lifelong and Continual Learning Dialogue Systems. Springer Nature.
-  • McCann et al.
-([n.
-d.]) Bryan McCann, Nitish Shirish Keskar, Caiming Xiong, and Richard Socher.
-[n.
-d.].
-The natural language decathlon: Multitask learning as question answering.
-arXiv 2018.
-arXiv preprint arXiv:1806.08730 ([n.
-d.]).
-  • Mehta et al. (2021) Sanket Vaibhav Mehta, Darshan Patil, Sarath Chandar, and Emma Strubell. 2021. An Empirical Investigation of the Role of Pre-training in Lifelong Learning. (2021).
-  • Michieli et al.
-(2023) Umberto Michieli, Pablo Peso Parada, and Mete Ozay.
-2023.
-Online Continual Learning in Keyword Spotting for Low-Resource Devices via Pooling High-Order Temporal Statistics.
-arXiv preprint arXiv:2307.12660 (2023).
-  • Min et al.
-(2023) Bonan Min, Hayley Ross, Elior Sulem, Amir Pouran Ben Veyseh, Thien Huu Nguyen, Oscar Sainz, Eneko Agirre, Ilana Heintz, and Dan Roth.
-2023.
-Recent advances in natural language processing via large pre-trained language models: A survey.
-Comput.
-Surveys 56, 2 (2023), 1–40.
-  • Mok et al.
-(2023) Jisoo Mok, Jaeyoung Do, Sungjin Lee, Tara Taghavi, Seunghak Yu, and Sungroh Yoon.
-2023.
-Large-scale Lifelong Learning of In-context Instructions and How to Tackle It.
-In ACL.
-12573–12589.
-  • Monaikul et al. (2021) Natawut Monaikul, Giuseppe Castellucci, Simone Filice, and Oleg Rokhlenko. 2021. Continual learning for named entity recognition. In AAAI, Vol. 35. 13570–13577.
-  • Moon et al.
-(2023) Jun-Yeong Moon, Keon-Hee Park, Jung Uk Kim, and Gyeong-Moon Park.
-2023.
-Online Class Incremental Learning on Stochastic Blurry Task Boundary via Mask and Visual Prompt Tuning.
-In ICCV.
-11731–11741.
-  • Mundra et al.
-(2024) Nandini Mundra, Sumanth Doddapaneni, Raj Dabre, Anoop Kunchukuttan, Ratish Puduppully, and Mitesh M Khapra.
-2024.
-A Comprehensive Analysis of Adapter Efficiency.
-In IKDD.
-136–154.
-  • Ni et al. (2023) Zixuan Ni, Longhui Wei, Siliang Tang, Yueting Zhuang, and Qi Tian. 2023. Continual vision-language representation learning with off-diagonal information. In ICML. 26129–26149.
-  • Novikova et al. (2017) Jekaterina Novikova, Ondřej Dušek, and Verena Rieser. 2017. The E2E dataset: New challenges for end-to-end generation. arXiv preprint arXiv:1706.09254 (2017).
-  • Ororbia et al.
-(2022) Alex Ororbia, Ankur Mali, C Lee Giles, and Daniel Kifer.
-2022.
-Lifelong neural predictive coding: Learning cumulatively online without forgetting.
-NeurIPS 35 (2022), 5867–5881.
-  • Ouyang et al.
-(2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al.
-2022.
-Training language models to follow instructions with human feedback.
-NeurIPS 35 (2022), 27730–27744.
-  • Parisi et al.
-(2019) German I Parisi, Ronald Kemker, Jose L Part, Christopher Kanan, and Stefan Wermter.
-2019.
-Continual lifelong learning with neural networks: A review.
-Neural networks 113 (2019), 54–71.
-  • PENG et al. (2024) Bohao PENG, Zhuotao Tian, Shu Liu, Ming-Chang Yang, and Jiaya Jia. 2024. Scalable Language Model with Generalized Continual Learning. In ICLR.
-  • Poth et al.
-(2023) Clifton Poth, Hannah Sterz, Indraneil Paul, Sukannya Purkayastha, Leon Engländer, Timo Imhof, Ivan Vulić, Sebastian Ruder, Iryna Gurevych, and Jonas Pfeiffer.
-2023.
-Adapters: A Unified Library for Parameter-Efficient and Modular Transfer Learning.
-In EMNLP.
-149–160.
-  • Qian et al.
-(2023) Zi Qian, Xin Wang, Xuguang Duan, Pengda Qin, Yuhong Li, and Wenwu Zhu.
-2023.
-Decouple before interact: Multi-modal prompt learning for continual visual question answering.
-In ICCV.
-2953–2962.
-  • Qin and Joty (2022a) Chengwei Qin and Shafiq Joty. 2022a. Continual Few-shot Relation Learning via Embedding Space Regularization and Data Augmentation. In ACL. 2776–2789.
-  • Qin and Joty (2022b) Chengwei Qin and Shafiq Joty. 2022b. LFPT5: A Unified Framework for Lifelong Few-shot Language Learning Based on Prompt Tuning of T5. In ICLR.
-  • Qin et al.
-(2023) Yujia Qin, Cheng Qian, Xu Han, Yankai Lin, Huadong Wang, Ruobing Xie, Zhiyuan Liu, Maosong Sun, and Jie Zhou.
-2023.
-Recyclable Tuning for Continual Pre-training.
-In ACL.
-11403–11426.
-  • Qin et al. (2022) Yujia Qin, Jiajie Zhang, Yankai Lin, Zhiyuan Liu, Peng Li, Maosong Sun, and Jie Zhou. 2022. ELLE: Efficient Lifelong Pre-training for Emerging Data. In ACL. 2789–2810.
-  • Qu et al.
-(2021) Haoxuan Qu, Hossein Rahmani, Li Xu, Bryan Williams, and Jun Liu.
-2021.
-Recent advances of continual learning in computer vision: An overview.
-arXiv preprint arXiv:2109.11369 (2021).
-  • Radford et al.
-(2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al.
-2021.
-Learning transferable visual models from natural language supervision.
-In ICML.
-PMLR, 8748–8763.
-  • Rajpurkar et al. (2016) Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016. SQuAD: 100,000+ Questions for Machine Comprehension of Text. In EMNLP. 2383–2392.
-  • Ramesh et al.
-(2021) Aditya Ramesh, Mikhail Pavlov, Gabriel Goh, Scott Gray, Chelsea Voss, Alec Radford, Mark Chen, and Ilya Sutskever.
-2021.
-Zero-shot text-to-image generation.
-In ICML.
-Pmlr, 8821–8831.
-  • Rastogi et al.
-(2020) Abhinav Rastogi, Xiaoxue Zang, Srinivas Sunkara, Raghav Gupta, and Pranav Khaitan.
-2020.
-Towards scalable multi-domain conversational agents: The schema-guided dialogue dataset.
-In AAAI, Vol.
-34.
-8689–8696.
-  • Razdaibiedina et al.
-(2023) Anastasia Razdaibiedina, Yuning Mao, Rui Hou, Madian Khabsa, Mike Lewis, and Amjad Almahairi.
-2023.
-Progressive Prompts: Continual Learning for Language Models.
-In ICLR.
-  • Rebuffi et al. (2017) Sylvestre-Alvise Rebuffi, Alexander Kolesnikov, Georg Sperl, and Christoph H. Lampert. 2017. iCaRL: Incremental Classifier and Representation Learning. In CVPR.
-  • Sang and De Meulder (2003) Erik F Sang and Fien De Meulder. 2003. Introduction to the CoNLL-2003 shared task: Language-independent named entity recognition. arXiv preprint cs/0306050 (2003).
-  • Scao et al.
-(2022) Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, Matthias Gallé, et al.
-2022.
-Bloom: A 176b-parameter open-access multilingual language model.
-arXiv preprint arXiv:2211.05100 (2022).
-  • Schuster et al.
-(2018) Sebastian Schuster, Sonal Gupta, Rushin Shah, and Mike Lewis.
-2018.
-Cross-lingual transfer learning for multilingual task oriented dialog.
-arXiv preprint arXiv:1810.13327 (2018).
-  • Scialom et al. (2022a) Thomas Scialom, Tuhin Chakrabarty, and Smaranda Muresan. 2022a. Fine-tuned Language Models are Continual Learners. In EMNLP. 6107–6122.
-  • Scialom et al. (2022b) Thomas Scialom, Tuhin Chakrabarty, and Smaranda Muresan. 2022b. Fine-tuned language models are continual learners. In EMNLP. 6107–6122.
-  • See et al. (2017) Abigail See, Peter J Liu, and Christopher D Manning. 2017. Get to the point: Summarization with pointer-generator networks. arXiv preprint arXiv:1704.04368 (2017).
-  • Semola et al.
-(2022) Rudy Semola, Vincenzo Lomonaco, and Davide Bacciu.
-2022.
-Continual-learning-as-a-service (claas): On-demand efficient adaptation of predictive models.
-arXiv preprint arXiv:2206.06957 (2022).
-  • Shaheen et al.
-(2022) Khadija Shaheen, Muhammad Abdullah Hanif, Osman Hasan, and Muhammad Shafique.
-2022.
-Continual learning for real-world autonomous systems: Algorithms, challenges and frameworks.
-Journal of Intelligent & Robotic Systems 105, 1 (2022), 9.
-  • Shen et al. (2019) Yilin Shen, Xiangyu Zeng, and Hongxia Jin. 2019. A progressive model to enable continual learning for semantic slot filling. In EMNLP. 1279–1284.
-  • Shim et al.
-(2021) Dongsub Shim, Zheda Mai, Jihwan Jeong, Scott Sanner, Hyunwoo Kim, and Jongseong Jang.
-2021.
-Online class-incremental continual learning with adversarial shapley value.
-In AAAI, Vol.
-35.
-9630–9638.
-  • Song et al.
-(2023) Chenyang Song, Xu Han, Zheni Zeng, Kuai Li, Chen Chen, Zhiyuan Liu, Maosong Sun, and Tao Yang.
-2023.
-ConPET: Continual Parameter-Efficient Tuning for Large Language Models.
-arXiv preprint arXiv:2309.14763 (2023).
-  • Sun et al. (2020a) Fan-Keng Sun, Cheng-Hao Ho, and Hung-Yi Lee. 2020a. LAMOL: LAnguage MOdeling for Lifelong Language Learning. In ICLR. OpenReview.net.
-  • Sun et al.
-(2020b) Yu Sun, Shuohuan Wang, Yukun Li, Shikun Feng, Hao Tian, Hua Wu, and Haifeng Wang.
-2020b.
-Ernie 2.0: A continual pre-training framework for language understanding.
-In AAAI, Vol.
-34.
-8968–8975.
-  • Touvron et al.
-(2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al.
-2023.
-Llama: Open and efficient foundation language models.
-arXiv preprint arXiv:2302.13971 (2023).
-  • Tufano et al.
-(2019) Michele Tufano, Cody Watson, Gabriele Bavota, Massimiliano Di Penta, Martin White, and Denys Poshyvanyk.
-2019.
-An empirical study on learning bug-fixing patches in the wild via neural machine translation.
-TOSEM 28, 4 (2019), 1–29.
-  • Van de Ven et al. (2022) Gido M Van de Ven, Tinne Tuytelaars, and Andreas S Tolias. 2022. Three types of incremental learning. Nature Machine Intelligence 4, 12 (2022), 1185–1197.
-  • Vander Eeckt et al. (2023) Steven Vander Eeckt et al. 2023. Rehearsal-Free Online Continual Learning for Automatic Speech Recognition. arXiv e-prints (2023), arXiv–2306.
-  • Varshney et al.
-(2022) Vaibhav Varshney, Mayur Patidar, Rajat Kumar, Lovekesh Vig, and Gautam Shroff.
-2022.
-Prompt augmented generative replay via supervised contrastive learning for lifelong intent detection.
-In NAACL.
-1113–1127.
-  • Verwimp et al.
-(2023) Eli Verwimp, Kuo Yang, Sarah Parisot, Lanqing Hong, Steven McDonagh, Eduardo Pérez-Pellitero, Matthias De Lange, and Tinne Tuytelaars.
-2023.
-Clad: A realistic continual learning benchmark for autonomous driving.
-Neural Networks 161 (2023), 659–669.
-  • Volske et al.
-(2017) Michael Volske, Martin Potthast, Shahbaz Syed, and Benno Stein.
-2017.
-TL;DR: Mining Reddit to Learn Automatic Summarization.
-In the Workshop on New Frontiers in Summarization.
-59–63.
-  • Wang et al.
-(2019a) Alex Wang, Yada Pruksachatkun, Nikita Nangia, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel Bowman.
-2019a.
-Superglue: A stickier benchmark for general-purpose language understanding systems.
-NeurIPS 32 (2019).
-  • Wang et al.
-(2018) Alex Wang, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel Bowman.
-2018.
-GLUE: A Multi-Task Benchmark and Analysis Platform for Natural Language Understanding.
-In EMNLP.
-353–355.
-  • Wang et al.
-(2021a) Chengyu Wang, Haojie Pan, Yuan Liu, Kehan Chen, Minghui Qiu, Wei Zhou, Jun Huang, Haiqing Chen, Wei Lin, and Deng Cai.
-2021a.
-Mell: Large-scale extensible user intent classification for dialogue systems with meta lifelong learning.
-In SIGKDD.
-3649–3659.
-  • Wang et al. (2022a) Han Wang, Ruiliu Fu, Xuejun Zhang, and Jun Zhou. 2022a. RVAE-LAMOL: Residual Variational Autoencoder to Enhance Lifelong Language Learning. In IJCNN. IEEE, 1–9.
-  • Wang et al. (2019b) Hong Wang, Wenhan Xiong, Mo Yu, Xiaoxiao Guo, Shiyu Chang, and William Yang Wang. 2019b. Sentence Embedding Alignment for Lifelong Relation Extraction. In NAACL. 796–806.
-  • Wang et al. (2021b) Jianren Wang, Xin Wang, Yue Shang-Guan, and Abhinav Gupta. 2021b. Wanderlust: Online continual object detection in the real world. In ICCV. 10829–10838.
-  • Wang et al. (2024b) Liyuan Wang, Xingxing Zhang, Hang Su, and Jun Zhu. 2024b. A comprehensive survey of continual learning: Theory, method and application. TPAMI (2024).
-  • Wang et al. (2023d) Quanziang Wang, Renzhen Wang, Yichen Wu, Xixi Jia, and Deyu Meng. 2023d. CBA: Improving Online Continual Learning via Continual Bias Adaptor. In ICCV. 19082–19092.
-  • Wang et al.
-(2023b) Xiao Wang, Guangyao Chen, Guangwu Qian, Pengcheng Gao, Xiao-Yong Wei, Yaowei Wang, Yonghong Tian, and Wen Gao.
-2023b.
-Large-scale multi-modal pre-trained models: A comprehensive survey.
-Machine Intelligence Research 20, 4 (2023), 447–482.
-  • Wang et al.
-(2023a) Xiao Wang, Tianze Chen, Qiming Ge, Han Xia, Rong Bao, Rui Zheng, Qi Zhang, Tao Gui, and Xuanjing Huang.
-2023a.
-Orthogonal Subspace Learning for Language Model Continual Learning.
-In EMNLP.
-10658–10671.
-  • Wang et al.
-(2024a) Xiao Wang, Yuansen Zhang, Tianze Chen, Songyang Gao, Senjie Jin, Xianjun Yang, Zhiheng Xi, Rui Zheng, Yicheng Zou, Tao Gui, Qi Zhang, and Xuanjing Huang.
-2024a.
-TRACE: A Comprehensive Benchmark for Continual Learning in Large Language Models.
-  • Wang et al.
-(2022b) Yabin Wang, Zhiwu Huang, and Xiaopeng Hong.
-2022b.
-S-prompts learning with pre-trained transformers: An occam’s razor for domain incremental learning.
-NeurIPS 35 (2022), 5682–5695.
-  • Wang et al.
-(2022c) Yizhong Wang, Swaroop Mishra, Pegah Alipoormolabashi, Yeganeh Kordi, Amirreza Mirzaei, Anjana Arunkumar, Arjun Ashok, Arut Selvan Dhanasekaran, Atharva Naik, David Stap, et al.
-2022c.
-Super-naturalinstructions: Generalization via declarative instructions on 1600+ nlp tasks.
-arXiv preprint arXiv:2204.07705 (2022).
-  • Wang et al.
-(2023c) Zhicheng Wang, Yufang Liu, Tao Ji, Xiaoling Wang, Yuanbin Wu, Congcong Jiang, Ye Chao, Zhencong Han, Ling Wang, Xu Shao, et al.
-2023c.
-Rehearsal-free Continual Language Learning via Efficient Parameter Isolation.
-In ACL.
-10933–10946.
-  • Wang et al. (2020) Zirui Wang, Sanket Vaibhav Mehta, Barnabas Poczos, and Jaime Carbonell. 2020. Efficient Meta Lifelong-Learning with Limited Memory. In EMNLP. 535–548.
-  • Wei et al. (2023) Yujie Wei, Jiaxin Ye, Zhizhong Huang, Junping Zhang, and Hongming Shan. 2023. Online Prototype Learning for Online Continual Learning. In ICCV. 18764–18774.
-  • Wen et al.
-(2015) Tsung-Hsien Wen, Milica Gasic, Nikola Mrksic, Pei-Hao Su, David Vandyke, and Steve Young.
-2015.
-Semantically conditioned lstm-based natural language generation for spoken dialogue systems.
-arXiv preprint arXiv:1508.01745 (2015).
-  • Wójcik et al.
-(2023) Mateusz Wójcik, Witold Kościukiewicz, Mateusz Baran, Tomasz Kajdanowicz, and Adam Gonczarek.
-2023.
-Domain-Agnostic Neural Architecture for Class Incremental Continual Learning in Document Processing Platform.
-In ACL.
-527–537.
-[…]
-  • Wu et al. (2023) Yuhao Wu, Tongjun Shi, Karthick Sharma, Chun Wei Seah, and Shuhao Zhang. 2023. Online Continual Knowledge Learning for Language Models. arXiv preprint arXiv:2311.09632 (2023).
-  • Xia et al.
-(2021) Congying Xia, Wenpeng Yin, Yihao Feng, and Philip Yu.
-2021.
-Incremental few-shot text classification with multi-round new classes: Formulation, dataset and system.
-arXiv preprint arXiv:2104.11882 (2021).
-  • Xie et al. (2023) Yong Xie, Karan Aggarwal, and Aitzaz Ahmad. 2023. Efficient continual pre-training for building domain specific large language models. arXiv preprint arXiv:2311.08545 (2023).
-  • Xu et al. (2019) Hu Xu, Bing Liu, Lei Shu, and Philip S Yu. 2019. BERT post-training for review reading comprehension and aspect-based sentiment analysis. arXiv preprint arXiv:1904.02232 (2019).
-  • Xu et al.
-(2017) Jiaming Xu, Bo Xu, Peng Wang, Suncong Zheng, Guanhua Tian, and Jun Zhao.
-2017.
-Self-taught convolutional neural networks for short text clustering.
-Neural Networks 88 (2017), 22–31.
-  • Yadav et al.
-(2023) Prateek Yadav, Qing Sun, Hantian Ding, Xiaopeng Li, Dejiao Zhang, Ming Tan, Xiaofei Ma, Parminder Bhatia, Ramesh Nallapati, Murali Krishna Ramanathan, et al.
-2023.
-Exploring continual learning for code generation models.
-arXiv preprint arXiv:2307.02435 (2023).
-  • Yang et al.
-(2021) An Yang, Junyang Lin, Rui Men, Chang Zhou, Le Jiang, Xianyan Jia, Ang Wang, Jie Zhang, Jiamang Wang, Yong Li, et al.
-2021.
-M6-t: Exploring sparse expert models and beyond.
-arXiv preprint arXiv:2105.15082 (2021).
-  • Yi et al.
-(2023) Huahui Yi, Ziyuan Qin, Qicheng Lao, Wei Xu, Zekun Jiang, Dequan Wang, Shaoting Zhang, and Kang Li.
-2023.
-Towards General Purpose Medical AI: Continual Learning Medical Foundation Model.
-arXiv preprint arXiv:2303.06580 (2023).
-  • Yin et al. (2021) Haiyan Yin, Ping Li, et al. 2021. Mitigating forgetting in online continual learning with neuron calibration. NeurIPS 34 (2021), 10260–10272.
-  • Yin et al. (2022) Wenpeng Yin, Jia Li, and Caiming Xiong. 2022. ConTinTin: Continual Learning from Task Instructions. In ACL. 3062–3072.
-  • Yogatama et al.
-(2019) Dani Yogatama, Cyprien de Masson d’Autume, Jerome Connor, Tomas Kocisky, Mike Chrzanowski, Lingpeng Kong, Angeliki Lazaridou, Wang Ling, Lei Yu, Chris Dyer, et al.
-2019.
-Learning and evaluating general linguistic intelligence.
-arXiv preprint arXiv:1901.11373 (2019).
-  • Yoon et al. (2021) Jaehong Yoon, Divyam Madaan, Eunho Yang, and Sung Ju Hwang. 2021. Online coreset selection for rehearsal-based continual learning. arXiv preprint arXiv:2106.01085 (2021).
-  • Young et al.
-(2014) Peter Young, Alice Lai, Micah Hodosh, and Julia Hockenmaier.
-2014.
-From image descriptions to visual denotations: New similarity metrics for semantic inference over event descriptions.
-TACL 2 (2014), 67–78.
-  • Yu et al.
-(2024) Jiazuo Yu, Yunzhi Zhuge, Lu Zhang, Ping Hu, Dong Wang, Huchuan Lu, and You He.
-2024.
-Boosting Continual Learning of Vision-Language Models via Mixture-of-Experts Adapters.
-In CVPR.
-  • Yuan et al.
-(2020) Fajie Yuan, Xiangnan He, Alexandros Karatzoglou, and Liguang Zhang.
-2020.
-Parameter-efficient transfer from sequential behaviors for user modeling and recommendation.
-In SIGIR.
-1469–1478.
-  • Yuan et al.
-(2021) Fajie Yuan, Guoxiao Zhang, Alexandros Karatzoglou, Joemon Jose, Beibei Kong, and Yudong Li.
-2021.
-One Person, One Model, One World: Learning Continual User Representation without Forgetting.
-696–705.
-  • Zellers et al. (2019) Rowan Zellers, Ari Holtzman, Hannah Rashkin, Yonatan Bisk, Ali Farhadi, Franziska Roesner, and Yejin Choi. 2019. Defending against neural fake news. NeurIPS 32 (2019).
-  • Zhang et al. (2021) Chen Zhang, Yu Xie, Hang Bai, Bin Yu, Weihong Li, and Yuan Gao. 2021. A survey on federated learning. Knowledge-Based Systems 216 (2021), 106775.
-  • Zhang et al.
-(2023b) Han Zhang, Lin Gui, Yuanzhao Zhai, Hui Wang, Yu Lei, and Ruifeng Xu.
-2023b.
-Copf: Continual learning human preference through optimal policy fitting.
-arXiv preprint arXiv:2310.15694 (2023).
-  • Zhang et al. (2024) Han Zhang, Yu Lei, Lin Gui, Min Yang, Yulan He, Hui Wang, and Ruifeng Xu. 2024. CPPO: Continual Learning for Reinforcement Learning with Human Feedback. In ICLR.
-  • Zhang and Kim (2023) Peiyan Zhang and Sunghun Kim. 2023. A Survey on Incremental Update for Neural Recommender Systems. arXiv preprint arXiv:2303.02851 (2023).
-  • Zhang et al. (2023d) Xi Zhang, Feifei Zhang, and Changsheng Xu. 2023d. Vqacl: A novel visual question answering continual learning setting. In CVPR. 19102–19112.
-  • Zhang et al. (2015) Xiang Zhang, Junbo Zhao, and Yann LeCun. 2015. Character-level convolutional networks for text classification. NeurIPS 28 (2015).
-  • Zhang et al.
-(2022a) Yaqian Zhang, Bernhard Pfahringer, Eibe Frank, Albert Bifet, Nick Jin Sean Lim, and Yunzhe Jia.
-2022a.
-A simple but strong baseline for online continual learning: Repeated augmented rehearsal.
-NeurIPS 35 (2022), 14771–14783.
-  • Zhang et al. (2022b) Yanzhe Zhang, Xuezhi Wang, and Diyi Yang. 2022b. Continual Sequence Generation with Adaptive Compositional Modules. In ACL. 3653–3667.
-  • Zhang et al.
-(2023c) Yating Zhang, Yexiang Wang, Fei Cheng, Sadao Kurohashi, et al.
-2023c.
-Reformulating Domain Adaptation of Large Language Models as Adapt-Retrieve-Revise.
-arXiv preprint arXiv:2310.03328 (2023).
-  • Zhang et al. (2023a) Zihan Zhang, Meng Fang, Ling Chen, and Mohammad-Reza Namazi-Rad. 2023a. Citb: A benchmark for continual instruction tuning. arXiv preprint arXiv:2310.14510 (2023).
-  • Zhao et al.
-(2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al.
-2023.
-A survey of large language models.
-arXiv preprint arXiv:2303.18223 (2023).
-  • Zhao et al.
-(2022) Yingxiu Zhao, Yinhe Zheng, Zhiliang Tian, Chang Gao, Jian Sun, and Nevin L.
-Zhang.
-2022.
-Prompt Conditioned VAE: Enhancing Generative Replay for Lifelong Learning in Task-Oriented Dialogue.
-In EMNLP.
-11153–11169.
-  • Zheng et al.
-(2023) Zangwei Zheng, Mingyuan Ma, Kai Wang, Ziheng Qin, Xiangyu Yue, and Yang You.
-2023.
-Preventing zero-shot transfer degradation in continual learning of vision-language models.
-In ICCV.
-19125–19136.
-  • Zhong et al.
-(2017) Victor Zhong, Caiming Xiong, and Richard Socher.
-2017.
-Seq2sql: Generating structured queries from natural language using reinforcement learning.
-arXiv preprint arXiv:1709.00103 (2017).
-  • Zhou et al. (2023b) Da-Wei Zhou, Qi-Wei Wang, Zhi-Hong Qi, Han-Jia Ye, De-Chuan Zhan, and Ziwei Liu. 2023b. Deep class-incremental learning: A survey. arXiv preprint arXiv:2302.03648 (2023).
-  • Zhou et al.
-(2023c) Da-Wei Zhou, Yuanhan Zhang, Jingyi Ning, Han-Jia Ye, De-Chuan Zhan, and Ziwei Liu.
-2023c.
-Learning without forgetting for vision-language models.
-arXiv preprint arXiv:2305.19270 (2023).
-  • Zhou et al.
-(2023a) Jie Zhou, Pei Ke, Xipeng Qiu, Minlie Huang, and Junping Zhang.
-2023a.
-ChatGPT: potential, prospects, and limitations.
-Frontiers of Information Technology & Electronic Engineering (2023), 1–6.
-  • Zhu et al.
-(2023) Hongguang Zhu, Yunchao Wei, Xiaodan Liang, Chunjie Zhang, and Yao Zhao.
-2023.
-Ctp: Towards vision-language continual pretraining via compatible momentum contrast and topology preservation.
-In ICCV.
-22257–22267.
-  • Zhu et al. (2022) Qi Zhu, Bing Li, Fei Mi, Xiaoyan Zhu, and Minlie Huang. 2022. Continual Prompt Tuning for Dialog State Tracking. In ACL. 1124–1137.
-  • Zhu et al.
-(2015) Yukun Zhu, Ryan Kiros, Rich Zemel, Ruslan Salakhutdinov, Raquel Urtasun, Antonio Torralba, and Sanja Fidler.
-2015.
-Aligning books and movies: Towards story-like visual explanations by watching movies and reading books.
-In ICCV.
-19–27.
 Appendix A Details of Metrics
@@
 The aforementioned metrics are utilized to assess key capabilities of PLMs, such as instruction following and safety, as discussed in Wang et al.
-(Wang et al., 2024a).
+(Wang et al., 2024b).
 A.2. Memory Stability.